### v1

# The Fraction of Clumpy Galaxies in JWST Surveys over $2<z<12$

Abstract
High-redshift galaxies exhibit compact regions of intense star formation, known as “clumps,” which are conspicuous in the rest-frame ultraviolet. Studying them can shed light on how they form and evolve and inform theoretical models of galaxy evolution. We examine the evolution of clumpy galaxies with redshift and stellar mass over $2<z<12$ with James Webb Space Telescope (JWST) imaging from the JWST Advanced Extragalactic Survey (JADES). Off-center clumps are detected in the rest-frame near-ultraviolet (NUV) using similar techniques to those in earlier studies based on Hubble Space Telescope (HST) images. This is done for a sample of 9,121 star-forming galaxies with stellar masses $\log\left(M_{\star}/M_{\odot}\right)\geq 8$. The fraction of clumpy galaxies, $f_{\rm{clumpy}}$, increases from $\sim 20\%$ at $z>6.5$ to $\sim 80\%$ at $z\sim 2.75$ at $\log\left(M_{\star}/M_{\odot}\right)\geq 9$. Our $f_{\rm{clumpy}}$ values are generally higher at fixed redshift and increase faster with decreasing redshift than what studies based on HST data found, which we attribute largely to the higher sensitivity of JWST. Clumps make up $15-20\%$ of the NUV luminosity of clumpy galaxies, with little dependence on redshift or stellar mass. $f_{\rm{clumpy}}$ correlates with stellar mass and there is little correlation with specific star-formation rate. Our $f_{\rm{clumpy}}$ measurements are compared with those from simulations as well as other observations. At low redshifts ($z\lesssim 5.75$) and intermediate-to-high stellar masses ($\log\left(M_{\star}/M_{\odot}\right)\geq 9$), our results suggest gas fragmentation due to violent disk instabilities to be the dominant mechanism for forming clumps. At high redshifts and low stellar masses, compression of gas during mergers appears to dominate.
Galaxy evolution(594) — Galaxy structure(622) — High-redshift galaxies(734) — Star forming regions(1565)
software: astropy (Astropy Collaboration et al., 2013, 2018, 2022), matplotlib (Hunter, 2007), numpy (Harris et al., 2020), photutils (Bradley et al., 2020; Bradley, 2023), scipy (Virtanen et al., 2020)
1 Introduction
Early Hubble Space Telescope (HST) imaging of high-redshift galaxies revealed systems with patchy, clumpy, kpc-scale substructures (Cowie et al., 1995; Giavalisco et al., 1996; van den Bergh et al., 1996; Elmegreen & Elmegreen, 2005; Elmegreen et al., 2007). Later work showed that these substructures (hereafter, “clumps”) are bright in the rest-frame ultraviolet (UV), are fairly massive ($\gtrsim 10^{7}~M_{\odot}$), have young stellar ages ($\sim 100$ Myr), and make up about $10-50\%$ of the total star-formation rates (SFRs) and about $4-20\%$ of the total stellar mass of their host galaxies, on average (Elmegreen et al., 2008, 2009; Förster Schreiber et al., 2011; Guo et al., 2012, 2015, 2018; Wisnioski et al., 2012; Wuyts et al., 2012, 2013; Livermore et al., 2015; Shibuya et al., 2016; Soto et al., 2017; Zanella et al., 2019; Mehta et al., 2021; Vanzella et al., 2021; Meštrić et al., 2022; Claeyssens et al., 2025). These results suggest that clumps were significant sites of star formation in galaxies over $0.5<z<3$, during the peak of cosmic star-formation activity (Madau & Dickinson, 2014).
Studies of observed and simulated clumps have proposed two main formation mechanisms. The first is violent disk instabilities (VDI), wherein gas collapses under gravitational instabilities in marginally stable gaseous disks (Noguchi, 1998, 1999; Immeli et al., 2004a, b; Bournaud et al., 2007, 2009; Agertz et al., 2009; Dekel et al., 2009a; Ceverino et al., 2010; Romeo et al., 2010; Romeo & Agertz, 2014; Inoue et al., 2016). The VDI is sustained by the high gas fractions in high-redshift galaxies (Tacconi et al., 2020, and references therein), which are fed by the accretion of cold gas from the intergalactic medium (Birnboim & Dekel, 2003; Kereš et al., 2005, 2009; Dekel & Birnboim, 2006; Bournaud & Elmegreen, 2009; Dekel et al., 2009b, a; Nelson et al., 2013). Observational support for the VDI has been found by examining integral field spectroscopic data, with which maps of gas kinematics may be measured (Förster Schreiber et al., 2011; Genzel et al., 2011, 2014; Wisnioski et al., 2011, 2012; Newman et al., 2012; Mieda et al., 2016). The second mechanism is mergers, during which gas is compressed by tidal forces and/or turbulence due to gravitational interactions between galaxies (Di Matteo et al., 2008; Renaud et al., 2015; Nakazato et al., 2024). Observational support for mergers has come from visual inspection of the morphologies of clumpy and non-clumpy galaxies (Straughn et al., 2015) and integral field spectroscopy (Puech, 2010; Menéndez-Delmestre et al., 2013).
Clumps could also evolve within their host galaxies and alter their morphologies. Theoretical studies and simulations have predicted that clumps experience dynamical friction and/or interactions, which cause them to migrate to the centers of their hosts and contribute to the growth of bulges (Noguchi, 1999; Immeli et al., 2004a, b; Dekel et al., 2009a; Krumholz & Dekel, 2010; Ceverino et al., 2012; Inoue & Saitoh, 2012; Perez et al., 2013; Bournaud et al., 2014; Mandelker et al., 2014, 2017; Dekel et al., 2022). However, other simulations find that clumps are short-lived and dissipate within $\sim 100$ Myr due to stellar feedback, which drives strong outflows of gas (Murray et al. 2010; Genel et al. 2012; Hopkins et al. 2012; Moody et al. 2014; Buck et al. 2017; Oklopčić et al. 2017; Meng & Gnedin 2020; see also Fensch & Bournaud 2021). In this scenario, stars formed within the clumps could contribute to the growth of the thick disk (Noguchi 1996; Inoue & Saitoh 2014; see also Bournaud et al. 2007).
The evolution of the fraction of clumpy galaxies ($f_{\rm{clumpy}}$) with redshift may also provide clues to the physical processes through which clumps form and evolve (Guo et al., 2015; Shibuya et al., 2016). Using HST imaging data, $f_{\rm{clumpy}}$ has been estimated at various redshifts by identifying compact, UV-bright regions within galaxies. The behavior of $f_{\rm{clumpy}}$ with redshift is well understood at $z\leq 3$, but unclear at higher redshifts. At $z\leq 3$, $f_{\rm{clumpy}}$ peaks at $\sim 60\%$ at $z\sim 2$ and declines at lower and higher redshifts (Elmegreen et al., 2007; Overzier et al., 2009; Puech, 2010; Guo et al., 2012; Murata et al., 2014; Guo et al., 2015; Hinojosa-Goñi et al., 2016; Shibuya et al., 2016; Adams et al., 2022; Sok et al., 2022; Martin et al., 2023; Sattari et al., 2023). These studies are largely consistent with predictions from simulations of the redshift evolution of galaxies undergoing VDI, based on simulations (Cacciato et al., 2012). Support for mergers was also found at lower redshift (Puech, 2010) and lower masses (Guo et al., 2015). At $z>3$, there is disagreement on the behavior of $f_{\rm{clumpy}}$ with redshift. One group of studies found that $f_{\rm{clumpy}}$ decreases from $\sim 40\%$ at $z\sim 3$ to $\sim 20\%$ at $z\sim 8$ (Ravindranath et al., 2006; Conselice & Arnold, 2009; Oesch et al., 2010; Shibuya et al., 2016). Another found that $f_{\rm{clumpy}}$ is roughly constant with redshift at $\sim 40\%$ from $z\sim 4$ to $z\sim 8$ (Jiang et al., 2013; Kawamata et al., 2015; Curtis-Lake et al., 2016; Bowler et al., 2017; Ribeiro et al., 2017).
Studies using data from the James Webb Space Telescope (JWST) have revealed a more complex picture of the clumps. Thanks to the higher angular resolution, greater sensitivity, and longer wavelength coverage of the JWST, sub-kpc substructures in faint galaxies at high redshift can now be studied in greater detail with its Near-Infrared Camera (NIRCam). JWST imaging of gravitationally lensed galaxies over $1<z<8.5$ revealed a significant population of clumps that were not detected by the HST (Claeyssens et al., 2023; Fujimoto et al., 2024). Un-lensed galaxies at $z>4$ also reveal clumps in JWST images (e.g., Tacchella et al., 2023; Hainline et al., 2024), some of which are not found in HST observations (e.g., Bik et al., 2024; Tanaka et al., 2024). On average, 76% of the rest-frame UV light from $z\simeq 6-8$ galaxies was found to come from the clumps (Chen et al., 2023), which is much higher than the $\sim 20-30\%$ measured over $0.5<z<3$ (Guo et al., 2015). The rest-optical and rest-NIR properties of clumps have been analyzed out to higher redshifts ($z\sim 2$) than what is possible with HST (Kalita et al., 2024, 2025; Mercier et al., 2025). Massive ($\log\left(M_{\star}/M_{\odot}\right)\geq 10$) star-forming galaxies at $z<3$ were found to have higher $f_{\rm{clumpy}}$ when undergoing starburst activity (Faisst et al., 2025). Finally, about 70% of spectroscopically selected UV-bright galaxies at $z\sim 7$ contain clumps (Harikane et al., 2025), in contrast with the 102010-20% inferred from pre-JWST studies (Shibuya et al., 2016, and references therein).
While recent studies show that clumpy galaxies might be more abundant than expected at high redshifts, interpretation of these results is difficult due to selection effects and occasional small sample sizes. The largest post-JWST sample at $z>6$ to date for which $f_{\rm{clumpy}}$ has been measured consists of 53 galaxies (Harikane et al., 2025). That sample was selected to be intrinsically bright with absolute UV magnitudes $23<M_{UV}<-19$ mag. Tan et al. (2024) measured $f_{\rm{clumpy}}$ over $0.25<z<5$ for a sample of 909 galaxies expected to be progenitors of Milky Way analogs. Their sample was selected with bins in stellar mass that evolve with redshift. This complicates comparisons with pre-JWST studies, which largely selected galaxies with a mass range that was constant with redshift. Claeyssens et al. (2025) examined the physical properties of 1,751 clumps in 377 galaxies over $0.7<z<5.5$ that are gravitationally lensed by the cluster Abell 2744. A measurement of $f_{\rm{clumpy}}$ as a function of redshift was not performed in that work.
In this paper, we measure $f_{\rm{clumpy}}$ over $2<z<12$ with JWST data for a sample of 9,121 star-forming galaxies. The selection of the sample is performed such that results from the JWST and HST data may be directly compared. This allows us to determine whether the redshift evolution of $f_{\rm{clumpy}}$ is significantly different with JWST data compared with HST data.
The paper is organized as follows. In Section 2, we describe the data used in this work. The selection criteria of our sample are given in Section 3. Section 4.1 describes the completeness of the routine used to identify clumps. The redshift evolution of $f_{\rm{clumpy}}$ is discussed in Section 5. The contribution of clumps to the rest-frame NUV luminosity is shown in Section 6. In Section 7, we examine the correlation between stellar mass and sSFR with $f_{\rm{clumpy}}$. Predictions from simulations of clumpy galaxies are compared with our observations in Section 8. We discuss the implications for clump formation mechanisms in Section 9. We summarize and conclude in Section 10.
Throughout this work, cosmological parameters measured by Planck Collaboration et al. (2016) are adopted, with $H_{0}=67.7~\rm{km}~\rm{s}^{-1}~\rm{kpc}^{-1},\Omega_{\rm{m}}=0.309$, and $\Omega_{\Lambda}=0.691$. All magnitudes are in the AB photometric system (Oke & Gunn, 1983).
2 Data
2.1 Images, Source Detection, and Photometry
[…]
Rieke et al. (2023b) detected sources in JADES and performed photometry in nine NIRCam and 14 ground- and space-based bandpasses. Briefly, a detection image was created by stacking the F277W, F335M, F356W, F410M, and F [^4] W images and weighting them by their inverse variances. Next, various source detection routines in the python module Photutils (Bradley, 2023) were used to detect and deblend sources and perform photometry. All images are aligned to the same reference frame and drizzled to a common pixel scale of 0.″03  per pixel. Source detection and photometry are performed on HST and JWST images convolved to the F [^4] W resolution. For more details, see Section 4 in Rieke et al. (2023b).
This work relies on photometry in HST and JWST bandpasses. Spatially integrated fluxes are measured in Kron apertures with a Kron parameter of $k=2.5$. Spatially integrated errors on the HST and JWST photometry are computed as follows. Two sources of error are considered: errors due to sky subtraction and a combination of source Poisson and instrumental errors. The former are measured in apertures of various sizes that are centered on empty regions of the sky (see Section 4.2 in Rieke et al. 2023b, also Labbé et al. 2005; Whitaker et al. 2011; and Skelton et al. 2014). For JWST/NIRCam images, the latter come from the ERR extension, and for HST images, they are measured following the procedure described in Sec. 4.1 by van der Wel et al. (2012). Total errors on the JWST photometry are computed by adding these two sources of error in quadrature. For the HST images, the JADES team have measured errors due to sky subtraction and report the average values of the weight, or inverse variance, within the Kron apertures. The error due to the weight for each galaxy is computed by multiplying the average weight value by the area of the Kron aperture in pixels. Total errors on the HST photometry are calculated by adding these two errors in quadrature. A final error is added to all fluxes for each bandpass to account for zeropoint errors and mismatches in the templates between model and observed SEDs. It is equal to 5% of the flux (see, e.g., Table 4 of Dahlen et al. 2013 and Table 11 of Skelton et al. 2014).
Error maps in the JWST bandpasses, which are needed to detect clumps, come from the ERR extension in the images provided by the JADES team. Those in the HST bandpasses are measured following the procedure described in Sec. 4.1 in van der Wel et al. (2012).
Axis ratios, half-light radii, segmentation maps, and central coordinates of galaxies were measured from the detection images by Rieke et al. (2023b). These measurements and data products are adopted for this analysis. They were calculated by applying photutils to the detection image in each field. Half-light radii were measured in circular apertures. Central coordinates were computed by taking the light-weighted centroids along the horizontal and vertical axes of the detection images.
2.2 Photometric and Spectroscopic Redshifts
Photometric redshifts in JADES come from Hainline et al. (2024) and are estimated using EAZY (Brammer et al., 2008). EAZY measures photometric redshifts by fitting non-negative linear combinations of templates to observed fluxes in order to derive probability distribution functions of the redshift. Templates are supplied by the user. Hainline et al. (2024) develop modified versions of the original galaxy templates used by Brammer et al. (2008) and seven new templates designed to fit the spectral energy distributions (SEDs) of high-redshift galaxies. For more details, see Section 3 by Hainline et al. (2024).
Spectroscopic redshifts come from the JADES and the compilation over the CANDELS fields by Kodra et al. (2023). Descriptions of the selection of the targets, reduction and calibration of the data, measurements of the redshifts, and presentation of the JADES spectra are provided by Bunker et al. (2024) and D’Eugenio et al. (2024). The provenances of the spectroscopic redshifts gathered by Kodra et al. (2023) and descriptions of the quality cuts applied can be found in that work and references therein.
Figure 1: Comparison of 5,780 spectroscopic and photometric redshifts in the JADES fields. The former are compiled from various sources (see Section 2.2) and the latter are measured using EAZY (Brammer et al., 2008). The bias, normalized median absolute deviation (NMAD) scatter, and outlier fraction are measured according to Dahlen et al. (2013).
The accuracy of the photometric redshifts measured by the JADES team is assessed by comparing them with spectroscopic redshifts. We find 2,774 galaxies with spectroscopic redshifts in GOODS-N and 3,006 in GOODS-S. Their redshifts span $0<z<13.35$. The accuracies of the photometric redshifts are quantified using the bias, normalized median absolute deviation (NMAD) scatter, and outlier fraction according to Dahlen et al. (2013). The bias is given by
$\textrm{bias}=\textrm{mean}\left(\Delta z/(1+z_{\rm{spec}})\right),$ (1)
where $\Delta z=z_{\rm{spec}}-z_{\rm{phot}}$. It is computed after excluding outliers, which are defined below. The NMAD scatter is:
$\textrm{NMAD}=1.48\times\textrm{median}\left(|\Delta z|/(1+z_{\rm{spec}})\right).$ (2)
The outlier fraction is given by the fraction of galaxies that satisfy $|\Delta z|/\left(1+z_{\rm{spec}})\right)>0.15$.
Spectroscopic and photometric redshifts are compared in Figure 1. We measure a bias of 0.002, an NMAD scatter of 0.034, and an outlier fraction of 12.4% for all 5,780 redshifts from both fields. There is good agreement between the spectroscopic and photometric redshifts, given the low bias and NMAD scatter. The outlier fraction is somewhat high, which may be due to the shallower depth of and fewer medium bands available in GOODS-N than those of GOODS-S. Similar NMAD scatter and outlier fractions were obtained by Puskás et al. (2025). They built a catalog of spectroscopic redshifts in GOODS-N and GOODS-S and classified them in terms of their quality. Of these, 5,382 in GOODS-S were found to have the highest quality, for which they measured an NMAD scatter of 0.065 and outlier fraction of 6.91%. For the 2,591 highest-quality redshifts in GOODS-N, an outlier fraction of 13.62% was measured.
Figure 2: The relation between stellar mass and SFR, also known as the star-formation main sequence (SFMS), for the full sample (Section 3.1) over $2<z<12$. These quantities are measured from fits to the combined HST and JWST SEDs using bagpipes (Carnall et al., 2018, see Section 2.3). Measurements for individual galaxies are shown as gray dots in various redshift bins, a different one of which is in each column. Completeness limits in stellar mass are indicated with vertical dashed lines. There is broad agreement between the SFMS presented here and that measured in other studies. The SFMS measured by Popesso et al. (2023) over $2<z<6$ from a compilation of studies is shown as a blue solid line in the leftmost four columns. That measured by Salmon et al. (2015) over $3.5<z<6.5$ is plotted in the second through fourth columns. The SFMS from Cole et al. (2025), measured over $4.5<z<12$, is shown as a green dashed line. The SFMS measurements by Ciesla et al. (2024) over $6<z<12$ are shown as magenta dash-dotted lines.
2.3 Stellar Masses and Star-Formation Rates
Stellar masses and star-formation rates for all galaxies in the full sample (Section 3) are measured by fitting their SEDs. The SFRs used in this work are averaged over the last 100 Myr. Where available, fits were performed to the HST/Advanced Camera for Surveys (ACS) bandpasses F435W, F606W, and F775W combined with all the NIRCam bandpasses listed in Section 2.1. The number of filters used in each fit ranges from five to 17, with a median of 13. The Bayesian SED-fitting tool bagpipes (Carnall et al., 2018) is used. It fits SEDs using the MultiNest Markov Chain Monte Carlo algorithm (Feroz et al., 2009) and returns marginalized posterior probability distribution functions for each parameter of interest. Stellar masses and SFRs are measured by taking the median stellar mass and SFR from their respective posteriors.
We adopt the priors listed in Table 1 in Carnall et al. (2020), adjusting those on the metallicity and the amount of dust attenuation in the rest-frame $V$-band, $A_{V}$. This is done based on arguments for physically motivated priors given by de la Vega et al. (2025). A Gaussian prior on the metallicity $Z/Z_{\odot}$ with a mean $\mu=0.3$ and standard deviation $\sigma=0.5$ is assumed. Another Gaussian prior on $A_{V}$ is assumed, with $\mu=0.3$ and $\sigma=1.0$, over the range $A_{V}\in[0,6]$. This is adopted from Leja et al. (2019).
Priors for other physical parameters are as follows. A uniform prior on the log of the stellar mass $\log\left(M_{\star}/M_{\odot}\right)$ is assumed over the range of 1 to 13. We adopt the flexible attenuation curve model by Salim et al. (2018). It modifies the shape of the attenuation curve measured by Calzetti et al. (2000) by multiplying it by a power law that varies with wavelength. For the slope of the power law, $\delta$, a Gaussian prior is set with $\mu=0$ and $\sigma=0.1$ over the range [-0.3, 0.3]. The Calzetti et al. (2000) curve is obtained by setting $\delta=0$. The strength of the dust absorption feature at rest-frame 2,175 Å is also modeled by Salim et al. (2018). We set a uniform prior on it over the range [0, 5]. Another prior is assumed for the differential attenuation between young ($\leq 10$ Myr) and old stars (see, e.g., Calzetti et al., 2000; Charlot & Fall, 2000). The former are generally dustier than the latter due to additional attenuation from birth clouds. A uniform prior on $\eta$, the ratio of $A_{V}$ for young stars to that for old stars, is assumed over the range [1, 5]. SED fits include nebular emission, the strength of which is determined with the log of the ionization parameter, $\log U$, and the nebular metallicity. A uniform prior is set on the former over the range [4,2-4,-2]. The latter is assumed to be the same as that of the stars.
Last, a double power-law star-formation history is assumed. It is given by
$\textrm{SFR}(t)\propto\left[\left(\frac{t}{\tau}\right)^{\alpha}+\left(\frac{t}{\tau}\right)^{-\beta}\right]^{-1},$ (3)
where $t$ is the time since star-formation began in a galaxy, $\tau$ is the time after the Big Bang when the SFR peaked, $\alpha$ is the declining slope, and $\beta$ is the rising slope. We assume log-uniform priors for $\alpha$ and $\beta$, each of which is over the range [0.01, 1000]. A uniform prior on $\tau$ is set, which ranges from 0.1 Gyr to the age of the Universe at the observed redshift of a given galaxy.
For the SED fits, we fix the redshift to its spectroscopic value, if available. Otherwise, the photometric redshift from JADES is assumed. We apply upper limits on photometry with signal-to-noise ratio (S/N) <3<3 following the procedure by de la Vega et al. (2025). This is relevant only for the HST and bluest NIRCam filters, especially at high redshift.
Stellar population synthesis templates from version 2.2.1 of the Binary Population and Spectral Synthesis (BPASS; Eldridge et al., 2017; Stanway & Eldridge, 2018) models are used. The initial mass function is set to the default ‘135_300’ model listed in Table 2 in Stanway & Eldridge (2018). The cloudy (Ferland et al., 2017) photoionization code is used to model nebular emission. All HST and JWST filter transmission curves, which are used in the SED fits, are taken from the SVO Filter Profile Service (Rodrigo et al., 2012; Rodrigo & Solano, 2020).
In Figure 2, the relation between stellar mass and SFR, also known as the star formation main sequence (SFMS; e.g., Noeske et al., 2007; Rodighiero et al., 2010; Whitaker et al., 2014; Salmon et al., 2015; Schreiber et al., 2015; Tomczak et al., 2016; Popesso et al., 2023; Ciesla et al., 2024; Cole et al., 2025) is shown. These properties are derived from the SED fits. The distribution of stellar masses and SFRs measured in this work agrees broadly with the SFMS measured in various studies.
2.4 Stellar Mass Completeness Limits
Here we compute 90% mass completeness limit for our selected sample. The limits are based on the method in Pozzetti et al. (2010) and follow what was done for galaxies in JADES by Simmonds et al. (2024). The stellar masses obtained with bagpipes are used.
The completness limits are estimated as follows. First, a limiting stellar mass, $M_{\rm{lim}}$, is computed for all galaxies in the full sample (Section 3). It is defined to be
$\log\left(M_{\rm{lim}}\right)=\log\left(M_{\star}\right)+0.4\left(m-m_{\rm{lim}}\right),$ (4)
where $\log\left(M_{\star}\right)$ is the stellar mass in solar units of a given galaxy, $m$ is its apparent magnitude, and $m_{\rm{lim}}$ is the limiting magnitude of the survey. In this work, the magnitude in the F [^4] W band is used to calculate $M_{\rm{lim}}$. This band is used as it is the longest NIRCam filter available in both GOODS-N and GOODS-S and thus best traces the stellar mass. After finding $M_{\rm{lim}}$ for all galaxies, they are divided into redshift bins (see Section 3.1). In each bin, the faintest 20 percent of galaxies are identified. Then, the 90th percentile of $M_{\rm{lim}}$ for these is found. We take $m_{\rm{lim}}$ to be 29.8, which is the average $5\sigma$ point-source depth across the GOODS-N and GOODS-S fields, as measured by Puskás et al. (2025).
Completeness limits span $\log\left(M_{\star}/M_{\odot}\right)=7.16-8.21$. They change monotonically with redshift. These limits are comparable to those derived by other studies where there is overlap in redshift (e.g., Ciesla et al., 2024; Simmonds et al., 2024; Puskás et al., 2025).
[…]
3 Sample Selection
3.1 The Full Sample
We start with all sources in the JADES GOODS-N and JADES GOODS-S photometric catalogs. There are 179,709 sources in both fields: 85,709 in GOODS-N and 94,000 in GOODS-S.
All sources that are brighter than or equal to 29th magnitude in the reddest broadband filter, F [^4] W, are selected. This is about 0.5 mag brighter than the limiting magnitudes for all NIRCam filters in JADES (Table 2 in Rieke et al., 2023b). This cut removes 82,606 objects with 97,103 remaining. Next, we remove point sources using the FLAG_ST flag. This cut removes 531 sources, leaving 96,572 galaxies. All sources that are contaminated by diffraction spikes from bright stars are removed with the FLAG_BS flag. Those that are contaminated by bright neighbors are removed using the FLAG_BN flag. The former cut removes 1,504 galaxies and the latter removes 23,672, with 71,396 left.
Galaxies over $2<z<12$ are selected. This is done to enable the identification of clumps in the rest-NUV at all redshifts, following pre-JWST studies (e.g., Wuyts et al., 2012; Guo et al., 2015; Shibuya et al., 2016), over the widest possible range in redshift. The lower limit of $z=2$ is chosen because the bluest NIRCam bandpass available, F090W, spans the rest-frame NUV starting at $z=2$. The upper limit of $z=12$ is chosen such that the reddest bandpasses sample the rest-frame Balmer break. The range $2<z<12$ thus spans the rest-UV to rest-optical at all redshifts for the NIRCam bandpasses in JADES. This rest-frame wavelength coverage is optimal for the measurement of stellar masses and SFRs (Pacifici et al., 2012; Pforr et al., 2012; Buat et al., 2014). Applying the redshift cut removes 38,947 galaxies and leaves 32,449.
[…]
Cuts are then applied on the morphologies of galaxies. These are done to exclude compact and edge-on galaxies and have often been applied in pre-JWST studies of clumpy galaxies (e.g., Guo et al., 2012, 2015; Shibuya et al., 2016). Galaxies with circular half-light radii of at least 0.″1 in the F [^4] W bandpass are selected. This removes 2,848 galaxies with 16,345 remaining. Galaxies with axis ratios of $b/a\geq 0.5$ are chosen. This excludes 933 more.
The full sample contains 15,412 galaxies: 7,167 in GOODS-N and 8,245 in GOODS-S. Galaxies in the final sample have a median S/N of about 4.0 in F090W and 13.2 in F277W. For each galaxy in the full sample, its stellar mass and SFR are measured by fitting its combined HST+JWST SED (Section 2.3). About 8% (1,195/15,412) of the full sample has spectroscopic redshifts. The distribution of the full sample in terms of redshift and stellar mass is shown in Figure 4. Individual galaxies are represented with small gray dots. The stellar mass completeness limit for each redshift bin is indicated with a light gray shaded region. Galaxies with spectroscopic redshifts are indicated with empty red squares.
3.2 The Final Sample
Galaxies with stellar masses $\geq 10^{8}~M_{\odot}$ are selected. This is based on our completeness limits, details which are provided in Section 2.4. In the highest redshift bin, $9<z<12$, this cut in mass is about 80% complete. However, we choose to apply it across all redshift bins for consistency. Applying the mass cut removes 6,074 galaxies and leaves 9,338.
Figure 4: The stellar masses as a function of redshift are shown for the full sample (Section 3). Individual galaxies with only photometric redshifts are plotted as small gray dots and those with spectroscopic redshifts are indicated with red squares. Top panel: the distribution of redshifts. Right panel: stellar mass distribution. In the top and right panels, the distributions for galaxies with spectroscopic redshifts are denoted with red dash-dot histograms. Clumps are identified in galaxies with stellar masses of at least $10^{8}~M_{\odot}$, which is indicated with the horizontal dashed line. It is higher than or close to the 90% stellar mass completeness limit of the sample at all redshifts, which are shown as light gray shaded regions and capped with dotted lines.
Finally, galaxies are selected to be star-forming. This is done following pre-JWST studies, which studied star-forming clumpy galaxies (e.g., Guo et al., 2015; Shibuya et al., 2016; Huertas-Company et al., 2020; Sattari et al., 2023). In this work, galaxies with sSFR $\geq 10^{-9.5}~\textrm{yr}^{-1}$ are defined to be star-forming. Cutting on sSFR removes 217 galaxies. Our final sample contains 9,121 star-forming galaxies.
Figure 5: The completeness of our clump detection algorithm in redshift bins (columns) and stellar mass bins of their host galaxies (thickness of the lines) is shown here. Fake clumps are added to rest-frame NUV images of real galaxies. More than 1,500 galaxies are randomly selected over redshifts $2<z<12$ and stellar masses $8<\log\left(M_{\star}/M_{\odot}\right)<11$, regardless whether they have clumps. The completeness is defined as the fraction of fake clumps recovered relative to the number inserted into each galaxy image. The relation between the ratio of fake clump luminosity to host galaxy luminosity and completeness is used to determine minimum fractional luminosities used to find clumps. A completeness of 50% is denoted with the horizontal dashed lines. Minimum fractional luminosities that meet this completeness are indicated with vertical dash-dot lines. Those for more massive galaxies are denoted with thicker lines. Completeness curves are shown only for ranges in mass that include at least 10 galaxies. Faint clumps are more likely to be detected in higher-mass galaxies at lower redshifts. For galaxies over $9<\log\left(M_{\star}/M_{\odot}\right)<10$, clumps are identified with 50% completeness at 2% of the host luminosity at $z\sim 2.75$ and at 7% of the host luminosity at $z\sim 7.75$. At $z\sim 2.75$ and $\log\left(M_{\star}/M_{\odot}\right)\geq 9$, clumps can be identified at nearly four times fainter luminosities with JWST than with HST. The error on the completeness is calculated using Bayesian binomial confidence intervals and shown as the light shading around each line.
Table 1: Minimum Fractional Luminosities Needed to Identify $\geq 50\%$ of Clumps as a Function of Stellar Mass and Redshift
Redshift Range $8<\log\left(M_{\star}/M_{\odot}\right)<9$ $9<\log\left(M_{\star}/M_{\odot}\right)<10$ $10<\log\left(M_{\star}/M_{\odot}\right)<11$
($L_{\rm{clump}}/L_{\rm{host~galaxy}}$) ($L_{\rm{clump}}/L_{\rm{host~galaxy}}$) ($L_{\rm{clump}}/L_{\rm{host~galaxy}}$)
2–3.5 $\geq$7% $\geq$2% $\geq$2%
3.5–5 $\geq$10% $\geq$2% $\geq$2%
5–6.5 $\geq$15% $\geq$5% $\ldots$
6.5–9 $\geq$25% $\geq$7% $\ldots$
9–12 $\ldots$ $\ldots$ $\ldots$
Note. — Each fraction is in terms of the rest-frame NUV luminosity of the host galaxy. Empty values indicate that either too few (<10<10) galaxies lie in a given mass and redshift bin or clumps could not be detected with at least 50% completeness.
4 Selection of Clumpy Galaxies
4.1 Identification of Clumps
Clumps are identified for all galaxies in the mass-selected, star-forming sample (Section 3.2) using the technique developed by Sattari et al. (2023). Briefly, it involves taking the Fourier Transform of the rest-frame NUV image of a galaxy, filtering out the low-frequency signal that corresponds to the host galaxy light, and performing source detection on the leftover high-frequency signal, which consists of clumps and noise (see Sec. 3 in Sattari et al. 2023 for details). To distinguish these, we apply stringent source detection criteria.
Objects are identified as clumps in JWST/NIRCam images if they span at least five contiguous pixels, each of which must be at least two standard deviations brighter than the background-subtracted image, and if the total S/N of the object is at least 3. These criteria are the same as those used by Shibuya et al. (2016), who used HST data, and similar to those used by Claeyssens et al. (2025), who used JWST data. We use photutils to detect clumps. For a given galaxy, clumps must lie at least 0.″1 from the center and within the segmentation map created by the JADES team. This corresponds to a distance of 0.86 kpc at $z=2$ and 0.37 kpc at $z=12$. The former criterion is used to avoid confusion with a bulge or active galactic nucleus and the latter ensures that all detected clumps lie within a single galaxy and not within its neighbors projected on the sky.
Figure 6: Example clumpy galaxies are shown in this figure for each of the stellar mass and redshift bins considered in this work (Figure 5 and Section 4.2). Stellar mass bins are indicated with the columns and increase from left to right. Redshift bins are denoted with the rows and increase from top to bottom. For each galaxy in a given mass and redshift bin, images in the rest-frame NUV are displayed in the left two columns and a RGB image is shown in the rightmost column. Images are shown with an arcsinh stretch. In the middle columns, detected clumps are circled in pink. Clumps that meet the completeness limits listed in Table 1 are denoted with solid lines and those that do not with dotted lines. The segmentation map of each galaxy is indicated with the red outline. The bandpasses used to create the RGB images are listed on the far left side of the figure.
4.2 Completeness of Clump Identification
To ensure that a complete sample of clumps is identified, we measure how faint clumps can be in order to detect at least half of them, following pre-JWST studies (Guo et al., 2015; Shibuya et al., 2016; Sattari et al., 2023). This is done by adding fake clumps to images of randomly selected real galaxies, regardless whether they have clumps, and running the clump identification method described above.
Fake clumps are modeled as point sources following previous works (e.g., Guo et al., 2012, 2015, 2018; Shibuya et al., 2016; Soto et al., 2017; Sattari et al., 2023). We measure empirical point spread functions (PSFs) for all NIRCam bandpasses in GOODS-N and GOODS-S. This is described in Appendix A. Positions of the fake clumps are randomly selected to lie within the segmentation map of a given galaxy. Their brightnesses vary randomly between 0.1% and 75% of that of the host galaxy. In order to have sufficient statistics to determine the completeness, 100 galaxies are randomly selected in each of three stellar mass bins ($8<\log\left(M_{\star}/M_{\odot}\right)<9,9<\log\left(M_{\star}/M_{\odot}\right)<10$, and $10<\log\left(M_{\star}/M_{\odot}\right)<11$) for each redshift bin. If fewer than 100 galaxies exist within a given bin in mass and redshift, all galaxies within that bin are selected. In total, 1,541 galaxies are randomly selected. A fake clump is added to the rest-frame NUV image of each galaxy using the ImagePSF routine in photutils. The clump identification algorithm is then run on the modified image to determine whether the fake clump is recovered. For each randomly selected galaxy, this procedure is repeated 50 times. The position and brightness of the fake clumps are randomly varied in each iteration. Uncertainties on the completeness are computed in a Bayesian fashion and assume a noninformative Jeffreys prior on the binomial distribution (Cameron, 2011).
The completeness of the fake clump simulations is shown in Figure 5. In each panel, the fraction of clumps detected by the clump identification algorithm is shown as a function of the fake clump luminosity. This is further shown as a function of stellar mass (lines of various thicknesses) and redshift (columns). Clumps are more likely to be detected in higher-mass systems and lower redshifts. The minimum fractional luminosities at which clumps are detected at 50% completeness are indicated with the vertical dashed-dotted lines in Figure 5. For galaxies at $z\sim 2.75$ and over $9<\log\left(M_{\star}/M_{\odot}\right)<10$, 50% of clumps are detected when they are at least 2% of the NUV luminosities of their hosts. At $z\sim 7.75$ and over the same range in mass, the same completeness is obtained for a minimum fractional luminosity of 7%. At lower masses, minimum fractional luminosities are higher, and at higher masses, they are similar. At $z\sim 10.5$, no clumps are detected with at least 50% completeness. Because of this, galaxies in the two highest redshift bins at $z\sim 7.75$ and $z\sim 10.5$ will be combined into one redshift bin in the following Sections. The minimum fractional luminosities at $z\sim 7.75$ will be used to find clumps over $6.5<z<12$.
Table 1 lists the minimum fractional luminosities as a function of stellar mass and redshift. At $z\sim 2.75$ and $\log\left(M_{\star}/M_{\odot}\right)>9$, clump identification is 50% complete at four times fainter luminosities with JWST (2%) than with HST (8%; Shibuya et al. 2016).
A galaxy is defined to be clumpy if it has at least one off-center clump with a rest-frame NUV luminosity that is greater than or equal to the completeness limits listed in Table 1. These depend on stellar mass and redshift. Example clumpy galaxies are shown in Figure 6. Clumps that are brighter than the completeness limits are indicated with solid circles and those that are fainter with dotted circles.
Figure 7: The fraction of clumpy galaxies, $f_{\rm{clumpy}}$, over $2<z<12$. It peaks at $z\sim 2.75$ and declines by a factor of $\sim 4$ from $z\sim 2.75$ to $z>6.5$ at a nearly constant rate. Galaxies are considered clumpy if they have at least one off-center clump. Measurements from this work for galaxies with stellar masses of at least $10^{9}~M_{\odot}$ are shown as black hexagons connected by a solid black line. $f_{\rm{clumpy}}$ measurements for galaxies in GOODS-N are shown to the left of the solid black line. Those for galaxies in GOODS-S are to the right of it. The minimum fractional luminosity changes with redshift according to Table 1. We compile measurements from the literature and plot them alongside ours. These are shown as various symbols and lines and are listed in Table 2 (see Section 5 for details). The results are consistent: $f_{\rm{clumpy}}$ rises gradually from $z\sim 8$ to $z\sim 1.5-2$, when it peaks, and declines at lower redshifts. This is qualitatively consistent with the evolution of the cosmic SFR density, which is shown as a thin blue solid line and has been arbitrarily scaled vertically. The results of this study are consistent with the view that clumpy galaxies became more common at later times. However, the $f_{\rm{clumpy}}$ calculated in this work is higher at fixed redshift and increases more steeply with decreasing redshift than the bulk of previous studies. This suggests that clumpy galaxies were more common at high redshift than previously measured, but with consistent trends towards lower redshifts.
4.3 Aperture Photometry of Clumps
We measure the fluxes of detected clumps by performing fixed aperture photometry with photutils. Each aperture is circular with a radius of 3 pixels. The background is measured in annuli with inner radii of 4 pixels and outer radii of 10 pixels. To perform background subtraction, the mean surface brightness of the light in the annuli is measured, multiplied by the area of the aperture, and subtracted. In each of these steps, for a given galaxy, any nearby clumps are masked using a circular mask of radius equal to 4 pixels. The sizes of the apertures, background annuli, and masks are adopted from Guo et al. (2018), who carried out extensive tests on photometry of clumps in HST images.
Because clumps are modeled as point sources, the circular apertures used here enclose the majority of the total light, about 50-70%, depending on the bandpass. We account for this by applying aperture corrections that vary with wavelength. Clump aperture fluxes are multiplied by factors of 1.434, 1.400, 1.422, 1.477, and 1.886 in GOODS-N and 1.471, 1.429, 1.469, 1.502, and 1.806 in GOODS-S for the F090W, F115W, F150W, F200W, and F277W bandpasses, respectively. These are determined using the empirical PSFs that are measured in this work (see Appendix A).
5 The Fraction of Clumpy Galaxies over the Redshift Range $2<z<12$
We now study the fraction of clumpy galaxies ($f_{\rm{clumpy}}$) as a function of redshift. For a galaxy to be considered clumpy, it must have at least one off-center clump identified in its rest-frame NUV image. This is done using the technique developed by Sattari et al. (2023). Clumps are identified using the minimum fractional luminosities listed in Table 1). The fraction of clumpy galaxies is defined to be the number of galaxies with clumps divided by the total number of galaxies in a given bin in stellar mass, redshift, and/or sSFR. All uncertainties are calculated assuming Bayesian confidence intervals on the binomial distribution, unless otherwise noted.
The evolution of $f_{\rm{clumpy}}$ with redshift is examined in Section 5.1. In Section 5.2, measurements from this work are compared with those from other studies that overlap with ours in redshift. We find that our results disagree with those from most studies over $2<z<6.5$. To understand why this is, the redshift evolution of $f_{\rm{clumpy}}$ is measured from HST and JWST images and examined in Section 5.3.
Table 2: Summary of Previous Studies on Clumpy Galaxies
Reference Sample $\left(N_{\rm{galaxy}}\right)$ Stellar Mass Range Redshift Range Method Detection Band
$\log\left(M_{\star}/M_{\odot}\right)$
Lotz et al. (2006) LBGsaafootnotemark: bbfootnotemark: (82) $\gtrsim 9.5$ $\sim 4$ $CAGM_{20}$ccfootnotemark: rest-frame UV
Ravindranath et al. (2006) SFGsddfootnotemark: (5719) $\gtrsim 9.6$ 0.95–4.9 Visual rest-frame UV
Elmegreen et al. (2007) Starbursts (1003) N/A 0–5 Visual F775W
Conselice & Arnold (2009) LBGs (133) $\gtrsim 9$ 4–6 $CASGM_{20}$eefootnotemark: rest-frame UV
Oesch et al. (2010) LBGs (21) $\gtrsim 8$ 7–8 Visual rest-frame UV
Law et al. (2012) SFGs (309) 9–11 1.5–3.6 $CAGM_{20}$ rest-frame UV
Wuyts et al. (2012) SFGs (649) >10>10 0.5–2.5 Algorithm rest-frame 2800 Å
Jiang et al. (2013) LBGs (24) $\gtrsim 9.2$ 5.7–7 $CAGM_{20}$ rest-frame UV
[…]
Guo et al. (2015) SFGs (3239) 9–11.5 0.5–3 Algorithm rest-frame NUV
Kawamata et al. (2015) LBGs (39) $\gtrsim 8$ 6–8 Visual rest-frame UV
Curtis-Lake et al. (2016) LBGs (1318) $\gtrsim 8.5$ 4–8 Asymmetry rest-frame UV
Shibuya et al. (2016) SFGs and LBGs (13787) 9–12 and $\gtrsim 8.5$ 0-8 Algorithm rest-frame UV and Optical
Bowler et al. (2017) LBGs (25) $\gtrsim 10$ $\simeq 7$ Visual rest-frame UV
Ribeiro et al. (2017) SFGs (1242) 9.35–11.5 2–6 Algorithm F814W
Huertas-Company et al. (2020) SFGs (1575) 9–12 1–3 Algorithm rest-frame UV and Optical
Sattari et al. (2023) SFGs (6767) >9.5>9.5 0.5–3 Algorithm rest-frame 1600 Å
Tan et al. (2024) MWAfffootnotemark: progenitors (909) 8.5–10.7 0.25–5 Algorithm rest-frame $U$ band
Harikane et al. (2025) SFGs (12) $\gtrsim 8.5$ 6.595–7.154 Algorithm F115W
This work SFGs (9121) $\geq 8$ 2–12 Algorithm rest-frame NUV
Note. — aLyman-break Galaxies. bStellar mass ranges for samples comprising Lyman-break galaxies are estimated using the relations between absolute UV magnitude and stellar mass shown in Fig. 1 in Shibuya et al. (2015). cC:C:”concentration; “A:A:”asymmetry; “G:G:” Gini coefficient; $M_{20}:$ second-order moment parameter. dStar-forming galaxies. eS:S:” clumpiness. fMilky Way Analog. See Conselice (2014) for definitions of the $CASGM_{20}$ morphological parameters.
5.1 Evolution with Redshift
Figure 7 shows $f_{\rm{clumpy}}$ as a function of redshift. Measurements from this work for galaxies with stellar masses greater than $10^{9}~M_{\odot}$ are shown as black hexagons connected by a thick solid line. This limit in mass is constant with redshift and chosen for two reasons. First, at a fixed redshift, clumps are identified using a uniform cut in luminosity at masses above this limit (see Table 1). Second, it is comparable to the mass limits used by other studies (see Table 2).
The fraction of clumpy galaxies is highest ($\sim 80\%$) at $z\sim 2.75$ and decreases with increasing redshift to $\sim 20$% at $z>6.5$. The fraction decreases at a roughly constant rate with increasing redshift: it is $\sim 60\%$ at $z\sim 4.25$ and $\sim 40\%$ at $z\sim 5.75$. $f_{\rm{clumpy}}$ measurements in either GOODS-N or GOODS-S agree with this trend and differ from it by $\leq 4\%$. These are indicated with empty hexagons to the left and right of the solid black line in Figure 7.
Our $f_{\rm{clumpy}}$ measurements are also compared with the cosmic SFR density (SFRD), as measured by Madau & Dickinson (2014). This is shown as a thin blue solid line in Figure 7. The fraction of clumpy galaxies is observed to rise with decreasing redshift more steeply than the SFRD. However, the net change in $f_{\rm{clumpy}}$ from $z>6.5$ to $z\sim 2.75$, an increase of about a factor of four, is similar to that in the SFRD over the same range in redshift. This seemingly suggests a connection between the frequency of clumpy galaxies and the volume-averaged SFR at high redshift.
Figure 8: Clumpy galaxies are found to be more common over $2<z<6$ in this work compared with earlier studies based on HST due to the higher sensitivity of JWST. This figure shows $f_{\rm{clumpy}}$ measured with HST (dashed lines) and JWST (solid lines) as a function of redshift for three choices of the minimum fractional luminosity $(L_{\rm{clump}}/L_{\rm{host~galaxy}})$. This is the minimum fraction of the host galaxy NUV luminosity a clump must have. A slight offset in redshift is applied to each set of $f_{\rm{clumpy}}$ measurements for clarity. In the left panel, the limit is set to 8%, which is adopted from pre-JWST studies. Measurements from both telescopes agree to within 9% and are statistically consistent. At the intermediate limit of 5%, shown in the middle panel, fractions differ by 6–18%, with those from JWST being higher at all redshifts. Estimates are statistically consistent for only one redshift bin. Finally, at the lowest limit of 2%, shown in the right panel, JWST-based measurements are higher than those from HST by 22–29% and statistically inconsistent at all redshifts.
5.2 Comparison with Other Studies
Our $f_{\rm{clumpy}}$ measurements are also compared with those from other studies in Figure 7. The samples and methods of these are summarized in Table 2. The uncertainties on $f_{\rm{clumpy}}$ were not published in some studies. They were calculated in this work assuming Bayesian binomial confidence intervals, wherever feasible. Details on the extraction of $f_{\rm{clumpy}}$ measurements from earlier studies are provided in Appendix B.
We now examine the evolution of $f_{\rm{clumpy}}$ over $0<z<8$ in the rest-frame UV from previous studies, shown in Figure 7. This was found to be $\sim 30-40\%$ at $z\sim 0.5$, peak at $\sim 60-80\%$ at $z\sim 2$, and decline to $\sim 40-60\%$ at $z\sim 3$ (Elmegreen et al. 2007; Wuyts et al. 2012; Tadaki et al. 2014; Guo et al. 2015; Shibuya et al. 2016; Sattari et al. 2023; although lower fractions of $\sim 20-40\%$ were obtained by Law et al. 2012). The decline continues to $\sim 10-30\%$ at $z\sim 4$ (Lotz et al. 2006; Ravindranath et al. 2006; Conselice & Arnold 2009; Shibuya et al. 2016; Ribeiro et al. 2017; although Elmegreen et al. 2007 and Curtis-Lake et al. 2016 measured higher fractions of $\sim 40-50\%$). At $z\gtrsim 4$, previous measurements of $f_{\rm{clumpy}}$ largely divide into two groups: lower fractions of $\sim 10-20\%$ that are approximately constant with redshift (Conselice & Arnold 2009; Oesch et al. 2010; Shibuya et al. 2016; and Kawamata et al. 2015, although their $f_{\rm{clumpy}}$ was measured to be $\sim 50$% at $z\sim 8$) and higher fractions of $\sim 30-70\%$ over $4<z<8$ (Elmegreen et al., 2007; Jiang et al., 2013; Curtis-Lake et al., 2016; Bowler et al., 2017; Ribeiro et al., 2017; Harikane et al., 2025). Another measurement of $f_{\rm{clumpy}}$ as a function of redshift was recently made by Tan et al. (2024). Galaxies were selected in bins of stellar mass that evolve with redshift to track the redshift evolution of progenitors of Milky Way-like systems. Their evolution of $f_{\rm{clumpy}}$ with redshift is consistent with the bulk of studies discussed here: $f_{\rm{clumpy}}$ was measured to be $\sim 20\%$ over $2<z<5$ and increased with redshift to $\sim 60\%$ at $z\sim 0.5$.
The results of this work over $2<z<6.5$ are inconsistent with those from most of the previous studies at similar redshifts. We find that $f_{\rm{clumpy}}$ rose quickly with redshift at early times, from $\sim 20\%$ at $z>6.5$ to $\sim 80\%$ at $z\sim 2.75$ at a roughly constant rate with redshift. In contrast, most studies found that $f_{\rm{clumpy}}$ changed by $\lesssim 10\%$ over $4<z<8$ and began to increase at $z\lesssim 4$.
Our measurements are consistent with those from Wuyts et al. (2012), Jiang et al. (2013), and Bowler et al. (2017) at the redshifts where they overlap. Our results are also consistent with those from Elmegreen et al. (2007) and Ribeiro et al. (2017) at $z\sim 4.25$ and those from Shibuya et al. (2016) at $z>6.5$.
5.3 Comparing Clumpy Galaxy Fractions between the HST and JWST Data
Our results on the redshift evolution of $f_{\rm{clumpy}}$ disagree with those of most studies over $2<z<6.5$. All but two of these are based on HST data, whereas this work is based on JWST data. To determine the cause of the disagreement, we compute $f_{\rm{clumpy}}$ using both HST and JWST images for a subset of our sample and vary the minimum fractional luminosity of the clumps. As we demonstrate in this Section, it is largely the difference in sensitivity between the two datasets that is responsible for the disagreement in the $f_{\rm{clumpy}}$ measurements shown here. Other factors may have secondary effects on the measurement of $f_{\rm{clumpy}}$, such as the differences between photometric redshifts, stellar masses, and SFRs from earlier studies and those used in this work.
Galaxies were selected for this comparison as follows. We match the mass-selected, star-forming sample (Section 3.2) with photometric catalogs from the CANDELS survey, which were created by Barro et al. (2019) for GOODS-N and Guo et al. (2013) for GOODS-S. We use a match radius of 0.″5 and match HST/Wide Field Camera 3 (WFC3) F160W magnitudes to within 0.5 mag. There are 2,007 matches in GOODS-N and 2,189 in GOODS-S. As studies based on HST data limited their analyses to galaxies with stellar masses $\gtrsim 10^{9}~M_{\odot}$ (see Table 2), we only consider those with masses $\geq 10^{9}~M_{\odot}$. This mass cut leaves a total of 1,245 galaxies (592 in GOODS-N and 653 in GOODS-S) that will be used for our comparison.
Clumps are identified in HST images using the algorithm by Sattari et al. (2023) and the detection criteria from Shibuya et al. (2016). These are the same algorithm and criteria as those used to search for clumps in the JWST images (Section 4.1). HST images have a pixel scale of 0.″06 pixel-1. We note that the criterion on the minimum number of contiguous pixels that must be spanned by a candidate clump depends on the HST instrument used: it is 5 for ACS images and 7 for WFC3 images. We search for clumps in HST data over $2<z<6.5$ to ensure that the coverage in rest-frame wavelength is as similar as possible between HST and JWST (see Section 3). Clumps are identified in HST bandpasses that span the rest-frame NUV at various redshifts. They are identified in the HST instruments/bandpasses as follows: ACS/F850LP over $2<z<3.5$; WFC3/F125W over $3.5<z<5$; and WFC3/F160W over $5<z<6.5$. Cutting on redshift removes 18 galaxies, with 1,227 remaining. We make sure that $f_{\rm{clumpy}}$ is measured using the same galaxies for both HST and JWST images. Therefore, any difference in $f_{\rm{clumpy}}$ is due to discrepancies in depth and resolution between HST and JWST images and not the selection of the sample of galaxies used to compare these data.
Figure 9: The fraction of the NUV luminosity per galaxy that originates in clumps. It is calculated only for clumpy galaxies and plotted as a function of stellar mass of the host galaxy in each panel. Clumps must be at least as bright as the lowest minimum fractional luminosity listed in Table 1 at a given redshift. Redshift increases from left to right and is indicated at the top of each panel. Measurements for individual galaxies are shown as small gray dots. At each redshift, medians are calculated in three bins in mass and are indicated with large black hexagons and horizontal error bars. The 16th and 84th percentiles of the fraction of NUV light in each bin are shown as the vertical error bars. Medians and percentiles are only calculated in mass bins that have at least 10 galaxies. The median fraction of NUV light from clumps is $\sim 15-20\%$ and changes little with redshift or the stellar mass of the host galaxy. Our measurements are consistent within uncertainties with those from Guo et al. (2015) over $2<z<3$, which are shown as black diamonds in the leftmost panel (see text for details).
In Figure 8, $f_{\rm{clumpy}}$ is plotted as a function of redshift for HST (indicated with dashed lines) and JWST data (indicated with solid lines). The limit on the luminosities of the clumps with respect to their hosts is varied in each panel. This is the minimum fraction of the host galaxy NUV luminosity a clump must have. It is highest in the left panel and selected to be 8% that of the host luminosity. This value is adopted from pre-JWST studies (Guo et al., 2015; Shibuya et al., 2016). It is lowered to 5% in the middle panel and 2% in the right panel. The latter is the lowest clump minimum fractional luminosity measured in this work (see Section 4.2).
[…]
We attribute the results discussed above to the difference in sensitivity between HST and JWST. The latter can detect point sources that are $\sim 2-5$ times fainter than the former, depending on the wavelength [^1] https://jwst-docs.stsci.edu/jwst-opportunities-and-policies/jwst-general-science-policies/guidelines-for-proposals-where-jwst-and-hst-overlap-in-capabilities#gsc.tab=0. It is therefore not surprising that $f_{\rm{clumpy}}$ is systematically higher when measured with JWST images than with HST images, when the faintest clumps are considered. Furthermore, the clump fractions from the two datasets are consistent at the highest clump minimum fractional luminosity of 8%, despite the significant differences in angular resolution. The greatest difference in resolution is that between the NIRCam/F150W and WFC3/F160W bandpasses, the images in which are used to measure $f_{\rm{clumpy}}$ at $z\sim 5.75$. The full-width at half maximum of the PSF of the former is $\approx$0.″065 (Zhuang & Shen, 2024) and that of the latter is 0.″18 (Koekemoer et al., 2011), a factor of 3 difference.
6 Contribution of Clumps to the Rest-Frame NUV Light of Galaxies
We now measure the fraction of the NUV luminosity originating in clumps, $C_{\rm{UV}}$, for our sample of galaxies. For a given galaxy in a given redshift bin, it is calculated by summing the flux of all of its clumps in the NIRCam filter that corresponds to the rest-frame NUV (Section 3) and dividing by the flux of the host galaxy in the same filter. Only clumpy galaxies are used to compute this. All clumps are selected to be at least as bright as the lowest minimum fractional luminosity listed in Table 1 at a given redshift.
$C_{\rm{UV}}$ is shown as a function of stellar mass and redshift in Figure 9. Individual measurements are shown as small gray points. At a given redshift, medians are calculated in three bins in mass and are indicated with large black hexagons and horizontal error bars. In each bin, the 16th and 84th percentiles are also computed and shown as vertical error bars. Medians and percentiles are only calculated for mass bins with at least 10 galaxies.
At all masses and redshifts considered, the median $C_{\rm{UV}}$ is $15-20\%$. The 16th percentiles are as low as 4% and 84th percentiles are as high as 35%. They also vary little with mass. We note that several galaxies have very high fractions ($80\%$) of their NUV light emerging from clumps. These are found predominantly at the lowest masses considered.
Our measurements at $z\sim 2.75$ are consistent with results from Guo et al. (2015) at overlapping redshifts and stellar masses. Their calculations of $C_{\rm{UV}}$ over $2<z<3$ are shown as small, black diamonds in the leftmost column of Figure 9. In that work, $C_{\rm{UV}}$ was computed as a function of stellar mass for galaxies in two survey fields, GOODS-S and UDS. The running average over both fields is shown in the Figure. Their $C_{\rm{UV}}$ values were corrected for incompleteness and averaged over all galaxies, including non-clumpy systems. In contrast, our $C_{\rm{UV}}$ values are not corrected for completeness and are computed only for clumpy galaxies. Their $C_{\rm{UV}}$ values are systematically higher than ours at fixed mass by $3-12\%$. Despite this, our measurements and theirs are consistent to within the uncertainties.
At higher redshifts, measurements of $C_{\rm{UV}}$ from other studies are scarce. Bowler et al. (2017) mentioned that clumps made up the majority of the UV light in their sample of 25 bright Lyman-break galaxies at $z\simeq 7$. Chen et al. (2023) examined the JWST/NIRCam images of a sample of 12 UV-bright galaxies over $z\simeq 6-8$. They found that clumps, on average, account for 76% of the light in F200W for the galaxies in their sample. These measurements exceed our median $C_{\rm{UV}}$ values of $15-20\%$ at similar redshifts. However, these previous works considered clumps at all radii in their samples of galaxies, whereas only off-center clumps are investigated in this work. The higher $C_{\rm{UV}}$ reported by these studies is likely the result of light from the centers of galaxies being included in their measurements.
7 Clumps and the Evolution of Galaxies
In this Section, we examine the correlations between two physical properties of galaxies, stellar mass and sSFR, and the fraction of clumpy galaxies. The former is shown in Figure 10 and the latter in Figure 11. Each is considered in its own subsection below. All uncertainties on $f_{\rm{clumpy}}$ are calculated assuming Bayesian confidence intervals on the binomial distribution.
7.1 Stellar Mass vs. Fraction of Clumpy Galaxies
Figure 10: $f_{\rm{clumpy}}$ as a function of host galaxy stellar mass using measurements from this work (left panel). Our measurements are shown in various colors for various redshift bins. The clump minimum fractional luminosities listed in Table 1 are used to compute $f_{\rm{clumpy}}$. Because they vary with stellar mass, $f_{\rm{clumpy}}$ is measured again for the lowest-mass galaxies, this time using the same as those for higher-mass galaxies. These alternative measurements are shown as fainter points connected with dashed lines. The present study finds a positive correlation between stellar mass and the fraction of clumpy galaxies, $f_{\rm{clumpy}}$, over $2<z<12$. In the right panel, our $f_{\rm{clumpy}}$ measurements at $z\sim 2.75$ are compared with those from pre-JWST studies over $2<z<3$, plotted as various symbols. Two studies (Guo et al. 2015 and Sattari et al. 2023) found an anti-correlation between stellar mass and $f_{\rm{clumpy}}$. Another (Huertas-Company et al., 2020) found a positive correlation, but lower $f_{\rm{clumpy}}$ values than ours by about a factor of 3.
The fraction of clumpy galaxies is measured as a function of stellar mass of their host galaxy and in various bins in redshift and plotted in the left panel of Figure 10. The fraction is calculated in three bins, spanning stellar masses $8<\log\left(M_{\star}/M_{\odot}\right)<9$, $9<\log\left(M_{\star}/M_{\odot}\right)<10$, and $10<\log\left(M_{\star}/M_{\odot}\right)<11$. These are shown as hexagons connected with solid lines in a different color for each redshift bin. Fractions are only computed for bins that have at least 10 galaxies. Clumps are selected according to the minimum fractional luminosities as a function of stellar mass and redshift listed in Table 1.
Figure 11: The correlation between the fraction of clumpy galaxies ($f_{\rm{clumpy}}$) and sSFR is plotted. The former is calculated in three bins in stellar mass, indicated with various line styles, and three bins in sSFR. Redshift increases from top to bottom and is indicated in the top left corner of each panel. At the lowest stellar masses, there is a mild correlation or anti-correlation between sSFR and $f_{\rm{clumpy}}$, depending on the redshift. At higher masses, there is no clear correlation, once uncertainties are taken into account. Measurements from Shibuya et al. (2016) over $2<z<3$ are shown in the top row as gray symbols.
At all redshifts, $f_{\rm{clumpy}}$ increases with the stellar mass. At fixed mass, as redshift increases, $f_{\rm{clumpy}}$ decreases. At the lowest redshifts, $z\sim 2.75$, $f_{\rm{clumpy}}$ increases from $\sim 30\%$ at $\log\left(M_{\star}/M_{\odot}\right)\sim 8.5$ to $\sim 70\%$ at $\log\left(M_{\star}/M_{\odot}\right)\sim 9.5$ and $\sim 90\%$ at $\log\left(M_{\star}/M_{\odot}\right)\sim 10.5$. The rate at which $f_{\rm{clumpy}}$ increases with stellar mass is similar at higher redshifts. It grows from $\sim 20\%$ at the lowest masses to $\sim 70\%$ at the highest masses at $z\sim 4.25$. At $z\sim 5.75$, it rises from $\sim 10\%$ at $\log\left(M_{\star}/M_{\odot}\right)\sim 8.5$ to $\sim 40\%$ at $\log\left(M_{\star}/M_{\odot}\right)\sim 9.5$. Over the same range in mass at $z>6.75$, it increases from $\sim 3\%$ to $\sim 15\%$.
The results above were found by selecting clumps using the minimum fractional luminosities listed in Table 1, which vary with stellar mass and redshift. However, the correlation between stellar mass and $f_{\rm{clumpy}}$ still holds when selecting clumps with the same minimum fractional luminosity at all masses. The fraction of clumpy galaxies is calculated again for galaxies with stellar mass $\log\left(M_{\star}/M_{\odot}\right)\sim 8.5$, with the minimum fractional luminosities this time being the same as those applied to higher-mass galaxies. These alternative measurements are plotted as faint hexagons connected with dashed lines in the left panel in Figure 10. The effect of lowering the minimum fractional luminosity for the lowest-mass galaxies is an increase in $f_{\rm{clumpy}}$ by $\sim 10\%$ at fixed redshift. At $z\lesssim 5.75$, this makes the slope between stellar mass and $f_{\rm{clumpy}}$ slightly shallower than in the case where the minimum fractional luminosities vary with mass. At $z>6.5$, the slope is nearly flat. However, the uncertainties on $f_{\rm{clumpy}}$ are the largest at these redshifts. More data are needed to determine the correlation between stellar mass and $f_{\rm{clumpy}}$ at the highest redshifts considered in this work.
We note that the correlation observed in the present work may result from selection effects. Clumps are identified based on their brightnesses relative to their hosts. Therefore, clumps detected in more massive galaxies must be brighter, and hence, easier to detect, than those in less massive galaxies. Indeed, recent work on lensed galaxies has found that more massive galaxies contain more massive clumps (Claeyssens et al., 2025). The physical properties of clumps that are identified in this work, such as stellar masses and SFRs, are needed to determine the impact of these selection effects.
In the right panel in Figure 10, we compare the correlation between stellar mass and $f_{\rm{clumpy}}$ at $z\sim 2.75$ with that measured by other studies over $2<z<3$. These measurements are shown as various symbols. Guo et al. (2015) and Sattari et al. (2023) found that, as stellar mass increases, $f_{\rm{clumpy}}$ decreases. It was found to be $\sim 60\%$ at $\log\left(M_{\star}/M_{\odot}\right)\sim 9.5$ and $\sim 40\%$ at $\log\left(M_{\star}/M_{\odot}\right)\sim 11$. Both studies used HST data covering the rest-frame UV and developed detection algorithms to search for clumps. In contrast, Huertas-Company et al. (2020) found that, as stellar mass increases, $f_{\rm{clumpy}}$ increases. From $\log\left(M_{\star}/M_{\odot}\right)\sim 9$ to $\log\left(M_{\star}/M_{\odot}\right)\sim 11$, it grows from $\sim 20\%$ to $\sim 30\%$. In that study, neural networks were trained on mock HST images from the VELA simulations (Ceverino et al., 2014) and used to identify clumps in real HST images spanning the rest-frame UV and optical. Their sample consisted of 1,575 galaxies over $0.5<z<3$ and $9<\log\left(M_{\star}/M_{\odot}\right)<12$.
Results from the present work at $z\sim 2.75$ disagree with those from other studies at similar redshifts. We find a positive correlation between stellar mass and $f_{\rm{clumpy}}$, such that, as stellar mass increases from $\log\left(M_{\star}/M_{\odot}\right)\sim 8.5$ to $\log\left(M_{\star}/M_{\odot}\right)\sim 10.5$, $f_{\rm{clumpy}}$ increases about threefold from $\sim 30\%$ to $\sim 90\%$. Guo et al. (2015) and Sattari et al. (2023) found an anti-correlation and, at fixed mass, overall lower fractions than those measured in this study. Huertas-Company et al. (2020) also found a positive correlation, but lower $f_{\rm{clumpy}}$ values than ours by a factor of 2-3 at fixed mass.
At $z>2.75$, results from this work qualitatively agree with those from other studies. We find that more massive galaxies are more likely to host clumps at all redshifts. Earlier studies that examined galaxies over $6\lesssim z\lesssim 8$ found that brighter galaxies were more likely to have clumps or multiple cores and fainter galaxies were more likely to have one component (Kawamata et al., 2015; Bowler et al., 2017).
7.2 Specific SFR vs. Fraction of Clumpy Galaxies
We now examine the correlation between sSFR and $f_{\rm{clumpy}}$. This is shown in Figure 11. It is examined in the same bins in stellar mass considered in the previous Section, which are indicated with various line styles in the Figure. For each bin in mass, $f_{\rm{clumpy}}$ is measured in three 0.5-dex-wide bins in sSFR. They are centered at $\log\left(\textrm{sSFR/yr}^{-1}\right)\sim-9.25,\sim-8.75,$ and $\sim-8.25$. Redshift increases from the top row to the bottom.
At the lowest stellar masses of $\left(\log\left(M_{\star}/M_{\odot}\right)\sim 8.5\right)$, there is a mild correlation or anti-correlation between sSFR and $f_{\rm{clumpy}}$, depending on the redshift. At higher stellar masses, there is little to no correlation. At $z\sim 2.75$ and $\log\left(M_{\star}/M_{\odot}\right)\sim 8.5$, as sSFR increases, $f_{\rm{clumpy}}$ increases from 20% to 35%. At higher stellar masses, as sSFR increases, $f_{\rm{clumpy}}$ varies by $\leq 10\%$. Similar behavior is observed at $z\sim 4.25$, once uncertainties are taken into account. For the lowest-mass galaxies, as sSFR increases, $f_{\rm{clumpy}}$ increases from 7% to 23%. For higher-mass galaxies, as sSFR increases, $f_{\rm{clumpy}}$ varies by up to $\sim 30\%$. However, the greatest uncertainties are also this large, particularly at the lowest sSFRs and highest masses. Therefore, changes in $f_{\rm{clumpy}}$ with sSFR are consistent with little to no correlation at these masses. At $z\sim 5.75$, for the lowest-mass galaxies, as sSFR increases, $f_{\rm{clumpy}}$ decreases from 22% to 10%. However, uncertainties on $f_{\rm{clumpy}}$ at the lowest sSFRs are relatively large, $\sim 10-20\%$. Therefore, this is not a significant anti-correlation. At higher masses, as sSFR increases, $f_{\rm{clumpy}}$ rises from zero to 43% and declines to 38%. This is due to the lowest bin in sSFR having only one galaxy and thus is not significant. Finally, at $z>6.5$, at all masses, as sSFR increases, $f_{\rm{clumpy}}$ increases from zero and plateaus. However, this is not a significant correlation because the lowest sSFR bins have at most three galaxies. Consequently, the uncertainties at the lowest sSFRs are very large (up to $\sim 50\%$) and are consistent with little to no trend with sSFR.
The lack of a significant correlation between sSFR and $f_{\rm{clumpy}}$ at $z\sim 2.75$ for galaxies of stellar mass $M_{\star}\geq 10^{9}M_{\odot}$ is generally consistent with what previous studies found at similar redshifts, masses, and sSFRs. As sSFR increases over $\sim-9.25<\log\left(\textrm{sSFR/yr}^{-1}\right)<\sim-8.75$, Shibuya et al. (2016) measured $f_{\rm{clumpy}}$ to increase by $5-10\%$ for galaxies with masses over $9<\log\left(M_{\star}/M_{\odot}\right)<11$ and redshifts over $2<z<3$. For higher-mass galaxies, $10<\log\left(M_{\star}/M_{\odot}\right)<11$, $f_{\rm{clumpy}}$ then drops by $\sim 20\%$ at higher sSFRs. However, the uncertainty on this is of similar size, therefore, this decline is not significant. Similar increases of $5-10\%$ are seen in our sample at similar masses and sSFRs, despite our  $f_{\rm{clumpy}}$ values being about $20-40\%$ higher. Our uncertainties on $f_{\rm{clumpy}}$ at these masses and sSFRs range from $3-7\%$. Those from Shibuya et al. (2016) are similar, about $5-7\%$. Therefore, this small change in $f_{\rm{clumpy}}$ with sSFR is not significant. Similarly, Huertas-Company et al. (2020) found no correlation between $f_{\rm{clumpy}}$ and the location of galaxies relative to the SFMS over $2<z<3$.
8 Comparison with Hydrodynamic Simulations
Here we discuss the implication of our results to galaxy formation scenarios. There are two key results of this work. First, the fraction of clumpy galaxies, $f_{\rm{clumpy}}$, increases with decreasing redshift over $2<z<12$. For galaxies with stellar mass $M{\star}\geq 10^{9}M_{\odot}$, $f_{\rm{clumpy}}$ is $\sim 20\%$ at $z>6.5$ and $\sim 80\%$ at $z\sim 2.75$. $f_{\rm{clumpy}}$ is found to be higher, at fixed redshift, and increase more rapidly with decreasing redshift than what was found in many previous studies (Section 5 and Figure 7). We interpret this as a consequence of the higher sensitivity of the JWST compared to the HST, which nearly all earlier studies relied on (Section 5.3). Second, at a fixed redshift, as stellar mass increases, the fraction of clumpy galaxies increases (Section 7.1 and Figure 10).
In this Section, we interpret our observations within the context of clump formation over $2<z<12$. Our results are compared with those from simulations of clumpy galaxies. In doing so, we identify what might be the dominant processes for clumpy galaxy formation as a function of mass and redshift. In Section 9, our results are studied in the context of other observations.
Figure 12: $f_{\rm{clumpy}}$ as a function of redshift in three stellar mass bins. This is indicated with thick black lines. Predictions for $f_{\rm{clumpy}}$ based on simulations are also shown. Their corresponding uncertainties are shown as shaded regions. The fraction of clumpy galaxies that formed from violent disk instabilities (VDI; Mandelker et al., 2017) at $z<6$ is shown as gray, red, and green shaded regions, in the order of increasing stellar mass. Predictions for clumpy galaxies that formed from major mergers over $6\lesssim z\lesssim 9$ (Nakazato et al., 2024) are shown as a purple line and shaded region. Observed pair fractions for galaxies of stellar mass over $8<\log\left(M_{\star}/M_{\odot}\right)<10$ and redshift over $3<z<9$ (Puskás et al., 2025) are shown as thin gray lines in various styles. Trends for the least massive, intermediate, and most massive galaxies are indicated with dotted, dash-dot, and solid lines, respectively. The results from this work are consistent with $f_{\rm{clumpy}}$ due to VDI at $z<6$ and $\log\left(M_{\star}/M_{\odot}\right)\geq 9$ and that due to major mergers at $z>6$ and $\log\left(M_{\star}/M_{\odot}\right)\leq 9$. The observed pair fractions at $\log\left(M_{\star}/M_{\odot}\right)\leq 9$ are also consistent with the observed $f_{\rm{clumpy}}$ measured here at similar masses and redshifts. Taking these comparisons together, our results are consistent with a physical scenario in which VDI is the dominant mechanism behind clumpy galaxies at low redshifts and intermediate and high masses and mergers dominate at high redshifts and the lowest masses.
The $f_{\rm{clumpy}}$ measurements are compared with those from zoom-in hydrodynamic cosmological simulations of clumpy galaxies examined by Mandelker et al. (2017) and Nakazato et al. (2024) in Figure 12. The former studied 34 galaxies over $1\lesssim z\lesssim 6$ from the VELA simulations (Ceverino et al., 2014; Zolotov et al., 2015), which had stellar masses of $\log\left(M_{\star}/M_{\odot}\right)\gtrsim 9$ at $z=2$. In that work, clumps formed largely through VDI. The latter analyzed 62 galaxies over $5.5<z<9.5$ from the FirstLight simulations (Ceverino et al., 2017), which had stellar masses in the range of $(0.5-6)\times 10^{10}M_{\odot}$ at $z=5$. Most clumps in these simulations formed from major mergers. The $f_{\rm{clumpy}}$ measurements from the present work are plotted as thick black lines in Figure 12. Predictions from simulations are shown as lines with shaded regions. $f_{\rm{clumpy}}$ is shown for various mass ranges, where applicable, and indicated with various line styles.
The $f_{\rm{clumpy}}$ measurements in the present work are consistent with those from Mandelker et al. (2017) at $z\lesssim 5.75$ and $\log\left(M_{\star}/M_{\odot}\right)\geq 9$. The $f_{\rm{clumpy}}$ measurements from Mandelker et al. (2017) increase rapidly with redshift from $\sim 20\%$ at $z\sim 1$ to $\sim 50\%$ $z\sim 2$, above which they asymptote. They have uncertainties of $\sim 20\%$, which result from Poisson noise due to a limited number of galaxies examined in that work. Our $f_{\rm{clumpy}}$ measurements peak at $\sim 70-90\%$ at $z\sim 2.75$ and decline with increasing redshift, while those from Mandelker et al. (2017) do not change with redshift within the uncertainties. The discrepancy in $f_{\rm{clumpy}}$ as a function of redshift between the theoretical predictions and our observations may arise from differences in the way clumps are identified. Mandelker et al. (2017) searched for clumps in projected density maps of their galaxies and computed $f_{\rm{clumpy}}$ using those with stellar masses $10^{7}~M_{\odot}$, SFRs of at least 0.1% those of their host galaxies’ disks. In this work, clumps are detected in rest-frame NUV images and selected to be compact objects that are at least as bright as a fraction of their hosts’ NUV luminosity, depending on redshift and stellar mass of the host galaxy. In both studies, clumps are selected to be off-center. However, at the redshifts and stellar masses where the two studies overlap, the measurements are largely consistent within the uncertainties.
Our $f_{\rm{clumpy}}$ measurements are also consistent with those from Nakazato et al. (2024) at $z\sim 5.75$ and $z>6.5$ and $\log\left(M_{\star}/M_{\odot}\right)<9$. As redshift decreases, our measured $f_{\rm{clumpy}}$ increase from 3% at $z>6.5$ to 11% at $z\sim 5.75$. Those from Nakazato et al. (2024) are roughly constant with redshift at $\sim 10\%$. Galaxies studied by that work had stellar masses of $\lesssim 10^{9}~M_{\odot}$ at $z\geq 8.5$. Therefore, trends in $f_{\rm{clumpy}}$ at the lowest masses considered in this work are in overall agreement with those from Nakazato et al. (2024) at redshifts where there is overlap. However, clumps were identified differently between these two methods. In Nakazato et al. (2024), clumps were found in stellar mass surface density maps of young (<10<10 Myr) stars. Detectable clumps had stellar masses of $\gtrsim 10^{7}~M_{\odot}$. This is comparable to the faintest clumps that can be detected in this work at $z\sim 5.75$ and over $8<\log\left(M_{\star}/M_{\odot}\right)<9$, assuming clumps have similar NUV mass-to-light ratios as their hosts at these redshifts and stellar masses. In Nakazato et al. (2024), galaxies were defined to be clumpy if they had at least two clumps. In this work, a galaxy must have at least one off-center clump to be considered clumpy.
It is unclear from the simulations which processes dominate the formation of clumpy galaxies at other combinations of redshift and stellar mass, namely, low redshift ($z<5.75$) and low mass ($\log\left(M_{\star}/M_{\odot}\right)<9$), and high redshift ($z>5.75$) and high mass ($\log\left(M_{\star}/M_{\odot}\right)\geq 9$). In the case of the former, only two galaxies in the sample of Mandelker et al. (2017) had such low masses at $z=2$. Predictions for $f_{\rm{clumpy}}$ for galaxies at these masses and redshifts are needed. In the case of the latter, predictions from simulations that produce clumps through VDI are needed out to higher redshifts in order to distinguish clumps formed through VDI and those from mergers. Our $f_{\rm{clumpy}}$ measurements at intermediate stellar masses ($9<\log\left(M_{\star}/M_{\odot}\right)<10$) are about four times higher than those from Nakazato et al. (2024) at $z\sim 5.75$. This suggests that major mergers may not significantly contribute to clumpy galaxy formation at these masses and redshifts (see below for support for this from observations). However, confirming this will require $f_{\rm{clumpy}}$ measurements from simulations of clumpy galaxies that form via other mechanisms at these redshifts.
9 Discussion
Comparison of our observations and predictions from simulations suggests that VDI and mergers are the likely dominant mechanisms for clumpy galaxy formation, depending on their redshift and stellar mass, in agreement with other studies (e.g., Guo et al., 2015; Shibuya et al., 2016). VDI appears to be the dominant mechanism for forming clumps at $z\lesssim 5.75$ and at intermediate-to-high stellar masses ($\log\left(M_{\star}/M_{\odot}\right)\geq 9$). Mergers likely dominate at $z\gtrsim 5.75$ and low stellar masses ($8<\log\left(M_{\star}/M_{\odot}\right)<9$). We now assess the plausibility of these mechanisms by considering observations of the morphologies, gas kinematics, and pair fractions of star-forming galaxies with similar redshifts and stellar masses as those studied in this work. We find that observations largely support the scenario described above.
For clumps to form through VDI, their host galaxies must possess marginally stable disk components (e.g., Bournaud et al., 2007; Agertz et al., 2009; Dekel et al., 2009a; Ceverino et al., 2010; Mandelker et al., 2014). Star-forming galaxies over $2<z\lesssim 6$ and $\log\left(M_{\star}/M_{\odot}\right)>8$ have typical Sérsic indices of $n\simeq 1$ (van der Wel et al., 2014; Shibuya et al., 2015, 2016; Nedkova et al., 2021; Ormerod et al., 2024; Sun et al., 2024; Varadaraj et al., 2024; Ward et al., 2024; Miller et al., 2025; Yang et al., 2025), which suggests they have disk-like structures. At higher redshifts, Sérsic indices are generally larger, which implies disk-like structure becomes less common (Sun et al. 2024; Miller et al. 2025; see also Ferreira et al. 2023; Kartaltepe et al. 2023).
[…]
For clumps to form mainly through mergers, it is expected that the fraction of clumpy galaxies and that of galaxies undergoing mergers are similar. The $f_{\rm{clumpy}}$ measurements from the present study are compared with pair fractions measured by Puskás et al. (2025) in Figure 12. The latter are shown as thin gray lines. Various bins in stellar mass are indicated with various line styles. In that work, the pair fraction was measured over $3<z<9$ and $8<\log\left(M_{\star}/M_{\odot}\right)<10$ in both of the JADES fields. Puskás et al. (2025) computed the frequency of galaxies with major companions at projected separations of 5305-30 kpc and stellar masses of at least $1/4$ those of primary galaxies. Their pair fractions agree with other recent measurements of the pair fraction at high redshift with JWST (Dalmasso et al., 2024; Duan et al., 2025). Our $f_{\rm{clumpy}}$ measurements over $8<\log\left(M_{\star}/M_{\odot}\right)<9$ are consistent with pair fractions from Puskás et al. (2025) over the same range in mass and the overlapping range in redshift. This suggests that major mergers are the dominant mechanism for clumpy galaxy formation at low stellar masses over the redshift range $3<z<9$.
At higher stellar masses ($9<\log\left(M_{\star}/M_{\odot}\right)<10$), $f_{\rm{clumpy}}$ is consistent with the pair fraction only at $z>6.5$. At lower redshifts, the former is three to four times higher than the latter. This implies that major mergers are not the dominant process through which clumps form at these masses and redshifts. This is also the case at the highest masses considered here. Duncan et al. (2019) calculated pair fractions over $0.5<z<6$ and $\log\left(M_{\star}/M_{\odot}\right)>9.7$ using HST data. The pair fraction at $\log\left(M_{\star}/M_{\odot}\right)>10.3$ was found to be $\sim 10-15\%$ over $2\lesssim z\lesssim 4$ and increases to $\sim 30-40\%$ at $z\sim 6$. This is inconsistent with our $f_{\rm{clumpy}}$ over $10<\log\left(M_{\star}/M_{\odot}\right)<11$, which decreases from $90\%$ at $z\sim 2.75$ to $73\%$ at $z\sim 4.25$.
10 Conclusions
We use JWST/NIRCam imaging from the JADES survey (Eisenstein et al., 2023a, b; Rieke et al., 2023b; Bunker et al., 2024; D’Eugenio et al., 2024) to identify clumpy star-forming galaxies over $2<z<12$. The evolution of the fraction of clumpy galaxies, $f_{\rm{clumpy}}$, with redshift and correlations with other physical properties were examined. A sample of 9,121 galaxies with stellar masses $\log\left(M_{\star}/M_{\odot}\right)\geq 8$ and $\log\left(\rm{sSFR/yr}^{-1}\right)\geq-9.5$ was selected by fitting their combined HST and JWST SEDs. Clumps were selected in a way similar to earlier studies based on HST (e.g., Guo et al., 2015, 2018; Shibuya et al., 2016; Sattari et al., 2023). The algorithm in Sattari et al. (2023) was used to find off-center clumps in rest-frame NUV images of galaxies. The completeness of the clump detection procedure was determined by injecting fake clumps into images of real galaxies. Clumps were selected to be $\geq 50\%$ complete as a function of redshift and stellar mass.
Our main conclusions are summarized below.
  1. 1.
    We find that $f_{\rm{clumpy}}$ increases with decreasing redshift from $\sim 20\%$ at $z>6.5$ to $\sim 80\%$ at $z\sim 2.75$ (Section 5 and Figure 7). This is similar to the relative increase in the cosmic volume-averaged star-formation density over the same range in redshift. Compared with the bulk of pre-JWST studies, $f_{\rm{clumpy}}$ from this work is higher at fixed redshift and increases more steeply with decreasing redshift.
  2. 2.
    The higher $f_{\rm{clumpy}}$ found in this work is due to the higher sensitivity of JWST compared with HST (Section 5.3 and Figure 8). We calculate $f_{\rm{clumpy}}$ for a sample of 1,227 galaxies that are detected in both HST and JWST images and compare measurements from both sets of images. When clumps are selected to have NUV luminosities that are $\geq 8\%$ those of their host galaxies, both measurements of $f_{\rm{clumpy}}$ are consistent with each other. When fainter clumps are selected, $f_{\rm{clumpy}}$ from JWST data are systematically higher, which is inconsistent with that from HST.
  3. 3.
    Off-center clumps make up $15-20\%$ of the rest-frame NUV luminosity of clumpy galaxies (Section 6 and Figure 9). There is no significant dependence on redshift or stellar mass of the host galaxy. At $z\sim 2.75$, our results are consistent with previous studies.
  4. 4.
    We examine correlations between $f_{\rm{clumpy}}$ and physical properties of galaxies. There is a strong correlation between stellar mass and $f_{\rm{clumpy}}$, such that more massive galaxies have higher $f_{\rm{clumpy}}$ (Section 7.1 and Figure 10). This is the opposite trend of most previous studies. We find a mild correlation between sSFR and $f_{\rm{clumpy}}$  at the lowest masses and little to none at higher masses (Section 7.2 and Figure 11). This is consistent with the results of earlier studies.
  5. 5.
    Our $f_{\rm{clumpy}}$ measurements are consistent with a physical scenario in which clumps form largely through violent disk instabilities at $z\lesssim 5.75$ and over $9<\log\left(M_{\star}/M_{\odot}\right)<11$ and through mergers at $z\gtrsim 5.75$ and over $8<\log\left(M_{\star}/M_{\odot}\right)<9$ (Section 8 and Figure 12). This is done by comparing $f_{\rm{clumpy}}$ measured from simulations with the observations. This physical scenario is bolstered when considering the observed morphologies, gas kinematics, and pair fractions of star-forming galaxies at redshifts and stellar masses that overlap with the sample examined in this work (Section 9).
We release a catalog of physical properties derived from SED fits, the empirical PSFs, and code to reproduce Figure 7 at https://github.com/alex-delavega/clumpy_galaxies_jades_2_z_12.
We thank the JADES team for designing and preparing their observations and releasing a rich public dataset. We also thank the STScI staff for enabling this science. This research has made use of the SVO Filter Profile Service “Carlos Rodrigo”, funded by MCIN/AEI/10.13039/501100011033/ through grant PID2023-146210NB-I00. All of the JWST data used in this paper can be found in MAST (Rieke et al., 2023a; Williams et al., 2023a), as can all of the HST data used in this paper (Faber, 2011).
Appendix A Empirical Point Spread Functions
Empirical PSFs were measured using photutils EPSFBuilder routine, which is based on the technique developed by Anderson & King (2000). We adopt the parameters used to construct PSFs in photutils from Zhuang & Shen (2024), who characterized JWST/NIRCam PSFs in Cycle 0 calibration programs. In the present work, PSFs were created using bright (F277W mag <22<22), unsaturated, and isolated point sources that were visually inspected in each NIRCam bandpass. For the wide bandpasses, 16 point sources in GOODS-N and 17 in GOODS-S were used to estimate PSFs. For the medium bandpasses, six to ten point sources in either field were used, depending on the bandpass.
Empirical PSFs and their curves of enclosed energy are shown in Figure 13. The latter are measured by summing the light from the PSFs in circular apertures at various radii and are plotted as solid lines in the Figure. At each radius, the value of the curve indicates the fraction of the total light enclosed. We compare the PSFs measured here with those from Ji et al. (2024), who calculated empirical PSFs in the JADES GOODS-S field. Their enclosed energy curves are shown as dashed lines in the second row. There is good agreement overall. At worst, the enclosed energy curves computed in this work for the F430M, F460M, and F480M bandpasses disagree with those from Ji et al. (2024) by 5% at most. For all other bandpasses, they disagree at most by about $2-3\%$.
Figure 13: Empirical point spread functions (PSFs) in 14 JWST/NIRCam bandpasses in JADES GOODS-N and GOODS-S are measured in this work and shown in this figure. In the top row are images of the PSFs in each bandpass in GOODS-S. Below it are the curves of enclosed energy as a function of radius for each PSF. This is repeated in the third and bottom rows for GOODS-N. The PSFs are estimated using the algorithm by Anderson & King (2000) and implemented in photutils (Bradley, 2023). The parameters that are used to construct the PSFs with photutils are adopted from Zhuang & Shen (2024). The curves of enclosed energy that are measured in this work agree to within 5% with those measured by Ji et al. (2024), who also calculated empirical PSFs in the JADES GOODS-S field. The enclosed energy curves from that work are shown as dashed lines in the second row.
Appendix B Measurements of the Fraction of Clumpy Galaxies from Previous Studies
In this Section, we describe how measurements of $f_{\rm{clumpy}}$ were extracted from the earlier studies listed in Table 2. Following Shibuya et al. (2016), $f_{\rm{clumpy}}$ from Lotz et al. (2006); Ravindranath et al. (2006); Conselice & Arnold (2009); Oesch et al. (2010); Law et al. (2012); Jiang et al. (2013); Kawamata et al. (2015) and Curtis-Lake et al. (2016) were measured using the fractions of galaxies undergoing mergers or exhibiting disturbed morphologies or multiple cores reported in these works. These come from the following tables, figures, or sections in these studies: Table 1 in Lotz et al. (2006); Table 2 in Ravindranath et al. (2006); Table 2 in Conselice & Arnold (2009); Section 4 in Oesch et al. (2010); Figure 17 in Law et al. (2012); Section 3.3 in Jiang et al. (2013); Section 3.1 in Kawamata et al. (2015); and Figure 17 in Curtis-Lake et al. (2016). $f_{\rm{clumpy}}$ from Bowler et al. (2017) was obtained from the fraction of mergers given in their Section 4.4.
Following Guo et al. (2015), $f_{\rm{clumpy}}$ from Elmegreen et al. (2007) was inferred using the abundances of clump cluster, spiral, and elliptical galaxies, given in their Figure 6. Other categories (chain galaxies, double nuclei, and tadpoles) were excluded as they generally have small axis ratios and would not have been included according to our sample selection (Section 3).
For all other studies listed in Table 2, $f_{\rm{clumpy}}$ measurements were given explicitly. These come from the following tables, figures, or sections in these works: Table 1 in Wuyts et al. (2012); Section 3.1 in Tadaki et al. (2014); Figure 10 in Guo et al. (2015); Figure 4 in Shibuya et al. (2016); Figure 8 in Ribeiro et al. (2017); Figure 16 in Huertas-Company et al. (2020); Figure 7 in Sattari et al. (2023); Figure 12 in Tan et al. (2024); and Section 4.1 in Harikane et al. (2025).

### v2

@@
-The Fraction of Clumpy Galaxies in JWST Surveys over $2<z<12$
+The Fraction of Clumpy Galaxies in JADES over $2<z<9$
 Abstract
@@
 Studying them can shed light on how they form and evolve and inform theoretical models of galaxy evolution.
-We examine the evolution of clumpy galaxies with redshift and stellar mass over $2<z<12$ with James Webb Space Telescope (JWST) imaging from the JWST Advanced Extragalactic Survey (JADES).
+We examine the evolution of clumpy galaxies with redshift and stellar mass over 2<z<9{2<z<9} with James Webb Space Telescope (JWST) imaging from the JWST Advanced Extragalactic Survey (JADES).
 Off-center clumps are detected in the rest-frame near-ultraviolet (NUV) using similar techniques to those in earlier studies based on Hubble Space Telescope (HST) images.
 This is done for a sample of 9,121 star-forming galaxies with stellar masses $\log\left(M_{\star}/M_{\odot}\right)\geq 8$.
-The fraction of clumpy galaxies, $f_{\rm{clumpy}}$, increases from $\sim 20\%$ at $z>6.5$ to $\sim 80\%$ at $z\sim 2.75$ at $\log\left(M_{\star}/M_{\odot}\right)\geq 9$.
+The fraction of clumpy galaxies, $f_{\rm{clumpy}}$, increases from $\sim 10\%$} at $z\sim 7.75$} to $\sim 70\%$} at $z\sim 2.75$ at $\log\left(M_{\star}/M_{\odot}\right)\geq 9$.
 Our $f_{\rm{clumpy}}$ values are generally higher at fixed redshift and increase faster with decreasing redshift than what studies based on HST data found, which we attribute largely to the higher sensitivity of JWST.
-Clumps make up $15-20\%$ of the NUV luminosity of clumpy galaxies, with little dependence on redshift or stellar mass.
-$f_{\rm{clumpy}}$ correlates with stellar mass and there is little correlation with specific star-formation rate.
+$f_{\rm{clumpy}}$ correlates with stellar mass.
 Our $f_{\rm{clumpy}}$ measurements are compared with those from simulations as well as other observations.
 At low redshifts ($z\lesssim 5.75$) and intermediate-to-high stellar masses ($\log\left(M_{\star}/M_{\odot}\right)\geq 9$), our results suggest gas fragmentation due to violent disk instabilities to be the dominant mechanism for forming clumps.
-At high redshifts and low stellar masses, compression of gas during mergers appears to dominate.
+At high redshifts and intermediate stellar masses, compression of gas during mergers appears to dominate.
 Galaxy evolution(594) — Galaxy structure(622) — High-redshift galaxies(734) — Star forming regions(1565)
-software: astropy (Astropy Collaboration et al., 2013, 2018, 2022), matplotlib (Hunter, 2007), numpy (Harris et al., 2020), photutils (Bradley et al., 2020; Bradley, 2023), scipy (Virtanen et al., 2020)
-1 Introduction
-Early Hubble Space Telescope (HST) imaging of high-redshift galaxies revealed systems with patchy, clumpy, kpc-scale substructures (Cowie et al., 1995; Giavalisco et al., 1996; van den Bergh et al., 1996; Elmegreen & Elmegreen, 2005; Elmegreen et al., 2007).
-Later work showed that these substructures (hereafter, “clumps”) are bright in the rest-frame ultraviolet (UV), are fairly massive ($\gtrsim 10^{7}~M_{\odot}$), have young stellar ages ($\sim 100$ Myr), and make up about $10-50\%$ of the total star-formation rates (SFRs) and about $4-20\%$ of the total stellar mass of their host galaxies, on average (Elmegreen et al., 2008, 2009; Förster Schreiber et al., 2011; Guo et al., 2012, 2015, 2018; Wisnioski et al., 2012; Wuyts et al., 2012, 2013; Livermore et al., 2015; Shibuya et al., 2016; Soto et al., 2017; Zanella et al., 2019; Mehta et al., 2021; Vanzella et al., 2021; Meštrić et al., 2022; Claeyssens et al., 2025).
-These results suggest that clumps were significant sites of star formation in galaxies over $0.5<z<3$, during the peak of cosmic star-formation activity (Madau & Dickinson, 2014).
+software: astropy (Astropy Collaboration et al., 2013, 2018, 2022), LMFIT (Newville et al., 2020), matplotlib (Hunter, 2007), numpy (Harris et al., 2020), photutils (Bradley et al., 2020; Bradley, 2023), scipy (Virtanen et al., 2020)
+I Introduction
+Early Hubble Space Telescope (HST) imaging of high-redshift galaxies revealed systems with patchy, clumpy, kpc-scale substructures (Cowie et al., 1995; Giavalisco et al., 1996; van den Bergh et al., 1996; Elmegreen and Elmegreen, 2005; Elmegreen et al., 2007).
+Later work showed that these substructures (hereafter, “clumps”) are bright in the rest-frame ultraviolet (UV), are fairly massive ($\gtrsim 10^{7}~M_{\odot}$), have young stellar ages ($\sim 100$ Myr), and, on average, make up about $10-50\%$ of the total star-formation rates (SFRs) and about $4-20\%$ of the total stellar mass of their host galaxies (Elmegreen et al., 2008, 2009; Förster Schreiber et al., 2011; Guo et al., 2012, 2015, 2018; Wisnioski et al., 2012; Wuyts et al., 2012, 2013; Livermore et al., 2015; Shibuya et al., 2016; Soto et al., 2017; Zanella et al., 2019; Mehta et al., 2021; Vanzella et al., 2021; Meštrić et al., 2022; Claeyssens et al., 2025; Sok et al., 2025).
+These results suggest that clumps were significant sites of star formation in galaxies over $0.5<z<3$, during the peak of cosmic star-formation activity (Madau and Dickinson, 2014).
 Studies of observed and simulated clumps have proposed two main formation mechanisms.
-The first is violent disk instabilities (VDI), wherein gas collapses under gravitational instabilities in marginally stable gaseous disks (Noguchi, 1998, 1999; Immeli et al., 2004a, b; Bournaud et al., 2007, 2009; Agertz et al., 2009; Dekel et al., 2009a; Ceverino et al., 2010; Romeo et al., 2010; Romeo & Agertz, 2014; Inoue et al., 2016).
-The VDI is sustained by the high gas fractions in high-redshift galaxies (Tacconi et al., 2020, and references therein), which are fed by the accretion of cold gas from the intergalactic medium (Birnboim & Dekel, 2003; Kereš et al., 2005, 2009; Dekel & Birnboim, 2006; Bournaud & Elmegreen, 2009; Dekel et al., 2009b, a; Nelson et al., 2013).
-Observational support for the VDI has been found by examining integral field spectroscopic data, with which maps of gas kinematics may be measured (Förster Schreiber et al., 2011; Genzel et al., 2011, 2014; Wisnioski et al., 2011, 2012; Newman et al., 2012; Mieda et al., 2016).
+The first is violent disk instabilities (VDI), wherein gas collapses under gravitational instabilities in marginally stable gaseous disks (Noguchi, 1998, 1999; Immeli et al., 2004a, b; Bournaud et al., 2007, 2009; Agertz et al., 2009; Dekel et al., 2009b; Ceverino et al., 2010; Romeo et al., 2010; Romeo and Agertz, 2014; Inoue et al., 2016).
+The VDI is sustained by the high gas fractions in high-redshift galaxies (Tacconi et al., 2020, and references therein), fed by the accretion of cold gas from the intergalactic medium (Birnboim and Dekel, 2003; Kereš et al., 2005, 2009; Dekel and Birnboim, 2006; Bournaud and Elmegreen, 2009; Dekel et al., 2009a, b; Nelson et al., 2013).
+Observational support for the VDI comes from integral field spectroscopic data, with which maps of gas kinematics may be measured (Förster Schreiber et al., 2011; Genzel et al., 2011, 2014; Wisnioski et al., 2011, 2012; Newman et al., 2012; Mieda et al., 2016).
 The second mechanism is mergers, during which gas is compressed by tidal forces and/or turbulence due to gravitational interactions between galaxies (Di Matteo et al., 2008; Renaud et al., 2015; Nakazato et al., 2024).
@@
 Clumps could also evolve within their host galaxies and alter their morphologies.
-Theoretical studies and simulations have predicted that clumps experience dynamical friction and/or interactions, which cause them to migrate to the centers of their hosts and contribute to the growth of bulges (Noguchi, 1999; Immeli et al., 2004a, b; Dekel et al., 2009a; Krumholz & Dekel, 2010; Ceverino et al., 2012; Inoue & Saitoh, 2012; Perez et al., 2013; Bournaud et al., 2014; Mandelker et al., 2014, 2017; Dekel et al., 2022).
+Theoretical studies and simulations have predicted that clumps experience dynamical friction and/or interactions, which cause them to migrate to the centers of their hosts and contribute to the growth of bulges (Noguchi, 1999; Immeli et al., 2004a, b; Dekel et al., 2009b; Krumholz and Dekel, 2010; Ceverino et al., 2012; Inoue and Saitoh, 2012; Perez et al., 2013; Bournaud et al., 2014; Mandelker et al., 2014, 2017; Dekel et al., 2022).
 However, other simulations find that clumps are short-lived and dissipate within $\sim 100$ Myr due to stellar feedback, which drives strong outflows of gas (Murray et al.
@@
 2017; Oklopčić et al.
-2017; Meng & Gnedin 2020; see also Fensch & Bournaud 2021).
-In this scenario, stars formed within the clumps could contribute to the growth of the thick disk (Noguchi 1996; Inoue & Saitoh 2014; see also Bournaud et al.
-2007).
+2017; Meng and Gnedin 2020; see also Fensch and Bournaud 2021).
+Regardless whether clumps migrate to the centers of their hosts, stars formed within the clumps could also contribute to the growth of the thick disk (Noguchi 1996; Bournaud et al.
+2007; Inoue and Saitoh 2014).
 The evolution of the fraction of clumpy galaxies ($f_{\rm{clumpy}}$) with redshift may also provide clues to the physical processes through which clumps form and evolve (Guo et al., 2015; Shibuya et al., 2016).
@@
 At $z\leq 3$, $f_{\rm{clumpy}}$ peaks at $\sim 60\%$ at $z\sim 2$ and declines at lower and higher redshifts (Elmegreen et al., 2007; Overzier et al., 2009; Puech, 2010; Guo et al., 2012; Murata et al., 2014; Guo et al., 2015; Hinojosa-Goñi et al., 2016; Shibuya et al., 2016; Adams et al., 2022; Sok et al., 2022; Martin et al., 2023; Sattari et al., 2023).
-These studies are largely consistent with predictions from simulations of the redshift evolution of galaxies undergoing VDI, based on simulations (Cacciato et al., 2012).
+These studies are largely consistent with predictions from simulations of the redshift evolution of galaxies undergoing VDI (Cacciato et al., 2012).
 Support for mergers was also found at lower redshift (Puech, 2010) and lower masses (Guo et al., 2015).
-At $z>3$, there is disagreement on the behavior of $f_{\rm{clumpy}}$ with redshift.
-One group of studies found that $f_{\rm{clumpy}}$ decreases from $\sim 40\%$ at $z\sim 3$ to $\sim 20\%$ at $z\sim 8$ (Ravindranath et al., 2006; Conselice & Arnold, 2009; Oesch et al., 2010; Shibuya et al., 2016).
-Another found that $f_{\rm{clumpy}}$ is roughly constant with redshift at $\sim 40\%$ from $z\sim 4$ to $z\sim 8$ (Jiang et al., 2013; Kawamata et al., 2015; Curtis-Lake et al., 2016; Bowler et al., 2017; Ribeiro et al., 2017).
-Studies using data from the James Webb Space Telescope (JWST) have revealed a more complex picture of the clumps.
+At $z>3$, there is disagreement on the behavior of $f_{\rm{clumpy}}$ with redshift, with one set of observations finding that $f_{\rm{clumpy}}$ decreases from $\sim 40\%$ at $z\sim 3$ to $\sim 20\%$ at $z\sim 8$ (Ravindranath et al., 2006; Conselice and Arnold, 2009; Oesch et al., 2010; Shibuya et al., 2016), and another finding $f_{\rm{clumpy}}$ to be roughly constant with redshift at $\sim 40\%$ from $z\sim 4$ to $z\sim 8$ (Jiang et al., 2013; Kawamata et al., 2015; Curtis-Lake et al., 2016; Bowler et al., 2017; Ribeiro et al., 2017).
+Studies using data from the James Webb Space Telescope (JWST) have revealed a more complex picture.
 Thanks to the higher angular resolution, greater sensitivity, and longer wavelength coverage of the JWST, sub-kpc substructures in faint galaxies at high redshift can now be studied in greater detail with its Near-Infrared Camera (NIRCam).
@@
 On average, 76% of the rest-frame UV light from $z\simeq 6-8$ galaxies was found to come from the clumps (Chen et al., 2023), which is much higher than the $\sim 20-30\%$ measured over $0.5<z<3$ (Guo et al., 2015).
-The rest-optical and rest-NIR properties of clumps have been analyzed out to higher redshifts ($z\sim 2$) than what is possible with HST (Kalita et al., 2024, 2025; Mercier et al., 2025).
-Massive ($\log\left(M_{\star}/M_{\odot}\right)\geq 10$) star-forming galaxies at $z<3$ were found to have higher $f_{\rm{clumpy}}$ when undergoing starburst activity (Faisst et al., 2025).
 Finally, about 70% of spectroscopically selected UV-bright galaxies at $z\sim 7$ contain clumps (Harikane et al., 2025), in contrast with the 102010-20% inferred from pre-JWST studies (Shibuya et al., 2016, and references therein).
@@
 That sample was selected to be intrinsically bright with absolute UV magnitudes $23<M_{UV}<-19$ mag.
+Claeyssens et al.
+(2025) examined the physical properties of 1,751 clumps in 377 galaxies over $0.7<z<5.5$ that are gravitationally lensed by the cluster Abell 2744.
+A measurement of $f_{\rm{clumpy}}$ as a function of redshift was not performed in that work, due to the intrinsic incompleteness of studies on lensed galaxies.
 Tan et al.
-(2024) measured $f_{\rm{clumpy}}$ over $0.25<z<5$ for a sample of 909 galaxies expected to be progenitors of Milky Way analogs.
+(2025) measured $f_{\rm{clumpy}}$ over $0.25<z<5$ for a sample of 877 galaxies expected to be progenitors of Milky Way analogs.
 Their sample was selected with bins in stellar mass that evolve with redshift.
 This complicates comparisons with pre-JWST studies, which largely selected galaxies with a mass range that was constant with redshift.
-Claeyssens et al.
-(2025) examined the physical properties of 1,751 clumps in 377 galaxies over $0.7<z<5.5$ that are gravitationally lensed by the cluster Abell 2744.
-A measurement of $f_{\rm{clumpy}}$ as a function of redshift was not performed in that work.
-In this paper, we measure $f_{\rm{clumpy}}$ over $2<z<12$ with JWST data for a sample of 9,121 star-forming galaxies.
+Sok et al.
+(2025) measured $f_{\rm{clumpy}}$  and stellar masses of clumps in $\approx 4,500$ lensed galaxies with magnifications $\mu\leq 2$.
+In this paper, we measure $f_{\rm{clumpy}}$ over 2<z<9{2<z<9} with JWST data for a sample of 9,121 star-forming galaxies.
 The selection of the sample is performed such that results from the JWST and HST data may be directly compared.
-This allows us to determine whether the redshift evolution of $f_{\rm{clumpy}}$ is significantly different with JWST data compared with HST data.
+This allows us to determine whether the redshift evolution of $f_{\rm{clumpy}}$ is significantly different between HST and JWST data.
 The paper is organized as follows.
-In Section 2, we describe the data used in this work.
-The selection criteria of our sample are given in Section 3.
-Section 4.1 describes the completeness of the routine used to identify clumps.
-The redshift evolution of $f_{\rm{clumpy}}$ is discussed in Section 5.
-The contribution of clumps to the rest-frame NUV luminosity is shown in Section 6.
-In Section 7, we examine the correlation between stellar mass and sSFR with $f_{\rm{clumpy}}$.
-Predictions from simulations of clumpy galaxies are compared with our observations in Section 8.
-We discuss the implications for clump formation mechanisms in Section 9.
-We summarize and conclude in Section 10.
+In Section II, we describe the data used in this work.
+The selection criteria of our sample are given in Section III.
+Section IV describes the routine used to identify clumpy galaxies and its completeness as a function of redshift and stellar mass.
+The redshift evolution of $f_{\rm{clumpy}}$ is discussed in Section V.
+In Section VI, we examine the correlation between stellar mass with $f_{\rm{clumpy}}$.
+Predictions from simulations of clumpy galaxies are compared with our observations in Section VII.
+We discuss the implications for clump formation mechanisms in Section VIII.
+We summarize and conclude in Section IX.
 Throughout this work, cosmological parameters measured by Planck Collaboration et al.
 (2016) are adopted, with $H_{0}=67.7~\rm{km}~\rm{s}^{-1}~\rm{kpc}^{-1},\Omega_{\rm{m}}=0.309$, and $\Omega_{\Lambda}=0.691$.
-All magnitudes are in the AB photometric system (Oke & Gunn, 1983).
-2 Data
-2.1 Images, Source Detection, and Photometry
+All magnitudes are in the AB photometric system (Oke and Gunn, 1983).
+II Data
+II.1 Images, Source Detection, and Photometry
 […]
-Rieke et al.
-(2023b) detected sources in JADES and performed photometry in nine NIRCam and 14 ground- and space-based bandpasses.
-Briefly, a detection image was created by stacking the F277W, F335M, F356W, F410M, and F [^4] W images and weighting them by their inverse variances.
+Source detection and photometry in nine NIRCam and 14 ground- and space-based bandpasses in JADES was described in Rieke et al.
+(2023a).
+Briefly, a detection image was created by convolving the F277W, F335M, F356W, F410M, and F [^4] W images to the resolution of the F [^4] W image, stacking them, and weighting them by their inverse variances.
 Next, various source detection routines in the python module Photutils (Bradley, 2023) were used to detect and deblend sources and perform photometry.
 All images are aligned to the same reference frame and drizzled to a common pixel scale of 0.″03 per pixel.
-Source detection and photometry are performed on HST and JWST images convolved to the F [^4] W resolution.
+Photometry was then performed on HST and JWST images convolved to the F [^4] W resolution.
 For more details, see Section 4 in Rieke et al.
-(2023b).
+(2023a).
 This work relies on photometry in HST and JWST bandpasses.
@@
 The former are measured in apertures of various sizes that are centered on empty regions of the sky (see Section 4.2 in Rieke et al.
-2023b, also Labbé et al.
+2023a, also Labbé et al.
 2005; Whitaker et al.
@@
 Axis ratios, half-light radii, segmentation maps, and central coordinates of galaxies were measured from the detection images by Rieke et al.
-(2023b).
+(2023a).
 These measurements and data products are adopted for this analysis.
@@
 Central coordinates were computed by taking the light-weighted centroids along the horizontal and vertical axes of the detection images.
-2.2 Photometric and Spectroscopic Redshifts
+II.2 Photometric and Spectroscopic Redshifts
 Photometric redshifts in JADES come from Hainline et al.
@@
 Figure 1: Comparison of 5,780 spectroscopic and photometric redshifts in the JADES fields.
-The former are compiled from various sources (see Section 2.2) and the latter are measured using EAZY (Brammer et al., 2008).
+The former are compiled from various sources (see Section II.2) and the latter are measured using EAZY (Brammer et al., 2008).
 The bias, normalized median absolute deviation (NMAD) scatter, and outlier fraction are measured according to Dahlen et al.
@@
 The accuracy of the photometric redshifts measured by the JADES team is assessed by comparing them with spectroscopic redshifts.
-We find 2,774 galaxies with spectroscopic redshifts in GOODS-N and 3,006 in GOODS-S.
+We find 2,774 (3006) with spectroscopic redshifts in GOODS-N (GOODS-S).
 Their redshifts span $0<z<13.35$.
@@
 For the 2,591 highest-quality redshifts in GOODS-N, an outlier fraction of 13.62% was measured.
-Figure 2: The relation between stellar mass and SFR, also known as the star-formation main sequence (SFMS), for the full sample (Section 3.1) over $2<z<12$.
-These quantities are measured from fits to the combined HST and JWST SEDs using bagpipes (Carnall et al., 2018, see Section 2.3).
-Measurements for individual galaxies are shown as gray dots in various redshift bins, a different one of which is in each column.
-Completeness limits in stellar mass are indicated with vertical dashed lines.
+Figure 2: The relation between stellar mass and SFR, also known as the star-formation main sequence (SFMS), for the sample of galaxies (Section III) over $2<z<12$.
+These quantities are measured from fits to the combined HST and JWST SEDs using bagpipes (Carnall et al., 2018, see Section II.3).
+Measurements for individual galaxies are shown as gray dots in various redshift bins, a different one of which is in each panel.
+Completeness limits in stellar mass in each redshift bin are indicated with vertical dashed lines.
 There is broad agreement between the SFMS presented here and that measured in other studies.
 The SFMS measured by Popesso et al.
-(2023) over $2<z<6$ from a compilation of studies is shown as a blue solid line in the leftmost four columns.
+(2023) over $2<z<6$ from a compilation of studies is shown as a blue solid line in the leftmost four panels.
 That measured by Salmon et al.
@@
 (2024) over $6<z<12$ are shown as magenta dash-dotted lines.
-2.3 Stellar Masses and Star-Formation Rates
-Stellar masses and star-formation rates for all galaxies in the full sample (Section 3) are measured by fitting their SEDs.
-The SFRs used in this work are averaged over the last 100 Myr.
-Where available, fits were performed to the HST/Advanced Camera for Surveys (ACS) bandpasses F435W, F606W, and F775W combined with all the NIRCam bandpasses listed in Section 2.1.
+II.3 Stellar Masses and Star-Formation Rates
+Stellar masses and star-formation rates for all galaxies in the full sample (Section III) are measured by fitting their SEDs.
+The SFRs used in this work are taken to be the average values inferred from the most recent 100 Myr in the fit star-formation histories (SFHs).
+Where available, fits were performed to the HST/Advanced Camera for Surveys (ACS) bandpasses F435W, F606W, and F775W combined with all the NIRCam bandpasses listed in Section II.1.
 The number of filters used in each fit ranges from five to 17, with a median of 13.
@@
 We set a uniform prior on it over the range [0, 5].
-Another prior is assumed for the differential attenuation between young ($\leq 10$ Myr) and old stars (see, e.g., Calzetti et al., 2000; Charlot & Fall, 2000).
+Another prior is assumed for the differential attenuation between young ($\leq 10$ Myr) and old stars (see, e.g., Calzetti et al., 2000; Charlot and Fall, 2000).
 The former are generally dustier than the latter due to additional attenuation from birth clouds.
@@
 The latter is assumed to be the same as that of the stars.
-Last, a double power-law star-formation history is assumed. It is given by
+Last, a double power-law SFH is assumed, given by
 $\textrm{SFR}(t)\propto\left[\left(\frac{t}{\tau}\right)^{\alpha}+\left(\frac{t}{\tau}\right)^{-\beta}\right]^{-1},$ (3)
@@
 This is relevant only for the HST and bluest NIRCam filters, especially at high redshift.
-Stellar population synthesis templates from version 2.2.1 of the Binary Population and Spectral Synthesis (BPASS; Eldridge et al., 2017; Stanway & Eldridge, 2018) models are used.
-The initial mass function is set to the default ‘135_300’ model listed in Table 2 in Stanway & Eldridge (2018).
+Stellar population synthesis templates from version 2.2.1 of the Binary Population and Spectral Synthesis (BPASS; Eldridge et al., 2017; Stanway and Eldridge, 2018) models are used.
+The initial mass function is set to the default ‘135_300’ model listed in Table 2 in Stanway and Eldridge (2018).
 The cloudy (Ferland et al., 2017) photoionization code is used to model nebular emission.
-All HST and JWST filter transmission curves, which are used in the SED fits, are taken from the SVO Filter Profile Service (Rodrigo et al., 2012; Rodrigo & Solano, 2020).
+All HST and JWST filter transmission curves, which are used in the SED fits, are taken from the SVO Filter Profile Service (Rodrigo et al., 2012; Rodrigo and Solano, 2020).
 In Figure 2, the relation between stellar mass and SFR, also known as the star formation main sequence (SFMS; e.g., Noeske et al., 2007; Rodighiero et al., 2010; Whitaker et al., 2014; Salmon et al., 2015; Schreiber et al., 2015; Tomczak et al., 2016; Popesso et al., 2023; Ciesla et al., 2024; Cole et al., 2025) is shown.
@@
 The distribution of stellar masses and SFRs measured in this work agrees broadly with the SFMS measured in various studies.
-2.4 Stellar Mass Completeness Limits
+II.4 Stellar Mass Completeness Limits
 Here we compute 90% mass completeness limit for our selected sample.
@@
 The stellar masses obtained with bagpipes are used.
-The completness limits are estimated as follows. First, a limiting stellar mass, $M_{\rm{lim}}$, is computed for all galaxies in the full sample (Section 3). It is defined to be
+The completness limits are estimated as follows. First, a limiting stellar mass, $M_{\rm{lim}}$, is computed for all galaxies in the full sample (Section III). It is defined to be
 $\log\left(M_{\rm{lim}}\right)=\log\left(M_{\star}\right)+0.4\left(m-m_{\rm{lim}}\right),$ (4)
@@
 In this work, the magnitude in the F [^4] W band is used to calculate $M_{\rm{lim}}$.
-This band is used as it is the longest NIRCam filter available in both GOODS-N and GOODS-S and thus best traces the stellar mass.
-After finding $M_{\rm{lim}}$ for all galaxies, they are divided into redshift bins (see Section 3.1).
+This band is used as it is the longest-wavelength NIRCam filter available in both GOODS-N and GOODS-S and thus best traces the stellar mass.
+After finding $M_{\rm{lim}}$ for all galaxies, we divide them in redshift bins (Section III).
 In each bin, the faintest 20 percent of galaxies are identified.
@@
 […]
-3 Sample Selection
-3.1 The Full Sample
-We start with all sources in the JADES GOODS-N and JADES GOODS-S photometric catalogs. There are 179,709 sources in both fields: 85,709 in GOODS-N and 94,000 in GOODS-S.
+III Selection of the Sample of Galaxies
+We start with all sources in the JADES GOODS-N and GOODS-S photometric catalogs. There are 179,709 sources in both fields: 85,709 in GOODS-N and 94,000 in GOODS-S.
 All sources that are brighter than or equal to 29th magnitude in the reddest broadband filter, F [^4] W, are selected.
-This is about 0.5 mag brighter than the limiting magnitudes for all NIRCam filters in JADES (Table 2 in Rieke et al., 2023b).
+This is about 0.5 mag brighter than the limiting magnitudes for all NIRCam filters in JADES (Table 2 in Rieke et al., 2023a).
 This cut removes 82,606 objects with 97,103 remaining.
@@
 […]
-Cuts are then applied on the morphologies of galaxies.
-These are done to exclude compact and edge-on galaxies and have often been applied in pre-JWST studies of clumpy galaxies (e.g., Guo et al., 2012, 2015; Shibuya et al., 2016).
+Cuts are then applied to exclude compact and edge-on galaxies, following pre-JWST studies of clumpy galaxies (e.g., Guo et al., 2012, 2015; Shibuya et al., 2016; Sattari et al., 2023).
 Galaxies with circular half-light radii of at least 0.″1 in the F [^4] W bandpass are selected.
+This minimum size is slightly larger than the half-width at half maximum of the F [^4] W point spread function (PSF, Zhuang and Shen 2024) and is imposed in order to select galaxies that are resolved.
 This removes 2,848 galaxies with 16,345 remaining.
 Galaxies with axis ratios of $b/a\geq 0.5$ are chosen.
-This excludes 933 more.
+This excludes 933 more galaxies.
 The full sample contains 15,412 galaxies: 7,167 in GOODS-N and 8,245 in GOODS-S.
-Galaxies in the final sample have a median S/N of about 4.0 in F090W and 13.2 in F277W.
-For each galaxy in the full sample, its stellar mass and SFR are measured by fitting its combined HST+JWST SED (Section 2.3).
+Galaxies in the full sample have a median S/N of about 4.0 in F090W and 13.2 in F277W.
+For each galaxy in the sample, its stellar mass and SFR are measured by fitting its combined HST+JWST SED (Section II.3).
 About 8% (1,195/15,412) of the full sample has spectroscopic redshifts.
 The distribution of the full sample in terms of redshift and stellar mass is shown in Figure 4.
-Individual galaxies are represented with small gray dots.
-The stellar mass completeness limit for each redshift bin is indicated with a light gray shaded region.
-Galaxies with spectroscopic redshifts are indicated with empty red squares.
-3.2 The Final Sample
-Galaxies with stellar masses $\geq 10^{8}~M_{\odot}$ are selected.
-This is based on our completeness limits, details which are provided in Section 2.4.
+To identify clumps within a mass-complete sample, galaxies with stellar masses $\geq 10^{8}~M_{\odot}$ are selected.
+This is based on our completeness limits (Section II.4).
 In the highest redshift bin, $9<z<12$, this cut in mass is about 80% complete.
@@
 Applying the mass cut removes 6,074 galaxies and leaves 9,338.
-Figure 4: The stellar masses as a function of redshift are shown for the full sample (Section 3).
+Figure 4: The stellar masses as a function of redshift are shown for the full sample (Section III).
 Individual galaxies with only photometric redshifts are plotted as small gray dots and those with spectroscopic redshifts are indicated with red squares.
@@
 Our final sample contains 9,121 star-forming galaxies.
-Figure 5: The completeness of our clump detection algorithm in redshift bins (columns) and stellar mass bins of their host galaxies (thickness of the lines) is shown here.
-Fake clumps are added to rest-frame NUV images of real galaxies.
-More than 1,500 galaxies are randomly selected over redshifts $2<z<12$ and stellar masses $8<\log\left(M_{\star}/M_{\odot}\right)<11$, regardless whether they have clumps.
-The completeness is defined as the fraction of fake clumps recovered relative to the number inserted into each galaxy image.
-The relation between the ratio of fake clump luminosity to host galaxy luminosity and completeness is used to determine minimum fractional luminosities used to find clumps.
+Figure 5: Example clumpy galaxies are shown in each redshift (vertical) and stellar mass bin (horizontal), where a complete sample of clumps is detected.
+For each galaxy, images in the rest-frame NUV are shown to the left and an RGB image is on the right.
+Unmodified images are shown to the top left; smoothed images on the top right; contrast images on the bottom left, and detected clumps are circled in pink in the bottom right.
+Clumps that meet the completeness limits listed in Table 1 are denoted with solid lines, otherwise, with dotted lines.
+The segmentation map of each galaxy is indicated with the red outline.
+Figure 6: Surface brightness profiles of detected clumps.
+In each panel, the surface brightness profiles of individual clumps are shown as thin gray lines.
+Those of the PSFs of the filters used to detect clumps are shown with dark blue dotted lines.
+Median profiles are plotted as thick, black, solid lines and running 16th and 84th percentiles are shown as lower and upper dashed lines.
+Sérsic profiles are fit to the median profiles and shown as red dash-dot lines.
+The effective radii and indices of the fits are written in red text.
+All surface brightness profiles are normalized to that of the innermost aperture used to measure them.
+Panels are arranged in blocks of 3x3 panels that are indicated with bins in redshift.
+Only panels with at least 5 detected clumps are shown.
+Within each block, host galaxy stellar mass increases from left to right and brightness of the clump compared with its host increases from top to bottom.
+The latter quantity is measured by assuming the clumps behave as point sources.
+In all panels, clumps are more extended than the PSFs.
+The angular sizes of clumps are nearly constant with redshift, but their physical sizes increase by a factor of $\sim 2$ from $z\sim 7.75$ to $z\sim 2.75$.
+IV Selection of Clumpy Galaxies
+IV.1 Identification of Clumps
+Clumps are identified for all galaxies in the mass-selected, star-forming sample as follows.
+For each galaxy in the sample, a 6″ by 6″ postage stamp is created in the bandpass corresponding to the rest-NUV.
+This image is then smoothed with a boxcar filter of width equal to 10 pixels.
+Our choice of filter width is justified with injection-recovery tests (Section IV.3).
+The smoothed image is subtracted from the original image to create a contrast image.
+Objects are then detected in the contrast image.
+This is done with the photutils library.
+The $3\sigma$-clipped standard deviation of the background is measured and source detection is performed.
+Deblending is done with photutils to separate overlapping clumps, with a minimum contrast of 0.001 and 32 thresholds.
+Objects are selected to be candidate clumps if they span at least five contiguous pixels, each of which must be at least two standard deviations brighter than the background-subtracted image, and if the total S/N of the object is at least 3.
+The minimum physical area of a candidate clump is 0.33 kpc2 at $z=2$ and 0.06 kpc2 at $z=12$.
+To be selected as clumps, candidates must lie at least 0.″1 from the center and within the segmentation map created by the JADES team.
+This corresponds to a distance of 0.86 kpc at $z=2$ and 0.37 kpc at $z=12$.
+The former criterion is used to avoid confusion with a bulge or active galactic nucleus and the latter ensures that all detected clumps lie within a single galaxy and not within its neighbors projected on the sky.
+Examples of the detection technique and identified clumps are shown in Figure 5.
+The technique and selection criteria above broadly follow many pre-JWST studies (Murata et al., 2014; Guo et al., 2015; Shibuya et al., 2016; Soto et al., 2017; Martin et al., 2023; Sattari et al., 2023).
+IV.2 Radial Surface Brightness Profiles and Aperture Photometry of Clumps
+In the HST data, clumps were largely found to be point sources (Guo et al., 2015; Shibuya et al., 2016; Sattari et al., 2023).
+Given the higher spatial resolution of JWST, this assumption may no longer hold.
+The surface brightness profiles of clumps are shown in Figure 6, divided into bins of redshift, host galaxy stellar mass, and UV brightness of the clumps relative to their host galaxies.
+Profiles were measured with concentric circular apertures centered on the centroids determined by our clump detection routine.
+Clumps detected in this work have spatially extended surface brightness profiles, which are inconsistent with point sources.
+Empirical PSFs are measured and described in Appendix A.
+Figure 6 shows that in all bins, the median profiles are more extended than those of the PSFs.
+To measure the flux of each clump, a typical surface brightness profile shape is first assumed.
+Then, the flux is directly measured within a small circular aperture.
+Background flux from the host galaxy is removed and an aperture correction is performed.
+This follows previous works (e.g., Guo et al., 2015; Shibuya et al., 2016; Sattari et al., 2023), except for our treatment of the clumps as extended sources.
+A Sérsic function is assumed for clump profiles, which is then convolved with the PSF and fit to the median observed profiles (Figure 6).
+The effective radii and Sérsic indices of the fits are estimated.
+Relative errors on the effective radii are generally small, of only a few percent.
+When fitting the median profiles, the background light from the host is not subtracted, but is included in the fit.
+We assume the light profile over 6106-10 pixels in radius is dominated by the background light, as most profiles flatten out at these radii (see also Guo et al.
+2015).
+There is little variation in the median clump profile with redshift, host galaxy mass, or clump brightness.
+Over all bins shown in Figure 6, the median effective radius and Sérsic index from the fits are $R_{\rm{eff}}=$0.″046 and $n=1.04$.
+The 16–84th percentile range in $R_{\rm{eff}}$ is 0.″044 to 0.″049 and that for $n$ is 0.94 to 1.10.
+There is a weak evolution in physical size with redshift, such that the median effective radius of clumps grows from 0.23 kpc (0.″045) at $z\sim 7.75$ to 0.39 kpc (0.″048) at $z\sim 2.75$.
+This is a factor of a few larger than the median size of clumps within lensed galaxies, which was found to be 0.093 kpc over $3.5<z<5.5$ (Claeyssens et al., 2025).
+The larger sizes reported in this work may be due to our selection of unlensed galaxies.
+Individual clumps that may only be detectable in lensed galaxies could aggregate into larger ones in unlensed systems.
+Aperture photometry of the clumps is performed with photutils as follows.
+First, the flux of a given clump is measured with a circular aperture of radius 4 pixels.
+Beyond this radius, the typical clump light profile is dominated by host galaxy light (see Figure 6).
+The background light is subtracted by measuring the flux in circular annuli over the radii 6-10 pixels.
+It is then divided by the area of the annulus to yield a mean surface brightness.
+This is multiplied by the area of the inner aperture with radius of 4 pixels to yield a background flux, which is then subtracted from the original flux measured.
+In each step, all other clumps are masked with circular masks of radius 4 pixels.
+The sizes of the apertures, annuli, and masks used here are also the same as those used in Guo et al.
+(2015).
+Last, aperture corrections are applied to background-subtracted fluxes of clumps.
+These are calculated by assuming clump light profiles broadly follow the median Sérsic profile $(R_{\rm{eff}}=$0.″046, $n=1.04)$.
+This profile is then convolved with a PSF.
+The total flux of the convolved profile and that within a circular aperture of radius equal to 4 pixels are measured.
+The aperture correction is taken to be the ratio of the total flux to that within the aperture.
+Clump aperture fluxes are multiplied by factors of 1.511, 1.473, 1.477, 1.514, and 1.784 in GOODS-N and 1.533, 1.492, 1.509, 1.530, and 1.738 in GOODS-S for the F090W, F115W, F150W, F200W, and F277W bandpasses, respectively.
+These are determined using the empirical PSFs (Appendix A).
+Figure 7: Clump completeness functions in redshift (vertical) and stellar mass (horizontal) bins.
+The smoothing length used to detect clumps are also shown.
+Completeness is measured by injecting fake clumps to rest-frame NUV images of real galaxies and finding the fraction of fake clumps recovered relative to the number inserted into each galaxy image.
+The completeness is therefore measured as the ratio of fake clump luminosity to host galaxy luminosity above which >>50% of the fake clumps are detected.
 A completeness of 50% is denoted with the horizontal dashed lines.
@@
 Completeness curves are shown only for ranges in mass that include at least 10 galaxies.
-Faint clumps are more likely to be detected in higher-mass galaxies at lower redshifts.
-For galaxies over $9<\log\left(M_{\star}/M_{\odot}\right)<10$, clumps are identified with 50% completeness at 2% of the host luminosity at $z\sim 2.75$ and at 7% of the host luminosity at $z\sim 7.75$.
-At $z\sim 2.75$ and $\log\left(M_{\star}/M_{\odot}\right)\geq 9$, clumps can be identified at nearly four times fainter luminosities with JWST than with HST.
-The error on the completeness is calculated using Bayesian binomial confidence intervals and shown as the light shading around each line.
-Table 1: Minimum Fractional Luminosities Needed to Identify $\geq 50\%$ of Clumps as a Function of Stellar Mass and Redshift
-Redshift Range $8<\log\left(M_{\star}/M_{\odot}\right)<9$ $9<\log\left(M_{\star}/M_{\odot}\right)<10$ $10<\log\left(M_{\star}/M_{\odot}\right)<11$
-($L_{\rm{clump}}/L_{\rm{host~galaxy}}$) ($L_{\rm{clump}}/L_{\rm{host~galaxy}}$) ($L_{\rm{clump}}/L_{\rm{host~galaxy}}$)
-2–3.5 $\geq$7% $\geq$2% $\geq$2%
-3.5–5 $\geq$10% $\geq$2% $\geq$2%
-5–6.5 $\geq$15% $\geq$5% $\ldots$
-6.5–9 $\geq$25% $\geq$7% $\ldots$
-9–12 $\ldots$ $\ldots$ $\ldots$
-Note.
-— Each fraction is in terms of the rest-frame NUV luminosity of the host galaxy.
-Empty values indicate that either too few (<10<10) galaxies lie in a given mass and redshift bin or clumps could not be detected with at least 50% completeness.
-4 Selection of Clumpy Galaxies
-4.1 Identification of Clumps
-Clumps are identified for all galaxies in the mass-selected, star-forming sample (Section 3.2) using the technique developed by Sattari et al.
-(2023).
-Briefly, it involves taking the Fourier Transform of the rest-frame NUV image of a galaxy, filtering out the low-frequency signal that corresponds to the host galaxy light, and performing source detection on the leftover high-frequency signal, which consists of clumps and noise (see Sec.
-3 in Sattari et al.
-2023 for details).
-To distinguish these, we apply stringent source detection criteria.
-Objects are identified as clumps in JWST/NIRCam images if they span at least five contiguous pixels, each of which must be at least two standard deviations brighter than the background-subtracted image, and if the total S/N of the object is at least 3.
-These criteria are the same as those used by Shibuya et al.
-(2016), who used HST data, and similar to those used by Claeyssens et al.
-(2025), who used JWST data.
-We use photutils to detect clumps.
-For a given galaxy, clumps must lie at least 0.″1 from the center and within the segmentation map created by the JADES team.
-This corresponds to a distance of 0.86 kpc at $z=2$ and 0.37 kpc at $z=12$.
-The former criterion is used to avoid confusion with a bulge or active galactic nucleus and the latter ensures that all detected clumps lie within a single galaxy and not within its neighbors projected on the sky.
-Figure 6: Example clumpy galaxies are shown in this figure for each of the stellar mass and redshift bins considered in this work (Figure 5 and Section 4.2).
-Stellar mass bins are indicated with the columns and increase from left to right.
-Redshift bins are denoted with the rows and increase from top to bottom.
-For each galaxy in a given mass and redshift bin, images in the rest-frame NUV are displayed in the left two columns and a RGB image is shown in the rightmost column.
-Images are shown with an arcsinh stretch.
-In the middle columns, detected clumps are circled in pink.
-Clumps that meet the completeness limits listed in Table 1 are denoted with solid lines and those that do not with dotted lines.
-The segmentation map of each galaxy is indicated with the red outline.
-The bandpasses used to create the RGB images are listed on the far left side of the figure.
-4.2 Completeness of Clump Identification
-To ensure that a complete sample of clumps is identified, we measure how faint clumps can be in order to detect at least half of them, following pre-JWST studies (Guo et al., 2015; Shibuya et al., 2016; Sattari et al., 2023).
-This is done by adding fake clumps to images of randomly selected real galaxies, regardless whether they have clumps, and running the clump identification method described above.
-Fake clumps are modeled as point sources following previous works (e.g., Guo et al., 2012, 2015, 2018; Shibuya et al., 2016; Soto et al., 2017; Sattari et al., 2023).
-We measure empirical point spread functions (PSFs) for all NIRCam bandpasses in GOODS-N and GOODS-S.
-This is described in Appendix A.
+The error on the completeness is shown as the light shading around each line.
+IV.3 Completeness of Clump Identification
+To ensure that a complete sample of clumps is identified, we measure how faint clumps can be in order for half of them to be detected (Guo et al., 2015; Shibuya et al., 2016; Sattari et al., 2023).
+This is done by adding fake clumps to images of randomly selected real galaxies, regardless whether they have clumps or not, and running the clump identification method described above.
+Fake clumps are modeled as Sérsic profiles with the same effective radius and index as the median profile of the clumps.
+They are convolved with the empirical PSFs measured in this work.
+The PSF used to convolve the profile varies with redshift as described in Section III.
 Positions of the fake clumps are randomly selected to lie within the segmentation map of a given galaxy.
@@
 In total, 1,541 galaxies are randomly selected.
-A fake clump is added to the rest-frame NUV image of each galaxy using the ImagePSF routine in photutils.
+A fake clump is added to the rest-frame NUV image of each galaxy.
 The clump identification algorithm is then run on the modified image to determine whether the fake clump is recovered.
@@
 Uncertainties on the completeness are computed in a Bayesian fashion and assume a noninformative Jeffreys prior on the binomial distribution (Cameron, 2011).
-The completeness of the fake clump simulations is shown in Figure 5.
+Since our clump detection procedure depends on the choice the smoothing length of the boxcar filter, we perform the above procedure for three choices of smoothing length: 5, 10, and 20 pixels (0.″15, 0.″30, and 0.″60, respectively).
+If the smoothing length is too small, only the brightest clumps will be detected.
+If it is too large, spurious clumps may be detected, as noted by Guo et al.
+(2015).
+In this work, the optimal length is found to be 10 pixels (see below).
+It yields maximal completeness while being intermediate in length, thereby minimizing the chance of detecting spurious clumps.
+The clump completeness functions are shown in Figure 7.
 In each panel, the fraction of clumps detected by the clump identification algorithm is shown as a function of the fake clump luminosity.
 This is further shown as a function of stellar mass (lines of various thicknesses) and redshift (columns).
+In each row, completeness curves are shown for a certain smoothing length that is used to detect clumps.
 Clumps are more likely to be detected in higher-mass systems and lower redshifts.
-The minimum fractional luminosities at which clumps are detected at 50% completeness are indicated with the vertical dashed-dotted lines in Figure 5.
-For galaxies at $z\sim 2.75$ and over $9<\log\left(M_{\star}/M_{\odot}\right)<10$, 50% of clumps are detected when they are at least 2% of the NUV luminosities of their hosts.
-At $z\sim 7.75$ and over the same range in mass, the same completeness is obtained for a minimum fractional luminosity of 7%.
-At lower masses, minimum fractional luminosities are higher, and at higher masses, they are similar.
-At $z\sim 10.5$, no clumps are detected with at least 50% completeness.
-Because of this, galaxies in the two highest redshift bins at $z\sim 7.75$ and $z\sim 10.5$ will be combined into one redshift bin in the following Sections.
-The minimum fractional luminosities at $z\sim 7.75$ will be used to find clumps over $6.5<z<12$.
-Table 1 lists the minimum fractional luminosities as a function of stellar mass and redshift.
-At $z\sim 2.75$ and $\log\left(M_{\star}/M_{\odot}\right)>9$, clump identification is 50% complete at four times fainter luminosities with JWST (2%) than with HST (8%; Shibuya et al.
+Completeness limits are generally the highest for the shortest smoothing length of 5 pixels.
+They are smallest at the intermediate length of 10 pixels and change little for the longest smoothing length of 20 pixels.
+The intermediate length is chosen to be optimal.
+For all lengths considered, no clumps at $z\sim 10.5$ are detected with at least 50% completeness.
+In the following Sections, we will only focus on the redshift range of $2<z<9$.
+Table 1 lists the minimum fractional luminosities as a function of stellar mass and redshift for the optimal smoothing length of 10 pixels.
+At $z\sim 2.75$ and $\log\left(M_{\star}/M_{\odot}\right)>9$, clump identification is 50% complete at about two times fainter luminosities with JWST (3.5%) than with HST (8%; Shibuya et al.
 2016).
@@
 These depend on stellar mass and redshift.
-Example clumpy galaxies are shown in Figure 6.
-Clumps that are brighter than the completeness limits are indicated with solid circles and those that are fainter with dotted circles.
-Figure 7: The fraction of clumpy galaxies, $f_{\rm{clumpy}}$, over $2<z<12$.
-It peaks at $z\sim 2.75$ and declines by a factor of $\sim 4$ from $z\sim 2.75$ to $z>6.5$ at a nearly constant rate.
+Figure 8: The fraction of clumpy galaxies, $f_{\rm{clumpy}}$, over 2<z<9{2<z<9}.
+It is highest at $z\sim 2.75$ and declines by a factor of $\sim 3$ from $z\sim 2.75$ to $z\sim 5.75$.
 Galaxies are considered clumpy if they have at least one off-center clump.
 Measurements from this work for galaxies with stellar masses of at least $10^{9}~M_{\odot}$ are shown as black hexagons connected by a solid black line.
-$f_{\rm{clumpy}}$ measurements for galaxies in GOODS-N are shown to the left of the solid black line.
-Those for galaxies in GOODS-S are to the right of it.
+$f_{\rm{clumpy}}$ measurements for galaxies in GOODS-N (GOODS-S) are shown to the left (right) of the solid black line.
 The minimum fractional luminosity changes with redshift according to Table 1.
 We compile measurements from the literature and plot them alongside ours.
-These are shown as various symbols and lines and are listed in Table 2 (see Section 5 for details).
-The results are consistent: $f_{\rm{clumpy}}$ rises gradually from $z\sim 8$ to $z\sim 1.5-2$, when it peaks, and declines at lower redshifts.
+The results from these studies are largely consistent with each other: $f_{\rm{clumpy}}$ rises gradually from $z\sim 8$ to $z\sim 1.5-2$, when it peaks, and declines at lower redshifts.
 This is qualitatively consistent with the evolution of the cosmic SFR density, which is shown as a thin blue solid line and has been arbitrarily scaled vertically.
 The results of this study are consistent with the view that clumpy galaxies became more common at later times.
-However, the $f_{\rm{clumpy}}$ calculated in this work is higher at fixed redshift and increases more steeply with decreasing redshift than the bulk of previous studies.
-This suggests that clumpy galaxies were more common at high redshift than previously measured, but with consistent trends towards lower redshifts.
-4.3 Aperture Photometry of Clumps
-We measure the fluxes of detected clumps by performing fixed aperture photometry with photutils.
-Each aperture is circular with a radius of 3 pixels.
-The background is measured in annuli with inner radii of 4 pixels and outer radii of 10 pixels.
-To perform background subtraction, the mean surface brightness of the light in the annuli is measured, multiplied by the area of the aperture, and subtracted.
-In each of these steps, for a given galaxy, any nearby clumps are masked using a circular mask of radius equal to 4 pixels.
-The sizes of the apertures, background annuli, and masks are adopted from Guo et al.
-(2018), who carried out extensive tests on photometry of clumps in HST images.
-Because clumps are modeled as point sources, the circular apertures used here enclose the majority of the total light, about 50-70%, depending on the bandpass.
-We account for this by applying aperture corrections that vary with wavelength.
-Clump aperture fluxes are multiplied by factors of 1.434, 1.400, 1.422, 1.477, and 1.886 in GOODS-N and 1.471, 1.429, 1.469, 1.502, and 1.806 in GOODS-S for the F090W, F115W, F150W, F200W, and F277W bandpasses, respectively.
-These are determined using the empirical PSFs that are measured in this work (see Appendix A).
-5 The Fraction of Clumpy Galaxies over the Redshift Range $2<z<12$
-We now study the fraction of clumpy galaxies ($f_{\rm{clumpy}}$) as a function of redshift.
-For a galaxy to be considered clumpy, it must have at least one off-center clump identified in its rest-frame NUV image.
-This is done using the technique developed by Sattari et al.
-(2023).
-Clumps are identified using the minimum fractional luminosities listed in Table 1).
-The fraction of clumpy galaxies is defined to be the number of galaxies with clumps divided by the total number of galaxies in a given bin in stellar mass, redshift, and/or sSFR.
+However, as redshift decreases, the $f_{\rm{clumpy}}$ calculated in this work rises more steeply than that measured from the bulk of previous studies.
+At $z\sim 2.75$ and $z\sim 4.25$, our $f_{\rm{clumpy}}$ measurements are higher than those of most studies by factors of $\sim 1.3-3$.
+V Fraction of Clumpy Galaxies and Their Evolution with Redshift
+For a galaxy to be clumpy, it must have at least one off-center clump identified in its rest-frame NUV image (Section IV.1).
+Each clump must have a minimum fractional luminosity relative to its host galaxy.
+These luminosities are listed in Table 1.
+The fraction of clumpy galaxies ($f_{\rm{clumpy}}$) is defined to be the number of galaxies with clumps divided by the total number of galaxies in a given bin in redshift and stellar mass.
 All uncertainties are calculated assuming Bayesian confidence intervals on the binomial distribution, unless otherwise noted.
-The evolution of $f_{\rm{clumpy}}$ with redshift is examined in Section 5.1.
-In Section 5.2, measurements from this work are compared with those from other studies that overlap with ours in redshift.
-We find that our results disagree with those from most studies over $2<z<6.5$.
-To understand why this is, the redshift evolution of $f_{\rm{clumpy}}$ is measured from HST and JWST images and examined in Section 5.3.
-Table 2: Summary of Previous Studies on Clumpy Galaxies
+Table 1: Minimum Fractional Luminosities Needed to Identify $\geq 50\%$ of Clumps as a Function of Stellar Mass and Redshift
+Redshift Range $8<\log\left(M_{\star}/M_{\odot}\right)<9$ $9<\log\left(M_{\star}/M_{\odot}\right)<10$ $10<\log\left(M_{\star}/M_{\odot}\right)<11$
+($L_{\rm{clump}}/L_{\rm{host~galaxy}}$) ($L_{\rm{clump}}/L_{\rm{host~galaxy}}$) ($L_{\rm{clump}}/L_{\rm{host~galaxy}}$)
+2–3.5 $\geq$13% $\geq$3.5% $\geq$3.5%
+3.5–5 $\geq$25% $\geq$6% $\geq$2.5%
+5–6.5 $\geq$40% $\geq$11% $\ldots$
+6.5–9 $\geq$50% $\geq$23% $\ldots$
+9–12 $\ldots$ $\ldots$ $\ldots$
+Note.
+— Each fraction is in terms of the rest-frame NUV luminosity of the host galaxy.
+Empty values indicate that either too few (<10<10) galaxies lie in a given mass and redshift bin or clumps could not be detected with at least 50% completeness.
+All minimum fractional luminosities here are reported for the optimal smoothing length of 10 pixels.
+Table 2: Summary of Literature on Clumpy Galaxies
 Reference Sample $\left(N_{\rm{galaxy}}\right)$ Stellar Mass Range Redshift Range Method Detection Band
@@
 Elmegreen et al. (2007) Starbursts (1003) N/A 0–5 Visual F775W
-Conselice & Arnold (2009) LBGs (133) $\gtrsim 9$ 4–6 $CASGM_{20}$eefootnotemark: rest-frame UV
+Conselice and Arnold (2009) LBGs (133) $\gtrsim 9$ 4–6 $CASGM_{20}$eefootnotemark: rest-frame UV
 Oesch et al. (2010) LBGs (21) $\gtrsim 8$ 7–8 Visual rest-frame UV
@@
 Sattari et al. (2023) SFGs (6767) >9.5>9.5 0.5–3 Algorithm rest-frame 1600 Å
-Tan et al. (2024) MWAfffootnotemark: progenitors (909) 8.5–10.7 0.25–5 Algorithm rest-frame $U$ band
 Harikane et al. (2025) SFGs (12) $\gtrsim 8.5$ 6.595–7.154 Algorithm F115W
-This work SFGs (9121) $\geq 8$ 2–12 Algorithm rest-frame NUV
+This work SFGs (9121) $\geq 8$ 2– 9 Algorithm rest-frame NUV
 Note.
@@
 eS:S:” clumpiness.
-fMilky Way Analog.
 See Conselice (2014) for definitions of the $CASGM_{20}$ morphological parameters.
-5.1 Evolution with Redshift
+V.1 Evolution with Redshift
 Figure 8 shows $f_{\rm{clumpy}}$ as a function of redshift.
 Measurements from this work for galaxies with stellar masses greater than $10^{9}~M_{\odot}$ are shown as black hexagons connected by a thick solid line.
-This limit in mass is constant with redshift and chosen for two reasons.
-First, at a fixed redshift, clumps are identified using a uniform cut in luminosity at masses above this limit (see Table 1).
-Second, it is comparable to the mass limits used by other studies (see Table 2).
-The fraction of clumpy galaxies is highest ($\sim 80\%$) at $z\sim 2.75$ and decreases with increasing redshift to $\sim 20$% at $z>6.5$.
-The fraction decreases at a roughly constant rate with increasing redshift: it is $\sim 60\%$ at $z\sim 4.25$ and $\sim 40\%$ at $z\sim 5.75$.
-$f_{\rm{clumpy}}$ measurements in either GOODS-N or GOODS-S agree with this trend and differ from it by $\leq 4\%$.
-These are indicated with empty hexagons to the left and right of the solid black line in Figure 7.
-Our $f_{\rm{clumpy}}$ measurements are also compared with the cosmic SFR density (SFRD), as measured by Madau & Dickinson (2014).
-This is shown as a thin blue solid line in Figure 7.
-The fraction of clumpy galaxies is observed to rise with decreasing redshift more steeply than the SFRD.
-However, the net change in $f_{\rm{clumpy}}$ from $z>6.5$ to $z\sim 2.75$, an increase of about a factor of four, is similar to that in the SFRD over the same range in redshift.
-This seemingly suggests a connection between the frequency of clumpy galaxies and the volume-averaged SFR at high redshift.
-Figure 8: Clumpy galaxies are found to be more common over $2<z<6$ in this work compared with earlier studies based on HST due to the higher sensitivity of JWST.
-This figure shows $f_{\rm{clumpy}}$ measured with HST (dashed lines) and JWST (solid lines) as a function of redshift for three choices of the minimum fractional luminosity $(L_{\rm{clump}}/L_{\rm{host~galaxy}})$.
-This is the minimum fraction of the host galaxy NUV luminosity a clump must have.
-A slight offset in redshift is applied to each set of $f_{\rm{clumpy}}$ measurements for clarity.
-In the left panel, the limit is set to 8%, which is adopted from pre-JWST studies.
-Measurements from both telescopes agree to within 9% and are statistically consistent.
-At the intermediate limit of 5%, shown in the middle panel, fractions differ by 6–18%, with those from JWST being higher at all redshifts.
-Estimates are statistically consistent for only one redshift bin.
-Finally, at the lowest limit of 2%, shown in the right panel, JWST-based measurements are higher than those from HST by 22–29% and statistically inconsistent at all redshifts.
-5.2 Comparison with Other Studies
+This limit in mass is constant with redshift and chosen because it is comparable to the mass limits used by other studies (see Table 2).
+The fraction of clumpy galaxies is highest ($\sim 70\%$) at $z\sim 2.75$ and decreases with increasing redshift to $\sim 20$% at $z\sim 7.75$.
+It changes the most from $\sim 30\%$ at $z\sim 5.75$ to $\sim 60\%$ at $z\sim 4.25$.
+$f_{\rm{clumpy}}$ measurements in either GOODS-N or GOODS-S agree with this trend within $\leq 3\%$.
+Our $f_{\rm{clumpy}}$ measurements are also compared with the cosmic SFR density (SFRD), as measured by Madau and Dickinson (2014), in Figure 8.
+It has been rescaled along the y-axis to line up with the majority of the points in the figure.
+The fraction of clumpy galaxies rise s with decreasing redshift more steeply than the SFRD.
+However, the net change in $f_{\rm{clumpy}}$ from $z\sim 7.75$} to $z\sim 2.75$, an increase of almost a factor of four, is similar to that in the SFRD over the same range in redshift.
+This seemingly suggests a connection between the frequency of clumpy galaxies and the volume-averaged SFR as a function of redshift.
+V.2 Comparison with Other Studies
 Our $f_{\rm{clumpy}}$ measurements are also compared with those from other studies in Figure 8.
@@
 They were calculated in this work assuming Bayesian binomial confidence intervals, wherever feasible.
-Details on the extraction of $f_{\rm{clumpy}}$ measurements from earlier studies are provided in Appendix B.
-We now examine the evolution of $f_{\rm{clumpy}}$ over $0<z<8$ in the rest-frame UV from previous studies, shown in Figure 7.
-This was found to be $\sim 30-40\%$ at $z\sim 0.5$, peak at $\sim 60-80\%$ at $z\sim 2$, and decline to $\sim 40-60\%$ at $z\sim 3$ (Elmegreen et al.
-2007; Wuyts et al.
-2012; Tadaki et al.
-2014; Guo et al.
-2015; Shibuya et al.
-2016; Sattari et al.
-2023; although lower fractions of $\sim 20-40\%$ were obtained by Law et al.
+Details on the extraction of $f_{\rm{clumpy}}$ measurements from earlier studies are provided below.
+Following Shibuya et al.
+(2016), we include $f_{\rm{clumpy}}$ measurements from Lotz et al.
+(2006); Ravindranath et al.
+(2006); Conselice and Arnold (2009); Oesch et al.
+(2010); Law et al.
+(2012); Jiang et al.
+(2013); Kawamata et al.
+(2015) and Curtis-Lake et al.
+(2016).
+These were calculated using the fractions of galaxies undergoing mergers and/or exhibiting disturbed morphologies or multiple cores.
+In Ravindranath et al.
+(2006), Oesch et al.
+(2010), and Kawamata et al.
+(2015), morphologies were determined via visual inspection.
+In the other works, morphological parameters were calculated.
+Following Guo et al.
+(2015), we also include $f_{\rm{clumpy}}$ measurements from Elmegreen et al.
+(2007).
+In the latter work, galaxies were visually inspected and categorized.
+$f_{\rm{clumpy}}$ was inferred using the abundances of clump cluster, spiral, and elliptical galaxies.
+Other categories (chain galaxies, double nuclei, and tadpoles) were excluded as they generally have small axis ratios and would not have been included according to our sample selection (Section III).
+$f_{\rm{clumpy}}$ from Bowler et al. (2017) was obtained from their merger fraction, which was determined via visual inspection.
+For all other studies listed in Table 2, $f_{\rm{clumpy}}$ measurements were given explicitly.
+In Wuyts et al.
+(2012), $f_{\rm{clumpy}}$ was measured by searching for galactic regions with blue colors and high surface brightnesses, based on visual inspection of co-added surface brightness profiles.
+Clumps in Tadaki et al.
+(2014) were identified with the clumpfind code (Williams et al., 1994), which uses multiple isophotes to find clumps.
+In Guo et al.
+(2015) and Shibuya et al.
+(2016), clumps are found using algorithms that are nearly identical to the one described in Section IV.1.
+Ribeiro et al.
+(2017) detected clumps by identifying disconnected regions at various isophotal thresholds.
+Huertas-Company et al.
+(2020) used neural networks to detect clumps with a training set of simulated images of galaxies.
+Sattari et al.
+(2023) identified clumps by modeling the smooth, background light of host galaxies with Fourier transforms and subtracting this model to reveal clumps.
+Detection techniques similar to those used in Guo et al.
+(2015) and Shibuya et al.
+(2016) were then applied to the background-subtracted images.
+Last, clumps were detected in Harikane et al.
+(2025) by performing source detection on images in the F115W bandpass.
+Subcomponents detected with $5\sigma$ significance were then visually inspected.
+For the rest of the studies shown in Figure 8, we find $f_{\rm{clumpy}}$ to be $\sim 30-40\%$ at $z\sim 0.5$, peak at $\sim 60-80\%$ at $z\sim 2$, and decline to $\sim 40-60\%$ at $z\sim 3$ (although lower fractions of $\sim 20-40\%$ were obtained by Law et al.
 2012).
-The decline continues to $\sim 10-30\%$ at $z\sim 4$ (Lotz et al.
-2006; Ravindranath et al.
-2006; Conselice & Arnold 2009; Shibuya et al.
-2016; Ribeiro et al.
-2017; although Elmegreen et al.
+The decline continues to $\sim 10-30\%$ at $z\sim 4$ (although Elmegreen et al.
 2007 and Curtis-Lake et al.
 2016 measured higher fractions of $\sim 40-50\%$).
-At $z\gtrsim 4$, previous measurements of $f_{\rm{clumpy}}$ largely divide into two groups: lower fractions of $\sim 10-20\%$ that are approximately constant with redshift (Conselice & Arnold 2009; Oesch et al.
-2010; Shibuya et al.
-2016; and Kawamata et al.
-2015, although their $f_{\rm{clumpy}}$ was measured to be $\sim 50$% at $z\sim 8$) and higher fractions of $\sim 30-70\%$ over $4<z<8$ (Elmegreen et al., 2007; Jiang et al., 2013; Curtis-Lake et al., 2016; Bowler et al., 2017; Ribeiro et al., 2017; Harikane et al., 2025).
-Another measurement of $f_{\rm{clumpy}}$ as a function of redshift was recently made by Tan et al.
+At $z\gtrsim 4$, previous measurements of $f_{\rm{clumpy}}$ largely divide into two groups: lower fractions of $\sim 10-20\%$ that are approximately constant with redshift and higher fractions of $\sim 30-70\%$ over $4<z<8$.
+The results of this work are consistent with those of most previous studies at $z\gtrsim 5.75$ and largely inconsistent at lower redshifts.
+We find that, as redshift decreases, $f_{\rm{clumpy}}$ rises more steeply in this work than in the bulk of earlier studies.
+At $z\gtrsim 5.75$, $f_{\rm{clumpy}}$ as measured in this work is $\sim 20\%$, similar to what was measured earlier at these redshifts.
+At $z\sim 4.25$, we find $f_{\rm{clumpy}}$ to be $\sim 60\%$, whereas most earlier works found values between 20 and 40%.
+At $z\sim 2.75$, our measure of $f_{\rm{clumpy}}$ is $\sim 70\%$, while most other works found it to be between 40 and 60%.
+V.3 Comparing Clumpy Galaxy Fractions between the HST and JWST Data
+Figure 9: To understand why $f_{\rm{clumpy}}$ is higher with JWST data compared with previous data from HST, it is calculated with images from both telescopes as a function of redshift in each panel, minimum fractional luminosity (a proxy for image depth, columns) and whether JWST images are at their original pixel scale and resolutions (top row) or those of HST images (bottom).
+$f_{\rm{clumpy}}$ from JWST data, as shown with the large hexagons and solid lines, lies systematically above that from HST data, plotted as small circles and dashed lines.
+As the minimum fractional luminosity gradually decreases, from left to right, $f_{\rm{clumpy}}$ from JWST data progressively increases.
+Importantly, this is seen in both sets of JWST images, which are consistent with each other to within 9%, at worst.
+However, $f_{\rm{clumpy}}$ from HST images barely changes.
+Taken together, this suggests that the increased depth of JWST is the primary reason for the higher $f_{\rm{clumpy}}$ measured in this work.
+Difference in resolution between HST and JWST plays a secondary role.
+Clumps are treated as point sources for only this comparison, as they were found to have light profiles consistent with point sources in work based on HST data.
+Our results on the redshift evolution of $f_{\rm{clumpy}}$ disagree with those of most studies listed in Table 2 at $z\sim 2.75$ and $z\sim 4.25$.
+All but one of these are based on the HST data, whereas this work is entirely based on JWST data.
+To determine the cause of the disagreement, we compute $f_{\rm{clumpy}}$ using both HST and JWST images for a subset of our sample.
+We measure $f_{\rm{clumpy}}$ as a function of redshift at various minimum fractional luminosities.
+The aim is to assess to what extent the difference in resolution and sensitivity between the telescopes affect the measurement of $f_{\rm{clumpy}}$.
+In the following, we demonstrate that it is largely the difference in sensitivity that is responsible for the disagreement in the $f_{\rm{clumpy}}$ measurements, with difference in resolution having a secondary effect.
+To select galaxies for this comparison, we match the mass-selected, star-forming sample (end of Section III) with photometric catalogs from GOODS-N (Barro et al., 2019) and GOODS-S (Guo et al., 2013).
+We use a match radius of 0.″5 and match HST/Wide Field Camera 3 (WFC3) F160W magnitudes from these catalogs and those from the JADES to within 0.5 mag.
+There are 2,007 matches in GOODS-N and 2,189 in GOODS-S.
+As studies based on HST data limited their analyses to galaxies with stellar masses $\gtrsim 10^{9}~M_{\odot}$ (see Table 2), we only consider this mass range.
+This leaves a total of 1,245 galaxies (592 in GOODS-N and 653 in GOODS-S) for our comparison.
+Clumps are identified in HST images using the same algorithm and detection criteria described in Section IV.1, with one exception.
+Instead of using the segmentation map from JADES as the outer boundary in detecting clumps, as is done in Section IV.1, a limit is placed on the distance of a given clump to the center of its host to be less than $8\times R_{\textrm{half, F444W}}$, the half-light radius in F [^4] W, following Shibuya et al.
 (2016).
-Galaxies were selected in bins of stellar mass that evolve with redshift to track the redshift evolution of progenitors of Milky Way-like systems.
-Their evolution of $f_{\rm{clumpy}}$ with redshift is consistent with the bulk of studies discussed here: $f_{\rm{clumpy}}$ was measured to be $\sim 20\%$ over $2<z<5$ and increased with redshift to $\sim 60\%$ at $z\sim 0.5$.
-The results of this work over $2<z<6.5$ are inconsistent with those from most of the previous studies at similar redshifts.
-We find that $f_{\rm{clumpy}}$ rose quickly with redshift at early times, from $\sim 20\%$ at $z>6.5$ to $\sim 80\%$ at $z\sim 2.75$ at a roughly constant rate with redshift.
-In contrast, most studies found that $f_{\rm{clumpy}}$ changed by $\lesssim 10\%$ over $4<z<8$ and began to increase at $z\lesssim 4$.
-Our measurements are consistent with those from Wuyts et al.
-(2012), Jiang et al.
-(2013), and Bowler et al.
-(2017) at the redshifts where they overlap.
-Our results are also consistent with those from Elmegreen et al.
-(2007) and Ribeiro et al.
-(2017) at $z\sim 4.25$ and those from Shibuya et al.
-(2016) at $z>6.5$.
-5.3 Comparing Clumpy Galaxy Fractions between the HST and JWST Data
-Our results on the redshift evolution of $f_{\rm{clumpy}}$ disagree with those of most studies over $2<z<6.5$.
-All but two of these are based on HST data, whereas this work is based on JWST data.
-To determine the cause of the disagreement, we compute $f_{\rm{clumpy}}$ using both HST and JWST images for a subset of our sample and vary the minimum fractional luminosity of the clumps.
-As we demonstrate in this Section, it is largely the difference in sensitivity between the two datasets that is responsible for the disagreement in the $f_{\rm{clumpy}}$ measurements shown here.
-Other factors may have secondary effects on the measurement of $f_{\rm{clumpy}}$, such as the differences between photometric redshifts, stellar masses, and SFRs from earlier studies and those used in this work.
-Galaxies were selected for this comparison as follows.
-We match the mass-selected, star-forming sample (Section 3.2) with photometric catalogs from the CANDELS survey, which were created by Barro et al.
-(2019) for GOODS-N and Guo et al.
-(2013) for GOODS-S.
-We use a match radius of 0.″5 and match HST/Wide Field Camera 3 (WFC3) F160W magnitudes to within 0.5 mag.
-There are 2,007 matches in GOODS-N and 2,189 in GOODS-S.
-As studies based on HST data limited their analyses to galaxies with stellar masses $\gtrsim 10^{9}~M_{\odot}$ (see Table 2), we only consider those with masses $\geq 10^{9}~M_{\odot}$.
-This mass cut leaves a total of 1,245 galaxies (592 in GOODS-N and 653 in GOODS-S) that will be used for our comparison.
-Clumps are identified in HST images using the algorithm by Sattari et al.
-(2023) and the detection criteria from Shibuya et al.
-(2016).
-These are the same algorithm and criteria as those used to search for clumps in the JWST images (Section 4.1).
+For the purposes of the comparison here, this outer boundary criterion is also applied to JWST images.
 HST images have a pixel scale of 0.″06 pixel-1.
-We note that the criterion on the minimum number of contiguous pixels that must be spanned by a candidate clump depends on the HST instrument used: it is 5 for ACS images and 7 for WFC3 images.
-We search for clumps in HST data over $2<z<6.5$ to ensure that the coverage in rest-frame wavelength is as similar as possible between HST and JWST (see Section 3).
-Clumps are identified in HST bandpasses that span the rest-frame NUV at various redshifts.
-They are identified in the HST instruments/bandpasses as follows: ACS/F850LP over $2<z<3.5$; WFC3/F125W over $3.5<z<5$; and WFC3/F160W over $5<z<6.5$.
+We note that the criterion on the minimum number of contiguous pixels that must be spanned by a candidate clump depends on the HST instrument used: it is 5 for ACS and 7 for WFC3 images (Shibuya et al., 2016).
+We search for clumps in HST data over $2<z<6.5$ to ensure that the coverage in rest-frame wavelength is as similar as possible between HST and JWST (see Section III).
+Clumps are identified in HST bandpasses that span the rest-frame NUV at various redshifts with ACS/F850LP ($2<z<3.5$); WFC3/F125W ($3.5<z<5$); and WFC3/F160W ($5<z<6.5$).
 Cutting on redshift removes 18 galaxies, with 1,227 remaining.
 We make sure that $f_{\rm{clumpy}}$ is measured using the same galaxies for both HST and JWST images.
-Therefore, any difference in $f_{\rm{clumpy}}$ is due to discrepancies in depth and resolution between HST and JWST images and not the selection of the sample of galaxies used to compare these data.
-Figure 9: The fraction of the NUV luminosity per galaxy that originates in clumps.
-It is calculated only for clumpy galaxies and plotted as a function of stellar mass of the host galaxy in each panel.
-Clumps must be at least as bright as the lowest minimum fractional luminosity listed in Table 1 at a given redshift.
-Redshift increases from left to right and is indicated at the top of each panel.
-Measurements for individual galaxies are shown as small gray dots.
-At each redshift, medians are calculated in three bins in mass and are indicated with large black hexagons and horizontal error bars.
-The 16th and 84th percentiles of the fraction of NUV light in each bin are shown as the vertical error bars.
-Medians and percentiles are only calculated in mass bins that have at least 10 galaxies.
-The median fraction of NUV light from clumps is $\sim 15-20\%$ and changes little with redshift or the stellar mass of the host galaxy.
-Our measurements are consistent within uncertainties with those from Guo et al.
-(2015) over $2<z<3$, which are shown as black diamonds in the leftmost panel (see text for details).
-In Figure 8, $f_{\rm{clumpy}}$ is plotted as a function of redshift for HST (indicated with dashed lines) and JWST data (indicated with solid lines).
-The limit on the luminosities of the clumps with respect to their hosts is varied in each panel.
-This is the minimum fraction of the host galaxy NUV luminosity a clump must have.
-It is highest in the left panel and selected to be 8% that of the host luminosity.
-This value is adopted from pre-JWST studies (Guo et al., 2015; Shibuya et al., 2016).
+Therefore, any difference in $f_{\rm{clumpy}}$ is due to discrepancies in depth and resolution between HST and JWST images and not the selection of the sample.
+When measuring the brightness of clumps in the HST and JWST images, they are assumed to be point sources, instead of extended sources, as was found in Section IV.2.
+This is done because clumps were found to largely resemble point sources in works based on HST data (Guo et al., 2015; Shibuya et al., 2016; Sattari et al., 2023).
+To treat clumps found in the JWST images on equal footing, they are also considered point sources only for the purpose of the comparison presented here.
+Details are given in Appendix B on the completeness of detecting clumps when treated as point sources and the aperture corrections needed.
+JWST images are matched to the resolution of HST images using empirical PSFs in the HST bandpasses adopted from the 3D-HST survey (Skelton et al., 2014).
+They have a common pixel scale of 0.″06.
+They are upscaled by a factor of two to match the pixel scale of the JWST images.
+PSF-matching kernels are then created with pypher (Boucaud et al., 2016).
+After matching the resolution of HST and JWST, we performed photometry on the clumps in the same way as described in Section IV.2.
+Aperture corrections for the HST images are 1.234, 1.337, 1.447 in GOODS-N and 1.225, 1.334, and 1.436 in GOODS-S for the F850LP, F125W, and F160W bandpasses, respectively.
+When matching JWST PSFs to those from HST, their aperture corrections agree to within 1%.
+When treating clumps as point sources in the JWST images, aperture corrections are 1.348, 1.307, and 1.299 in GOODS-N and 1.365, 1.328, and 1.330 in GOODS-S for the F090W, F115W, and F150W bandpasses, respectively.
+Comparison of evolutionary trends for $f_{\rm{clumpy}}$ between HST and JWST is presented in Figure 9.
+The minimum fractional luminosity used to identify clumps is varied in each panel.
+It is highest in the left panel, corresponding to 8% of the host luminosity (Guo et al., 2015; Shibuya et al., 2016).
 It is lowered to 5% in the middle panel and 2% in the right panel.
-The latter is the lowest clump minimum fractional luminosity measured in this work (see Section 4.2).
-[…]
-We attribute the results discussed above to the difference in sensitivity between HST and JWST.
-The latter can detect point sources that are $\sim 2-5$ times fainter than the former, depending on the wavelength [^1] https://jwst-docs.stsci.edu/jwst-opportunities-and-policies/jwst-general-science-policies/guidelines-for-proposals-where-jwst-and-hst-overlap-in-capabilities#gsc.tab=0.
-It is therefore not surprising that $f_{\rm{clumpy}}$ is systematically higher when measured with JWST images than with HST images, when the faintest clumps are considered.
-Furthermore, the clump fractions from the two datasets are consistent at the highest clump minimum fractional luminosity of 8%, despite the significant differences in angular resolution.
-The greatest difference in resolution is that between the NIRCam/F150W and WFC3/F160W bandpasses, the images in which are used to measure $f_{\rm{clumpy}}$ at $z\sim 5.75$.
-The full-width at half maximum of the PSF of the former is $\approx$0.″065 (Zhuang & Shen, 2024) and that of the latter is 0.″18 (Koekemoer et al., 2011), a factor of 3 difference.
-6 Contribution of Clumps to the Rest-Frame NUV Light of Galaxies
-We now measure the fraction of the NUV luminosity originating in clumps, $C_{\rm{UV}}$, for our sample of galaxies.
-For a given galaxy in a given redshift bin, it is calculated by summing the flux of all of its clumps in the NIRCam filter that corresponds to the rest-frame NUV (Section 3) and dividing by the flux of the host galaxy in the same filter.
-Only clumpy galaxies are used to compute this.
-All clumps are selected to be at least as bright as the lowest minimum fractional luminosity listed in Table 1 at a given redshift.
-$C_{\rm{UV}}$ is shown as a function of stellar mass and redshift in Figure 9.
-Individual measurements are shown as small gray points.
-At a given redshift, medians are calculated in three bins in mass and are indicated with large black hexagons and horizontal error bars.
-In each bin, the 16th and 84th percentiles are also computed and shown as vertical error bars.
-Medians and percentiles are only calculated for mass bins with at least 10 galaxies.
-At all masses and redshifts considered, the median $C_{\rm{UV}}$ is $15-20\%$.
-The 16th percentiles are as low as 4% and 84th percentiles are as high as 35%.
-They also vary little with mass.
-We note that several galaxies have very high fractions ($80\%$) of their NUV light emerging from clumps.
-These are found predominantly at the lowest masses considered.
-Our measurements at $z\sim 2.75$ are consistent with results from Guo et al.
-(2015) at overlapping redshifts and stellar masses.
-Their calculations of $C_{\rm{UV}}$ over $2<z<3$ are shown as small, black diamonds in the leftmost column of Figure 9.
-In that work, $C_{\rm{UV}}$ was computed as a function of stellar mass for galaxies in two survey fields, GOODS-S and UDS.
-The running average over both fields is shown in the Figure.
-Their $C_{\rm{UV}}$ values were corrected for incompleteness and averaged over all galaxies, including non-clumpy systems.
-In contrast, our $C_{\rm{UV}}$ values are not corrected for completeness and are computed only for clumpy galaxies.
-Their $C_{\rm{UV}}$ values are systematically higher than ours at fixed mass by $3-12\%$.
-Despite this, our measurements and theirs are consistent to within the uncertainties.
-At higher redshifts, measurements of $C_{\rm{UV}}$ from other studies are scarce.
-Bowler et al.
-(2017) mentioned that clumps made up the majority of the UV light in their sample of 25 bright Lyman-break galaxies at $z\simeq 7$.
-Chen et al.
-(2023) examined the JWST/NIRCam images of a sample of 12 UV-bright galaxies over $z\simeq 6-8$.
-They found that clumps, on average, account for 76% of the light in F200W for the galaxies in their sample.
-These measurements exceed our median $C_{\rm{UV}}$ values of $15-20\%$ at similar redshifts.
-However, these previous works considered clumps at all radii in their samples of galaxies, whereas only off-center clumps are investigated in this work.
-The higher $C_{\rm{UV}}$ reported by these studies is likely the result of light from the centers of galaxies being included in their measurements.
-7 Clumps and the Evolution of Galaxies
-In this Section, we examine the correlations between two physical properties of galaxies, stellar mass and sSFR, and the fraction of clumpy galaxies.
-The former is shown in Figure 10 and the latter in Figure 11.
-Each is considered in its own subsection below.
-All uncertainties on $f_{\rm{clumpy}}$ are calculated assuming Bayesian confidence intervals on the binomial distribution.
-7.1 Stellar Mass vs. Fraction of Clumpy Galaxies
+The latter is the lowest clump minimum fractional luminosity at which a complete sample of clumps can be detected in the JWST images, when treating them as point sources.
+$f_{\rm{clumpy}}$ measurements from JWST that are calculated with images at the original resolution are shown in the top row, while those at the resolution of HST are shown on the bottom.
+Two conclusions may be drawn from Figure 9.
+First, at fixed minimum fractional luminosity, $f_{\rm{clumpy}}$ from JWST data changes little with resolution.
+Second, as the minimum fractional luminosity is progressively lowered, $f_{\rm{clumpy}}$ from HST barely changes, while that from JWST gradually changes.
+To assess how much overlap there is in the clumps detected between the HST and JWST datasets, we count the number of clumps in the HST images with minimum fractional luminosity $\geq 8\%$.
+Next, for each clump detected with HST, we compare its $x$ and $y$ coordinates with those of clumps detected in JWST images with the same pixel scale as HST for the same host galaxy.
+We then count the number of clumps in the JWST images that lie within a given separation in pixels from its counterpart in the HST images.
+At a separation of 1.5 pixels, which corresponds to the PSF FWHM of F850LP (0.″09, Guo et al., 2013), 49% of clumps are matched between the two datasets.
+At a larger separation of 3 pixels, which is equal to the FWHM of the F160W PSF (0.″18), 72% of the clumps are matched.
+There are a few sources of uncertainty in these measurements.
+The center of the galaxy is determined separately for the HST and JWST data and may not be identical between the two.
+Detection images from JWST are significantly deeper than the HST, which could affect the determination of the centers of galaxies.
+Similarly, since a clump must possess sufficient contrast with the local background to be detected, differences in depth will affect how well clumps are detected and deblended.
+This is evident in our measurements.
+There are about 2.5 times more clumps in the convolved JWST images than in the HST (2,512 vs.
+989) for the same galaxies and at a minimum fractional luminosity of $\geq 8\%$.
 Figure 10: $f_{\rm{clumpy}}$ as a function of host galaxy stellar mass using measurements from this work (left panel).
@@
 Another (Huertas-Company et al., 2020) found a positive correlation, but lower $f_{\rm{clumpy}}$ values than ours by about a factor of 3.
-The fraction of clumpy galaxies is measured as a function of stellar mass of their host galaxy and in various bins in redshift and plotted in the left panel of Figure 10.
+We attribute the results discussed above mostly to the difference in sensitivity between HST and JWST.
+The latter can detect point sources that are $\sim 2-5$ times fainter than the former, depending on the wavelength [^1] https://jwst-docs.stsci.edu/jwst-opportunities-and-policies/jwst-general-science-policies/guidelines-for-proposals-where-jwst-and-hst-overlap-in-capabilities#gsc.tab=0.
+It is therefore not surprising that $f_{\rm{clumpy}}$ is systematically higher when measured with JWST, when the faintest clumps are considered.
+Furthermore, our finding that resolution has a secondary ($\leq$9 %) effect on $f_{\rm{clumpy}}$ is similar to what Guo et al.
+(2012) found with HST images for a sample of 40 bright clumps in star-forming galaxies over $1.5<z<2.5$ and with stellar masses $M_{\star}\gtrsim 10^{10}M_{\odot}$.
+VI Clumpy Fraction as a Function of Stellar Mass
+The fraction of clumpy galaxies as a function of stellar mass of their host galaxy is plotted in redshift bins in Figure 10, left panel.
 The fraction is calculated in three bins, spanning stellar masses $8<\log\left(M_{\star}/M_{\odot}\right)<9$, $9<\log\left(M_{\star}/M_{\odot}\right)<10$, and $10<\log\left(M_{\star}/M_{\odot}\right)<11$.
-These are shown as hexagons connected with solid lines in a different color for each redshift bin.
 Fractions are only computed for bins that have at least 10 galaxies.
-Clumps are selected according to the minimum fractional luminosities as a function of stellar mass and redshift listed in Table 1.
-Figure 11: The correlation between the fraction of clumpy galaxies ($f_{\rm{clumpy}}$) and sSFR is plotted.
-The former is calculated in three bins in stellar mass, indicated with various line styles, and three bins in sSFR.
-Redshift increases from top to bottom and is indicated in the top left corner of each panel.
-At the lowest stellar masses, there is a mild correlation or anti-correlation between sSFR and $f_{\rm{clumpy}}$, depending on the redshift.
-At higher masses, there is no clear correlation, once uncertainties are taken into account.
-Measurements from Shibuya et al.
-(2016) over $2<z<3$ are shown in the top row as gray symbols.
 At all redshifts, $f_{\rm{clumpy}}$ increases with the stellar mass.
-At fixed mass, as redshift increases, $f_{\rm{clumpy}}$ decreases.
-At the lowest redshifts, $z\sim 2.75$, $f_{\rm{clumpy}}$ increases from $\sim 30\%$ at $\log\left(M_{\star}/M_{\odot}\right)\sim 8.5$ to $\sim 70\%$ at $\log\left(M_{\star}/M_{\odot}\right)\sim 9.5$ and $\sim 90\%$ at $\log\left(M_{\star}/M_{\odot}\right)\sim 10.5$.
-The rate at which $f_{\rm{clumpy}}$ increases with stellar mass is similar at higher redshifts.
-It grows from $\sim 20\%$ at the lowest masses to $\sim 70\%$ at the highest masses at $z\sim 4.25$.
-At $z\sim 5.75$, it rises from $\sim 10\%$ at $\log\left(M_{\star}/M_{\odot}\right)\sim 8.5$ to $\sim 40\%$ at $\log\left(M_{\star}/M_{\odot}\right)\sim 9.5$.
-Over the same range in mass at $z>6.75$, it increases from $\sim 3\%$ to $\sim 15\%$.
-The results above were found by selecting clumps using the minimum fractional luminosities listed in Table 1, which vary with stellar mass and redshift.
+At a fixed mass, as redshift increases, $f_{\rm{clumpy}}$ decreases.
+At the lowest redshifts, $z\sim 2.75$, $f_{\rm{clumpy}}$ increases from $\sim 20\%$} at $\log\left(M_{\star}/M_{\odot}\right)\sim 8.5$ to $\sim 70\%$ at $\log\left(M_{\star}/M_{\odot}\right)\sim 9.5$ and $\sim 85\%$ at $\log\left(M_{\star}/M_{\odot}\right)\sim 10.5$.
+At higher redshifts, $f_{\rm{clumpy}}$ increases with stellar mass at a rate similar to that at $z\sim 2.75$.
+At $z\sim 4.75$, it grows from $\sim 5\%$ at the lowest masses to $\sim 70\%$ at the highest masses.
+At $z\sim 5.75$, it rises from $\sim 1\%$ at $\log\left(M_{\star}/M_{\odot}\right)\sim 8.5$ to $\sim 20\%$ at $\log\left(M_{\star}/M_{\odot}\right)\sim 9.5$.
+At $z\sim 6.75$, a complete sample of clumps is only detected at $\log\left(M_{\star}/M_{\odot}\right)\sim 9.5$, for which $f_{\rm{clumpy}}$ is $\sim 17\%$.
+The above results were found by selecting clumps using the minimum fractional luminosities listed in Table 1, which vary with stellar mass and redshift.
 However, the correlation between stellar mass and $f_{\rm{clumpy}}$ still holds when selecting clumps with the same minimum fractional luminosity at all masses.
 The fraction of clumpy galaxies is calculated again for galaxies with stellar mass $\log\left(M_{\star}/M_{\odot}\right)\sim 8.5$, with the minimum fractional luminosities this time being the same as those applied to higher-mass galaxies.
-These alternative measurements are plotted as faint hexagons connected with dashed lines in the left panel in Figure 10.
-The effect of lowering the minimum fractional luminosity for the lowest-mass galaxies is an increase in $f_{\rm{clumpy}}$ by $\sim 10\%$ at fixed redshift.
-At $z\lesssim 5.75$, this makes the slope between stellar mass and $f_{\rm{clumpy}}$ slightly shallower than in the case where the minimum fractional luminosities vary with mass.
-At $z>6.5$, the slope is nearly flat.
-However, the uncertainties on $f_{\rm{clumpy}}$ are the largest at these redshifts.
-More data are needed to determine the correlation between stellar mass and $f_{\rm{clumpy}}$ at the highest redshifts considered in this work.
+The result is shown in the left panel in Figure 10.
+The effect of lowering the minimum fractional luminosity for the lowest-mass galaxies is an increase in $f_{\rm{clumpy}}$ by $\sim 10-20\%$ at fixed redshift.
 We note that the correlation observed in the present work may result from selection effects.
@@
 The physical properties of clumps that are identified in this work, such as stellar masses and SFRs, are needed to determine the impact of these selection effects.
-In the right panel in Figure 10, we compare the correlation between stellar mass and $f_{\rm{clumpy}}$ at $z\sim 2.75$ with that measured by other studies over $2<z<3$.
+In Figure 10, right panel, we compare the correlation between stellar mass and $f_{\rm{clumpy}}$ at $z\sim 2.75$ with that measured by other studies over $2<z<3$.
 These measurements are shown as various symbols.
@@
 In contrast, Huertas-Company et al.
-(2020) found that, as stellar mass increases, $f_{\rm{clumpy}}$ increases.
-From $\log\left(M_{\star}/M_{\odot}\right)\sim 9$ to $\log\left(M_{\star}/M_{\odot}\right)\sim 11$, it grows from $\sim 20\%$ to $\sim 30\%$.
-In that study, neural networks were trained on mock HST images from the VELA simulations (Ceverino et al., 2014) and used to identify clumps in real HST images spanning the rest-frame UV and optical.
-Their sample consisted of 1,575 galaxies over $0.5<z<3$ and $9<\log\left(M_{\star}/M_{\odot}\right)<12$.
-Results from the present work at $z\sim 2.75$ disagree with those from other studies at similar redshifts.
-We find a positive correlation between stellar mass and $f_{\rm{clumpy}}$, such that, as stellar mass increases from $\log\left(M_{\star}/M_{\odot}\right)\sim 8.5$ to $\log\left(M_{\star}/M_{\odot}\right)\sim 10.5$, $f_{\rm{clumpy}}$ increases about threefold from $\sim 30\%$ to $\sim 90\%$.
-Guo et al.
-(2015) and Sattari et al.
-(2023) found an anti-correlation and, at fixed mass, overall lower fractions than those measured in this study.
-Huertas-Company et al.
-(2020) also found a positive correlation, but lower $f_{\rm{clumpy}}$ values than ours by a factor of 2-3 at fixed mass.
-At $z>2.75$, results from this work qualitatively agree with those from other studies.
-We find that more massive galaxies are more likely to host clumps at all redshifts.
-Earlier studies that examined galaxies over $6\lesssim z\lesssim 8$ found that brighter galaxies were more likely to have clumps or multiple cores and fainter galaxies were more likely to have one component (Kawamata et al., 2015; Bowler et al., 2017).
-7.2 Specific SFR vs. Fraction of Clumpy Galaxies
-We now examine the correlation between sSFR and $f_{\rm{clumpy}}$.
-This is shown in Figure 11.
-It is examined in the same bins in stellar mass considered in the previous Section, which are indicated with various line styles in the Figure.
-For each bin in mass, $f_{\rm{clumpy}}$ is measured in three 0.5-dex-wide bins in sSFR.
-They are centered at $\log\left(\textrm{sSFR/yr}^{-1}\right)\sim-9.25,\sim-8.75,$ and $\sim-8.25$.
-Redshift increases from the top row to the bottom.
-At the lowest stellar masses of $\left(\log\left(M_{\star}/M_{\odot}\right)\sim 8.5\right)$, there is a mild correlation or anti-correlation between sSFR and $f_{\rm{clumpy}}$, depending on the redshift.
-At higher stellar masses, there is little to no correlation.
-At $z\sim 2.75$ and $\log\left(M_{\star}/M_{\odot}\right)\sim 8.5$, as sSFR increases, $f_{\rm{clumpy}}$ increases from 20% to 35%.
-At higher stellar masses, as sSFR increases, $f_{\rm{clumpy}}$ varies by $\leq 10\%$.
-Similar behavior is observed at $z\sim 4.25$, once uncertainties are taken into account.
-For the lowest-mass galaxies, as sSFR increases, $f_{\rm{clumpy}}$ increases from 7% to 23%.
-For higher-mass galaxies, as sSFR increases, $f_{\rm{clumpy}}$ varies by up to $\sim 30\%$.
-However, the greatest uncertainties are also this large, particularly at the lowest sSFRs and highest masses.
-Therefore, changes in $f_{\rm{clumpy}}$ with sSFR are consistent with little to no correlation at these masses.
-At $z\sim 5.75$, for the lowest-mass galaxies, as sSFR increases, $f_{\rm{clumpy}}$ decreases from 22% to 10%.
-However, uncertainties on $f_{\rm{clumpy}}$ at the lowest sSFRs are relatively large, $\sim 10-20\%$.
-Therefore, this is not a significant anti-correlation.
-At higher masses, as sSFR increases, $f_{\rm{clumpy}}$ rises from zero to 43% and declines to 38%.
-This is due to the lowest bin in sSFR having only one galaxy and thus is not significant.
-Finally, at $z>6.5$, at all masses, as sSFR increases, $f_{\rm{clumpy}}$ increases from zero and plateaus.
-However, this is not a significant correlation because the lowest sSFR bins have at most three galaxies.
-Consequently, the uncertainties at the lowest sSFRs are very large (up to $\sim 50\%$) and are consistent with little to no trend with sSFR.
-The lack of a significant correlation between sSFR and $f_{\rm{clumpy}}$ at $z\sim 2.75$ for galaxies of stellar mass $M_{\star}\geq 10^{9}M_{\odot}$ is generally consistent with what previous studies found at similar redshifts, masses, and sSFRs.
-As sSFR increases over $\sim-9.25<\log\left(\textrm{sSFR/yr}^{-1}\right)<\sim-8.75$, Shibuya et al.
-(2016) measured $f_{\rm{clumpy}}$ to increase by $5-10\%$ for galaxies with masses over $9<\log\left(M_{\star}/M_{\odot}\right)<11$ and redshifts over $2<z<3$.
-For higher-mass galaxies, $10<\log\left(M_{\star}/M_{\odot}\right)<11$, $f_{\rm{clumpy}}$ then drops by $\sim 20\%$ at higher sSFRs.
-However, the uncertainty on this is of similar size, therefore, this decline is not significant.
-Similar increases of $5-10\%$ are seen in our sample at similar masses and sSFRs, despite our  $f_{\rm{clumpy}}$ values being about $20-40\%$ higher.
-Our uncertainties on $f_{\rm{clumpy}}$ at these masses and sSFRs range from $3-7\%$.
-Those from Shibuya et al.
-(2016) are similar, about $5-7\%$.
-Therefore, this small change in $f_{\rm{clumpy}}$ with sSFR is not significant.
-Similarly, Huertas-Company et al.
-(2020) found no correlation between $f_{\rm{clumpy}}$ and the location of galaxies relative to the SFMS over $2<z<3$.
-8 Comparison with Hydrodynamic Simulations
+(2020) found that it has a positive slope, with $f_{\rm{clumpy}}$ increasing from $\sim 20\%$ to $\sim 30\%$ with as stellar mass increases from $\log\left(M_{\star}/M_{\odot}\right)\sim 9$ to 11.
+Results from the present work at $z\sim 2.75$ disagree with those from other studies at similar redshifts, finding a positive correlation between stellar mass and $f_{\rm{clumpy}}$.
+VII Comparison with Hydrodynamic Simulations
 Here we discuss the implication of our results to galaxy formation scenarios.
-There are two key results of this work.
-First, the fraction of clumpy galaxies, $f_{\rm{clumpy}}$, increases with decreasing redshift over $2<z<12$.
-For galaxies with stellar mass $M{\star}\geq 10^{9}M_{\odot}$, $f_{\rm{clumpy}}$ is $\sim 20\%$ at $z>6.5$ and $\sim 80\%$ at $z\sim 2.75$.
-$f_{\rm{clumpy}}$ is found to be higher, at fixed redshift, and increase more rapidly with decreasing redshift than what was found in many previous studies (Section 5 and Figure 7).
-We interpret this as a consequence of the higher sensitivity of the JWST compared to the HST, which nearly all earlier studies relied on (Section 5.3).
-Second, at a fixed redshift, as stellar mass increases, the fraction of clumpy galaxies increases (Section 7.1 and Figure 10).
-In this Section, we interpret our observations within the context of clump formation over $2<z<12$.
-Our results are compared with those from simulations of clumpy galaxies.
-In doing so, we identify what might be the dominant processes for clumpy galaxy formation as a function of mass and redshift.
-In Section 9, our results are studied in the context of other observations.
-Figure 12: $f_{\rm{clumpy}}$ as a function of redshift in three stellar mass bins.
-This is indicated with thick black lines.
-Predictions for $f_{\rm{clumpy}}$ based on simulations are also shown.
-Their corresponding uncertainties are shown as shaded regions.
+There are two key results.
+First, the fraction of clumpy galaxies, $f_{\rm{clumpy}}$, increases with decreasing redshift over 2<z<9{2<z<9}.
+For galaxies with stellar mass $M{\star}\geq 10^{9}M_{\odot}$, $f_{\rm{clumpy}}$ is $\sim 10\%$} at $z\sim 7.75$} and $\sim 70\%$} at $z\sim 2.75$.
+We find relatively higher $f_{\rm{clumpy}}$ in JWST-selected samples compared to those selected by HST.
+This is due to the higher sensitivity of the JWST observations (Section V and Figure 8).
+Second, at a fixed redshift, as stellar mass increases, the fraction of clumpy galaxies increases too (Section VI and Figure 10).
+In this Section, we compare our results with simulations of clumpy galaxies.
+This is done to identify potential physical mechanisms for their formation.
+We compare with zoom-in hydrodynamic cosmological simulations from Mandelker et al.
+(2017) and Nakazato et al.
+(2024).
+The former studied 34 galaxies over $1\lesssim z\lesssim 6$ from the VELA simulations (Ceverino et al., 2014; Zolotov et al., 2015), which had stellar masses of $\log\left(M_{\star}/M_{\odot}\right)\gtrsim 9$ at $z=2$, with clumps largely forming through VDI.
+The latter analyzed 62 galaxies over $5.5<z<9.5$ from the FirstLight simulations (Ceverino et al., 2017), which had stellar masses in the range $(0.5-6)\times 10^{10}M_{\odot}$ at $z=5$.
+Most clumps in these simulations formed from major mergers.
+The $f_{\rm{clumpy}}$  estimates and predictions are compared in Figure 11.
+Figure 11: Observed $f_{\rm{clumpy}}$ values as a function of redshift in three stellar mass bins compared with predictions from simulations.
+Uncertainties are shown as shaded regions.
 The fraction of clumpy galaxies that formed from violent disk instabilities (VDI; Mandelker et al., 2017) at $z<6$ is shown as gray, red, and green shaded regions, in the order of increasing stellar mass.
 Predictions for clumpy galaxies that formed from major mergers over $6\lesssim z\lesssim 9$ (Nakazato et al., 2024) are shown as a purple line and shaded region.
-Observed pair fractions for galaxies of stellar mass over $8<\log\left(M_{\star}/M_{\odot}\right)<10$ and redshift over $3<z<9$ (Puskás et al., 2025) are shown as thin gray lines in various styles.
-Trends for the least massive, intermediate, and most massive galaxies are indicated with dotted, dash-dot, and solid lines, respectively.
-The results from this work are consistent with $f_{\rm{clumpy}}$ due to VDI at $z<6$ and $\log\left(M_{\star}/M_{\odot}\right)\geq 9$ and that due to major mergers at $z>6$ and $\log\left(M_{\star}/M_{\odot}\right)\leq 9$.
+Observed pair fractions as a function of stellar mass and redshift are shown as thin gray lines (Puskás et al., 2025).
+The results from this work are consistent with $f_{\rm{clumpy}}$ from simulations based on VDI and those based on major mergers at redshifts and masses where they overlap.
 The observed pair fractions at $\log\left(M_{\star}/M_{\odot}\right)\leq 9$ are also consistent with the observed $f_{\rm{clumpy}}$ measured here at similar masses and redshifts.
-Taking these comparisons together, our results are consistent with a physical scenario in which VDI is the dominant mechanism behind clumpy galaxies at low redshifts and intermediate and high masses and mergers dominate at high redshifts and the lowest masses.
-The $f_{\rm{clumpy}}$ measurements are compared with those from zoom-in hydrodynamic cosmological simulations of clumpy galaxies examined by Mandelker et al.
-(2017) and Nakazato et al.
-(2024) in Figure 12.
-The former studied 34 galaxies over $1\lesssim z\lesssim 6$ from the VELA simulations (Ceverino et al., 2014; Zolotov et al., 2015), which had stellar masses of $\log\left(M_{\star}/M_{\odot}\right)\gtrsim 9$ at $z=2$.
-In that work, clumps formed largely through VDI.
-The latter analyzed 62 galaxies over $5.5<z<9.5$ from the FirstLight simulations (Ceverino et al., 2017), which had stellar masses in the range of $(0.5-6)\times 10^{10}M_{\odot}$ at $z=5$.
-Most clumps in these simulations formed from major mergers.
-The $f_{\rm{clumpy}}$ measurements from the present work are plotted as thick black lines in Figure 12.
-Predictions from simulations are shown as lines with shaded regions.
-$f_{\rm{clumpy}}$ is shown for various mass ranges, where applicable, and indicated with various line styles.
-The $f_{\rm{clumpy}}$ measurements in the present work are consistent with those from Mandelker et al.
+The $f_{\rm{clumpy}}$ measurements are consistent with predictions from Mandelker et al.
 (2017) at $z\lesssim 5.75$ and $\log\left(M_{\star}/M_{\odot}\right)\geq 9$.
-The $f_{\rm{clumpy}}$ measurements from Mandelker et al.
-(2017) increase rapidly with redshift from $\sim 20\%$ at $z\sim 1$ to $\sim 50\%$ $z\sim 2$, above which they asymptote.
-They have uncertainties of $\sim 20\%$, which result from Poisson noise due to a limited number of galaxies examined in that work.
-Our $f_{\rm{clumpy}}$ measurements peak at $\sim 70-90\%$ at $z\sim 2.75$ and decline with increasing redshift, while those from Mandelker et al.
-(2017) do not change with redshift within the uncertainties.
-The discrepancy in $f_{\rm{clumpy}}$ as a function of redshift between the theoretical predictions and our observations may arise from differences in the way clumps are identified.
+The predictions increase rapidly with redshift from $\sim 20\%$ at $z\sim 1$ to $\sim 50\%$ at $z\sim 2$, above which they asymptote.
+They have uncertainties of $\sim 20\%$, which result from Poisson noise due to a limited number of galaxies examined.
+While our $f_{\rm{clumpy}}$ measurements peak at $\sim 70-80\%$} at $z\sim 2.75$ and decline with increasing redshift, the model predictions do not change with redshift, when taking into account the uncertainties.
+This discrepancy is likely due to differences in the way clumps are identified.
 Mandelker et al.
-(2017) searched for clumps in projected density maps of their galaxies and computed $f_{\rm{clumpy}}$ using those with stellar masses $10^{7}~M_{\odot}$, SFRs of at least 0.1% those of their host galaxies’ disks.
+(2017) searched for clumps in projected density maps of their galaxies and computed $f_{\rm{clumpy}}$ using those with stellar masses $10^{7}~M_{\odot}$ and SFRs of at least 0.1% those of their host galaxies’ disks.
 In this work, clumps are detected in rest-frame NUV images and selected to be compact objects that are at least as bright as a fraction of their hosts’ NUV luminosity, depending on redshift and stellar mass of the host galaxy.
 In both studies, clumps are selected to be off-center.
-However, at the redshifts and stellar masses where the two studies overlap, the measurements are largely consistent within the uncertainties.
 Our $f_{\rm{clumpy}}$ measurements are also consistent with those from Nakazato et al.
-(2024) at $z\sim 5.75$ and $z>6.5$ and $\log\left(M_{\star}/M_{\odot}\right)<9$.
-As redshift decreases, our measured $f_{\rm{clumpy}}$ increase from 3% at $z>6.5$ to 11% at $z\sim 5.75$.
-Those from Nakazato et al.
-(2024) are roughly constant with redshift at $\sim 10\%$.
-Galaxies studied by that work had stellar masses of $\lesssim 10^{9}~M_{\odot}$ at $z\geq 8.5$.
-Therefore, trends in $f_{\rm{clumpy}}$ at the lowest masses considered in this work are in overall agreement with those from Nakazato et al.
-(2024) at redshifts where there is overlap.
-However, clumps were identified differently between these two methods.
+(2024) at $z\sim 7.75$} and $\log\left(M_{\star}/M_{\odot}\right)\sim 9.5$}.
+At these redshifts and stellar masses, we find $f_{\rm{clumpy}}$ to be 1313%.
+A similar, near-constant value of $\sim 10\%$ is obtained by Nakazato et al.
+(2024) over $6<z<9$.
+Galaxies studied in that work had stellar masses of $\lesssim 10^{9}~M_{\odot}$ at $z\geq 8.5$.
+Therefore, trends in $f_{\rm{clumpy}}$ at the intermediate masses considered in this work are consistent with those from Nakazato et al.
+(2024) at redshifts where they overlap.
+However, clumps were identified differently between these two works.
 In Nakazato et al.
@@
 Detectable clumps had stellar masses of $\gtrsim 10^{7}~M_{\odot}$.
-This is comparable to the faintest clumps that can be detected in this work at $z\sim 5.75$ and over $8<\log\left(M_{\star}/M_{\odot}\right)<9$, assuming clumps have similar NUV mass-to-light ratios as their hosts at these redshifts and stellar masses.
-In Nakazato et al.
+This is about an order of magnitude lower than the faintest clumps that can be detected in this work at $z\sim 7.75$} and $\log\left(M_{\star}/M_{\odot}\right)\sim 9.5$, assuming clumps have similar NUV mass-to-light ratios as their hosts at these redshifts and stellar masses.
+Furthermore, in Nakazato et al.
 (2024), galaxies were defined to be clumpy if they had at least two clumps.
-In this work, a galaxy must have at least one off-center clump to be considered clumpy.
-It is unclear from the simulations which processes dominate the formation of clumpy galaxies at other combinations of redshift and stellar mass, namely, low redshift ($z<5.75$) and low mass ($\log\left(M_{\star}/M_{\odot}\right)<9$), and high redshift ($z>5.75$) and high mass ($\log\left(M_{\star}/M_{\odot}\right)\geq 9$).
-In the case of the former, only two galaxies in the sample of Mandelker et al.
-(2017) had such low masses at $z=2$.
+It is unclear from the simulations which processes dominate the formation of clumpy galaxies at other combinations of redshift and stellar mass, particularly at the lowest masses.
 Predictions for $f_{\rm{clumpy}}$ for galaxies at these masses and redshifts are needed.
-In the case of the latter, predictions from simulations that produce clumps through VDI are needed out to higher redshifts in order to distinguish clumps formed through VDI and those from mergers.
-Our $f_{\rm{clumpy}}$ measurements at intermediate stellar masses ($9<\log\left(M_{\star}/M_{\odot}\right)<10$) are about four times higher than those from Nakazato et al.
+Predictions from simulations that produce clumps through VDI are needed out to higher redshifts in order to distinguish clumps formed through VDI and those from mergers.
+Our $f_{\rm{clumpy}}$ measurements at intermediate stellar masses ($9<\log\left(M_{\star}/M_{\odot}\right)<10$) are about two times higher than those from Nakazato et al.
 (2024) at $z\sim 5.75$.
@@
 However, confirming this will require $f_{\rm{clumpy}}$ measurements from simulations of clumpy galaxies that form via other mechanisms at these redshifts.
-9 Discussion
-Comparison of our observations and predictions from simulations suggests that VDI and mergers are the likely dominant mechanisms for clumpy galaxy formation, depending on their redshift and stellar mass, in agreement with other studies (e.g., Guo et al., 2015; Shibuya et al., 2016).
-VDI appears to be the dominant mechanism for forming clumps at $z\lesssim 5.75$ and at intermediate-to-high stellar masses ($\log\left(M_{\star}/M_{\odot}\right)\geq 9$).
-Mergers likely dominate at $z\gtrsim 5.75$ and low stellar masses ($8<\log\left(M_{\star}/M_{\odot}\right)<9$).
-We now assess the plausibility of these mechanisms by considering observations of the morphologies, gas kinematics, and pair fractions of star-forming galaxies with similar redshifts and stellar masses as those studied in this work.
-We find that observations largely support the scenario described above.
-For clumps to form through VDI, their host galaxies must possess marginally stable disk components (e.g., Bournaud et al., 2007; Agertz et al., 2009; Dekel et al., 2009a; Ceverino et al., 2010; Mandelker et al., 2014).
+VIII Discussion
+In this section, we discuss observations of the morphologies, gas kinematics, and pair fractions of star-forming galaxies with similar redshifts and stellar masses as those studied in this work.
+This is done to assess the plausibility of the mechanisms considered in the previous section, wherein our measurements of $f_{\rm{clumpy}}$ were compared with predictions from simulations.
+The comparison suggests that VDI and mergers may be the dominant mechanisms for clumpy galaxy formation, depending on redshift and stellar mass, in agreement with other studies (e.g., Guo et al., 2015; Shibuya et al., 2016).
+The dominant mechanism for forming clumps appears to be VDI at $z\lesssim 5.75$ and at intermediate-to-high stellar masses ($\log\left(M_{\star}/M_{\odot}\right)\geq 9$) and mergers at $z\gtrsim 7.75$} and intermediate stellar masses ($\log\left(M_{\star}/M_{\odot}\right)\sim 9.5$}).
+We find that the observations listed above largely support the scenario described here.
+For clumps to form through VDI, their host galaxies must possess marginally stable disk components (e.g., Bournaud et al., 2007; Agertz et al., 2009; Dekel et al., 2009b; Ceverino et al., 2010; Mandelker et al., 2014).
 Star-forming galaxies over $2<z\lesssim 6$ and $\log\left(M_{\star}/M_{\odot}\right)>8$ have typical Sérsic indices of $n\simeq 1$ (van der Wel et al., 2014; Shibuya et al., 2015, 2016; Nedkova et al., 2021; Ormerod et al., 2024; Sun et al., 2024; Varadaraj et al., 2024; Ward et al., 2024; Miller et al., 2025; Yang et al., 2025), which suggests they have disk-like structures.
@@
 (2025) in Figure 11.
-The latter are shown as thin gray lines.
-Various bins in stellar mass are indicated with various line styles.
 In that work, the pair fraction was measured over $3<z<9$ and $8<\log\left(M_{\star}/M_{\odot}\right)<10$ in both of the JADES fields.
@@
 Their pair fractions agree with other recent measurements of the pair fraction at high redshift with JWST (Dalmasso et al., 2024; Duan et al., 2025).
-Our $f_{\rm{clumpy}}$ measurements over $8<\log\left(M_{\star}/M_{\odot}\right)<9$ are consistent with pair fractions from Puskás et al.
-(2025) over the same range in mass and the overlapping range in redshift.
-This suggests that major mergers are the dominant mechanism for clumpy galaxy formation at low stellar masses over the redshift range $3<z<9$.
-At higher stellar masses ($9<\log\left(M_{\star}/M_{\odot}\right)<10$), $f_{\rm{clumpy}}$ is consistent with the pair fraction only at $z>6.5$.
-At lower redshifts, the former is three to four times higher than the latter.
+Our $f_{\rm{clumpy}}$ measurements are consistent with pair fractions from Puskás et al.
+(2025) at $z\sim 7.75$ and over $9<\log\left(M_{\star}/M_{\odot}\right)<10$.
+This suggests that major mergers are the dominant mechanism for clumpy galaxy formation at $z\sim 7.75$ and $\log\left(M_{\star}/M_{\odot}\right)\sim 9.5$}.
+At lower stellar masses, the pair fractions are generally much higher than $f_{\rm{clumpy}}$.
+This could suggest that clumps form inefficiently through mergers at these masses.
+However, given the lack of predictions in this mass regime, we leave interpretation to future work.
+At z<5.75{z<5.75} and $\log\left(M_{\star}/M_{\odot}\right)\sim 9.5$}, $f_{\rm{clumpy}}$ is three to four times higher than the pair fractions.
 This implies that major mergers are not the dominant process through which clumps form at these masses and redshifts.
@@
 The pair fraction at $\log\left(M_{\star}/M_{\odot}\right)>10.3$ was found to be $\sim 10-15\%$ over $2\lesssim z\lesssim 4$ and increases to $\sim 30-40\%$ at $z\sim 6$.
-This is inconsistent with our $f_{\rm{clumpy}}$ over $10<\log\left(M_{\star}/M_{\odot}\right)<11$, which decreases from $90\%$ at $z\sim 2.75$ to $73\%$ at $z\sim 4.25$.
-10 Conclusions
-We use JWST/NIRCam imaging from the JADES survey (Eisenstein et al., 2023a, b; Rieke et al., 2023b; Bunker et al., 2024; D’Eugenio et al., 2024) to identify clumpy star-forming galaxies over $2<z<12$.
+This is inconsistent with our $f_{\rm{clumpy}}$ over $10<\log\left(M_{\star}/M_{\odot}\right)<11$, which decreases from $81\%$} at $z\sim 2.75$ to $76\%$} at $z\sim 4.25$.
+We note that VDI and mergers are not the only mechanisms known to produce clumps.
+Another one, particularly active at high redshift, comes from turbulence in the interstellar medium driven by accretion of gas and stellar feedback (Sun et al., 2026).
+This can lead to strong gas temperature and density fluctuations that result in spatially extended, clumpy star formation.
+IX Conclusions
+We use JWST/NIRCam imaging from the JADES survey (Eisenstein et al., 2023b, a; Rieke et al., 2023a; Bunker et al., 2024; D’Eugenio et al., 2024) to identify clumpy star-forming galaxies over 2<z<9{2<z<9}.
 The evolution of the fraction of clumpy galaxies, $f_{\rm{clumpy}}$, with redshift and correlations with other physical properties were examined.
@@
 Clumps were selected in a way similar to earlier studies based on HST (e.g., Guo et al., 2015, 2018; Shibuya et al., 2016; Sattari et al., 2023).
-The algorithm in Sattari et al.
-(2023) was used to find off-center clumps in rest-frame NUV images of galaxies.
+The algorithm described in Section IV.1 was used to find off-center clumps in rest-frame NUV images of galaxies.
 The completeness of the clump detection procedure was determined by injecting fake clumps into images of real galaxies.
@@
   1. 1.
-    We find that $f_{\rm{clumpy}}$ increases with decreasing redshift from $\sim 20\%$ at $z>6.5$ to $\sim 80\%$ at $z\sim 2.75$ (Section 5 and Figure 7).
+    We find that $f_{\rm{clumpy}}$ increases with decreasing redshift from $\sim 10\%$} at $z\sim 7.75$} to $\sim 70\%$} at $z\sim 2.75$ (Section V and Figure 8).
 This is similar to the relative increase in the cosmic volume-averaged star-formation density over the same range in redshift.
@@
   2. 2.
-    The higher $f_{\rm{clumpy}}$ found in this work is due to the higher sensitivity of JWST compared with HST (Section 5.3 and Figure 8).
+    The higher $f_{\rm{clumpy}}$ found in this work is mostly due to the higher sensitivity of JWST compared with HST (Section V.3 and Figure 9).
 We calculate $f_{\rm{clumpy}}$ for a sample of 1,227 galaxies that are detected in both HST and JWST images and compare measurements from both sets of images.
-When clumps are selected to have NUV luminosities that are $\geq 8\%$ those of their host galaxies, both measurements of $f_{\rm{clumpy}}$ are consistent with each other.
-When fainter clumps are selected, $f_{\rm{clumpy}}$ from JWST data are systematically higher, which is inconsistent with that from HST.
+We repeat this for another set of JWST images that share the resolutions of the HST images.
+$f_{\rm{clumpy}}$ from JWST shows little change with resolution, but increases significantly when fainter clumps are selected.
   3. 3.
-    Off-center clumps make up $15-20\%$ of the rest-frame NUV luminosity of clumpy galaxies (Section 6 and Figure 9).
-There is no significant dependence on redshift or stellar mass of the host galaxy.
-At $z\sim 2.75$, our results are consistent with previous studies.
+    We examine the correlation between $f_{\rm{clumpy}}$ and the stellar masses of their host galaxies.
+There is a strong correlation between stellar mass and $f_{\rm{clumpy}}$, such that more massive galaxies have higher $f_{\rm{clumpy}}$ (Section VI and Figure 10).
+This is the opposite trend of most previous studies.
   4. 4.
-    We examine correlations between $f_{\rm{clumpy}}$ and physical properties of galaxies.
-There is a strong correlation between stellar mass and $f_{\rm{clumpy}}$, such that more massive galaxies have higher $f_{\rm{clumpy}}$ (Section 7.1 and Figure 10).
-This is the opposite trend of most previous studies.
-We find a mild correlation between sSFR and $f_{\rm{clumpy}}$  at the lowest masses and little to none at higher masses (Section 7.2 and Figure 11).
-This is consistent with the results of earlier studies.
-  5. 5.
-    Our $f_{\rm{clumpy}}$ measurements are consistent with a physical scenario in which clumps form largely through violent disk instabilities at $z\lesssim 5.75$ and over $9<\log\left(M_{\star}/M_{\odot}\right)<11$ and through mergers at $z\gtrsim 5.75$ and over $8<\log\left(M_{\star}/M_{\odot}\right)<9$ (Section 8 and Figure 12).
+    Our $f_{\rm{clumpy}}$ measurements are consistent with a physical scenario in which clumps form largely through violent disk instabilities at $z\lesssim 5.75$ and over $9<\log\left(M_{\star}/M_{\odot}\right)<11$ and through mergers at $z\sim 7.75$} and over $9<\log\left(M_{\star}/M_{\odot}\right)<10$} (Section VII and Figure 11).
 This is done by comparing $f_{\rm{clumpy}}$ measured from simulations with the observations.
-This physical scenario is bolstered when considering the observed morphologies, gas kinematics, and pair fractions of star-forming galaxies at redshifts and stellar masses that overlap with the sample examined in this work (Section 9).
-We release a catalog of physical properties derived from SED fits, the empirical PSFs, and code to reproduce Figure 7 at https://github.com/alex-delavega/clumpy_galaxies_jades_2_z_12.
-We thank the JADES team for designing and preparing their observations and releasing a rich public dataset.
+This physical scenario is bolstered when considering the observed morphologies, gas kinematics, and pair fractions of star-forming galaxies at redshifts and stellar masses that overlap with the sample examined in this work (Section VIII).
+We release a catalog of physical properties derived from SED fits, the empirical PSFs, and code to reproduce Figure 8 on Zenodo under an open-source Creative Commons Attribution license: https://doi.org/10.5281/zenodo.19444334 (catalog doi:10.5281/zenodo.19444334).
+We thank the anonymous referee for constructive comments which improved the paper.A.d.l.V.
+thanks Guochao Sun for enlightening discussions on formation of clumps in turbulent, high-redshift galaxies.
+A.d.l.V.
+also thanks Mark J.
+Achenbach for hitting the “submit” button on arXiv while A.d.l.V was driving in Hilo.We thank the JADES team for designing and preparing their observations and releasing a rich public dataset.
 We also thank the STScI staff for enabling this science.
 This research has made use of the SVO Filter Profile Service “Carlos Rodrigo”, funded by MCIN/AEI/10.13039/501100011033/ through grant PID2023-146210NB-I00.
-All of the JWST data used in this paper can be found in MAST (Rieke et al., 2023a; Williams et al., 2023a), as can all of the HST data used in this paper (Faber, 2011).
+All of the JWST data used in this paper can be found in MAST (Rieke et al., 2023b; Williams et al., 2023b), as can all of the HST data used in this paper (Faber, 2011).
 Appendix A Empirical Point Spread Functions
-Empirical PSFs were measured using photutils EPSFBuilder routine, which is based on the technique developed by Anderson & King (2000).
-We adopt the parameters used to construct PSFs in photutils from Zhuang & Shen (2024), who characterized JWST/NIRCam PSFs in Cycle 0 calibration programs.
+Empirical PSFs were measured using photutils EPSFBuilder routine, which is based on the technique developed by Anderson and King (2000).
+We adopt the parameters used to construct PSFs in photutils from Zhuang and Shen (2024), who characterized JWST/NIRCam PSFs in Cycle 0 calibration programs.
 In the present work, PSFs were created using bright (F277W mag <22<22), unsaturated, and isolated point sources that were visually inspected in each NIRCam bandpass.
@@
 This is repeated in the third and bottom rows for GOODS-N.
-The PSFs are estimated using the algorithm by Anderson & King (2000) and implemented in photutils (Bradley, 2023).
-The parameters that are used to construct the PSFs with photutils are adopted from Zhuang & Shen (2024).
+The PSFs are estimated using the algorithm by Anderson and King (2000) and implemented in photutils (Bradley, 2023).
+The parameters that are used to construct the PSFs with photutils are adopted from Zhuang and Shen (2024).
 The curves of enclosed energy that are measured in this work agree to within 5% with those measured by Ji et al.
@@
 The enclosed energy curves from that work are shown as dashed lines in the second row.
-Appendix B Measurements of the Fraction of Clumpy Galaxies from Previous Studies
-In this Section, we describe how measurements of $f_{\rm{clumpy}}$ were extracted from the earlier studies listed in Table 2.
-Following Shibuya et al.
-(2016), $f_{\rm{clumpy}}$ from Lotz et al.
-(2006); Ravindranath et al.
-(2006); Conselice & Arnold (2009); Oesch et al.
-(2010); Law et al.
-(2012); Jiang et al.
-(2013); Kawamata et al.
-(2015) and Curtis-Lake et al.
-(2016) were measured using the fractions of galaxies undergoing mergers or exhibiting disturbed morphologies or multiple cores reported in these works.
-These come from the following tables, figures, or sections in these studies: Table 1 in Lotz et al.
-(2006); Table 2 in Ravindranath et al.
-(2006); Table 2 in Conselice & Arnold (2009); Section 4 in Oesch et al.
-(2010); Figure 17 in Law et al.
-(2012); Section 3.3 in Jiang et al.
-(2013); Section 3.1 in Kawamata et al.
-(2015); and Figure 17 in Curtis-Lake et al.
-(2016).
-$f_{\rm{clumpy}}$ from Bowler et al.
-(2017) was obtained from the fraction of mergers given in their Section 4.4.
-Following Guo et al.
-(2015), $f_{\rm{clumpy}}$ from Elmegreen et al.
-(2007) was inferred using the abundances of clump cluster, spiral, and elliptical galaxies, given in their Figure 6.
-Other categories (chain galaxies, double nuclei, and tadpoles) were excluded as they generally have small axis ratios and would not have been included according to our sample selection (Section 3).
-For all other studies listed in Table 2, $f_{\rm{clumpy}}$ measurements were given explicitly.
-These come from the following tables, figures, or sections in these works: Table 1 in Wuyts et al.
-(2012); Section 3.1 in Tadaki et al.
-(2014); Figure 10 in Guo et al.
-(2015); Figure 4 in Shibuya et al.
-(2016); Figure 8 in Ribeiro et al.
-(2017); Figure 16 in Huertas-Company et al.
-(2020); Figure 7 in Sattari et al.
-(2023); Figure 12 in Tan et al.
-(2024); and Section 4.1 in Harikane et al.
-(2025).
+Appendix B Treating Clumps as Point Sources in JWST Images
+In this Appendix, we briefly describe how clumps are identified in JWST images when they are considered to be point sources.
+While clumps in JWST data typically have extended light profiles (Section IV.2), in order to compare $f_{\rm{clumpy}}$ measurements from HST and JWST images (Section V.3), clumps must be treated as point sources.
+This is because they were largely found to have light profiles consistent with point sources in earlier work based on HST.
+To identify a complete sample of clumps under the assumption that they resemble PSFs, we perform the same routines to identify clumps and determine the completeness thereof as described in Sections IV.1 and IV.3, with only two exceptions as follows.
+First, aperture corrections are applied assuming the typical light profile of a clump is that of the PSF.
+Aperture corrections are 1.348, 1.307, 1.299, 1.336, and 1.523 in GOODS-N and 1.365, 1.328, 1.330, 1.356, and 1.485 in GOODS-S for the F090W, F115W, F150W, F200W, and F277W bandpasses, respectively.
+Second, to determine the completeness as a function of redshift and stellar mass, PSFs are inserted randomly into real images of galaxies, instead of Sérsic profiles convolved with PSFs.
+The resulting completeness curves are shown in Figure 13.
+It is in the same format as Figure 7.
+The optimal smoothing length is 10 pixels, the same as that when treating clumps as extended objects.
+Complete samples of clumps can be detected at fainter minimum fractional luminosities (as low as $2\%$) when modeled as PSFs.
+This is largely because aperture corrections are smaller.
+Figure 13: Same as in Figure 7, except fake clumps are modeled as PSFs. The optimal smoothing length is 10 pixels (middle row), as was found when treating clumps as extended objects.